[Search for users] [Overall Top Noters] [List of all Conferences] [Download this site]

Conference asdg::cmp

Title:CMP
Notice:Welcome to the CMP notesfile
Moderator:FABSIX::B_TOWERLLE
Created:Tue Oct 31 1995
Last Modified:Fri Jun 06 1997
Last Successful Update:Fri Jun 06 1997
Number of topics:94
Total number of notes:6381

7.0. "CMP.A2 EE/PE PASSDOWNS" by SUBPAC::BJUBINVILLE (NFD Fire - Rescue) Fri Nov 17 1995 05:25

    
    
       This note is designated for the passdowns from Equipment
    and Process Engineering for:
    
                                CMP.A2
    
T.RTitleUserPersonal
Name
DateLines
7.86SUBPAC::LANDRYMon Nov 06 1995 19:4817
    Late this morning, manufacturing surrendered the system to EE so that
    the Monthly PM could be performed.  There still remains a few items
    that need to be performed (check sheet has been forwarded to the B
    shift people of remaining items).
    
    A couple of interesting things happened while doing the PM.  Down force
    was extremely difficult.  The left spindle down force (DF) was out of
    spec, it required three passes to fall within spec.  The right
    spindle DF was way out.  So far out, in fact, that the E-P regulator
    calibration procedure needed to be performed.  Once this was completed,
    the DF cal took ~5 passes to tweak into spec.  Unbelieveable!
    
    The slurry pumps will need to have the flow rates checked, as the lines
    were replaced at the pump head.  
    
    					A shift
    
7.87High Table Amps / wafer fly outsFABSIX::R_GEEWed Nov 08 1995 07:0327
	* After the system had been idle for sometime, breakin wafers
	  were run. On the first set of two a wafer slid out from under
	  the left carrier. The system was EMO'ed before it crashed.

	      System status at the time:

		- 24 wafers on the pad (2 suba IV, 1 IC1000)
		  10 wafers run before the counters were reset, counter 
		  is at 14 wafers.
		- 7 wafers on the carriers. (Exposure .0065 & .0070)checked.
		- Engineering carrier film mounted.
		- Table amps at the time spiked to 57 amps.
		- Checked slurry flow, slightly low 450 ml/min at 500 setting,
		  adjusted.  
		- Down force was 313 lbs
		- Slurry flow set at 417 ml/min
		- 20 sec into a 180 sec polish.
		- Back pressure 14.8
		- Believe the wafers to be oxide monitors.

	The system is now unscheduled down.

					B shift


	   

7.88passdownSUBPAC::LANDRYWed Nov 08 1995 19:3214


    Came in this morning to the system spittin' wafers from the left
    spindle.  Found broken wafer chips on right clean station brush???
    Replaced the brush on the right side. Rebuilt both carriers, Thanks 
    Earl. 
    
    Checked the E-P regulator pressures and they were  off
     on both spindles.  Brought them in and reperformed the down force
    calibration.  Completed,  and AGAIN its under 1%. Changed pad  Thanks 
    Barry.  MQC's passed   up and running
    
    				jim,doc,kyle,john_p. Brendan
7.89Wafer breakage....SUBPAC::BJUBINVILLENFD Fire - RescueThu Nov 09 1995 08:357
    
          The tool wasn't used for a while because of the
      stall outs.  We up'ed the slurry flow and before
      we could even get a look at some results.  The pad
      replaced as well as the carriers.  The tool needs
      full break-in and MQC's.
                                     D shift
7.90TABLE PROBLEMS STILL EXISTSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 05:359
    
        Another day, another problem...  When Step is all
     done, the table comes to a dead stop before the spindles
     even lift off.  The spindles are still turning as well.
     We went through all of the possible software calibration
     variables and software program settings and found nothing
     out of the ordinary.  The machine is still logged down.
    
                                         D shift
7.91Wafer Fly-Out ExperimentASDG::POIRIERFri Nov 10 1995 19:519
    
    Wafer Fly-Out Experiments:
    
    The system is currently UP to ENG. It experienced the same conditions
    as CMP.A1 (See Note #5 for CMP.A1) Please build one new carrier and
    mount it on the Right spindle. Leave the current pad on the table
    and follow the instructions from Eric Bodensieck.
    
                                                                  John.
7.92Table stall seeing an endSUBPAC::BJUBINVILLENFD Fire - RescueSat Nov 11 1995 07:1816
    
          We started off running the new pad break-in per Eric.
      We ran an additional 50 wafers for a total of 85 wafers with
      no problems.  We went to break and started up some more wafers
      before the MQC's were ran and the table stalled out on the first
      (2) wafers at 313 lbs.  The retaining ring on the right spindle,
      which was brand new was all shiny and appears to have been 
      polished.  After sitting in wet idle for an hour, the pad
      acted like a sponge and was wrapping around the wafers and
      as a result, the retaining ring was causing the table to stall.
      We ran 12 wafwers, (4) @ 200, (4) @ 250 and (4) @ 313l lbs and
      it ran fine.  Theyu ran MQC's and they passed everything and there
      were now stalls.  If you have the tool sitting idle, run the 12
      wafer sequence and it looks like it will be ok.....
    
                                                  D shift
7.93Oil level lowSUBPAC::BJUBINVILLENFD Fire - RescueThu Nov 16 1995 07:297
    
    
         we checked the oil level in the gear box and we couldn't even
      get an oil level.  It will have to be filled as well.  refer to
      5.last for more details.
    
                                                  D shift
7.1MECHANICAL ACCEPTANCESUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:45277
ASDG::POIRIER                                        12 lines  17-AUG-1994 22:09
                           -< Mechanical Acceptance >-
--------------------------------------------------------------------------------
    
    * The tool was powered up for the first time tonight. Eric checked the
      main power connections and took voltage readings.
    
    * I ran 3/8 tubing from the main CDA line to the back of the tool.
    
    * The robot was removed and dis-assembled on the bench. Eric and I 
      exchanged servo encoders with the robot assembly he brought from
      Strasbaugh's field service office. The robot was then re-installed.
    
    
                                                                   John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.9                            CMP.A2                             9 of 296
SUBPAC::BJUBINVILLE "The other side of life -- 3rd"  16 lines   8-SEP-1994 11:23
                             -< Passdown for 8/8 >-
--------------------------------------------------------------------------------
    
       Started working with Eric on setting up this machine.
    
       The machine wouldn't initialize because the upper
    structure sensor wasn't being made, adjusted same.
    
       The shuttle needed to be leveled
    
       We were about to try and indicate the carry arms
    but the motor on the right didn't come down far enough.
    It was was discovered that the motor was catching on the
    sensor bracket.  This bracket needed to be re-adjusted out
    of the way of the robot.  Eric is going to continue indicating
    everything in.
    
    BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.15                           CMP.A2                            15 of 296
ASDG::POIRIER                                        12 lines  15-SEP-1994 23:21
                        -< Mech Acceptance Activities >-
--------------------------------------------------------------------------------
    
    * The damaged tub in recieve station #2 was replaced with a new one.
    
    * Table Gearbox: Strasbaugh is concerned about the possibility of 
      excess heat in the gear box bearing now that the shims have been
      removed. This action has increased the pre-load on the bearing.
      The Strasbaugh rep is planning to run the table at 150 rpm
      for 8 hours on Monday and monitor the bearing temperature with
      a thermo probe.
    
    
                                                              John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.16                           CMP.A2                            16 of 296
ASDG::POIRIER                                        11 lines  16-SEP-1994 23:24
                            -< Start-up activities >-
--------------------------------------------------------------------------------
    
    * I attempted to level the tool tonight but ran into difficulties.
      The leveling bolts are not holding up to the load that the tool is
      placing on them. I tried turning the left front bolt with a 3/4"
      wrench and it bent over. I will ask the Strasbaugh rep to supply
      some new leveling bolts of at least Grade 5 quality on Monday.
    
    * I started indicating the spindles tonight. The right one is within
      .009". I'll continue with the alignments on Monday.
    
                                                                   John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.23                           CMP.A2                            23 of 296
ASDG::POIRIER                                        13 lines  29-SEP-1994 00:49
                              -< CMP.A2 Start Up >-
--------------------------------------------------------------------------------
    
    * The Strasbaugh rep replaced the bent shaft in the robot lower arm
      assembly but cracked the upper arm during reassembly. A new upper
      arm was shipped from Strasbaugh today and should arrive here tomorrow.
    
    * Elevator heights were reset on all 4 elevator stations. This was
      necessary due to the disassembly of the lower robot arm.
    
    * The Barksdale pressure transducer came in today and was installed
      on the right spindle.
    
                                                                      John.
     

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.24                           CMP.A2                            24 of 296
ASDG::POIRIER                                        23 lines  30-SEP-1994 00:06
                              -< CMP.A2 Start Up >-
--------------------------------------------------------------------------------
    
    * The elevator height calibrations were set.
    
    * The robot top arm was replaced and aligned. This arm is a replacement
      for the one Eric cracked yesterday.
    
    * The robot heights were calibrated.
    
    * The right spindle back pressure is consistantly 50% higher than 
      the the back pressure on the left spindle. We replaced the right
      spindle pressure transducer and followed the normal adjustment
      procedure only to find no change. Eric will continue trouble
      shooting this problem tomorrow.
    
    * We discovered a problem with the white plastic rings on the carriers.
      They appear to be warping. This causes the the carriers to not seat
      properly into the load stations. Some possible causes of the warping
      could be overtightening the screws, helicoils backing out or maybe
      the heads of the screws are dragging in the counterbores in the 
      rings. Can anyone shed any light on this subject? 
       
    
                                                                     J.P.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.25                           CMP.A2                            25 of 296
ASDG::POIRIER                                        10 lines  30-SEP-1994 23:54
                              -< CMP.A2 start up >-
--------------------------------------------------------------------------------
    
    * Replaced SOL #31 (D.I. water supply).
    
    * Re-installed the 4 elevator tubs.
    
    * The right spindle is experiencing excessive play. A major disassembly
      of the Right spindle is planned for Monday.
    
                                                                      
                                                                      John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.26                           CMP.A2                            26 of 296
ASDG::POIRIER                                        11 lines   4-OCT-1994 00:01
                              -< CMP.A2 start-up >-
--------------------------------------------------------------------------------
    
    * The load stations, cleaning stations, robot heights & elevator
      heights were "tweaked-in" today. The system was able to run 100
      wafers without an assist. We're planning to finish-up the fine
      tuning tomorrow and run a 500 wafer pre-test on Wednesday.
     
    * The problem with the LOOSE right spindle turned out to be an
      improperly seated 3" snap ring. Eric managed to knock it back into
      place. He will have to realign the right spindle tomorrow.
    
                                                                 John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.29                           CMP.A2                            29 of 296
ASDG::POIRIER                                        19 lines  12-OCT-1994 23:58
                        -< 1000 wafer run preparations >-
--------------------------------------------------------------------------------
    
    * A 150 wafer (pre test) run was sucessfully completed @ 10:30 p.m.
      
      The 1000 wafer marathon run is now scheduled for tomorrow. This run
      will include a 1 min polish on each wafer and will take approx
      28 hours to complete.
    
    * All monitor wafers used in the pre test were cleaned on the OnTrak
      scrubbers after use. They are now on the SEND elevators in CMP.A2.
      
    
    * The Strasbaugh rep replaced the robot end effector arm and related
      vacuum hoses. It appears that the installation of the hoses into the
      end effector (at the Factory) was not done properly. Eric feels that 
      this contributed to the intermittent end effector vacuum errors we
      experienced yesterday. The error did not occur during tonight's 150
      wafer run. 
    
                                                                    John. 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.30                           CMP.A2                            30 of 296
ASDG::POIRIER                                        19 lines  13-OCT-1994 23:58
                -< 1000 wfer run / Carrier clamp ring problems >-
--------------------------------------------------------------------------------
    
    * The 1000 wafer marathon run is in progress. 182 wafers have been 
      completed as of 11:30 p.m. Brian Jubinville is taking over for
      3rd shift. Please relieve him as early as possible tomorrow morning.
    
                                                                   
    
    * The plastic carrier clamp rings are starting to distort out of shape.
      We measured them at 9.765" and 9.770".  The spec is 9.750". 
      This extra .015" to .020" is starting to cause interference problems
      in the load stations. We had to change the "DOWN SOFT" calibration
      variable to compensate for the tight fit. One possible cause of the 
      distortion is over tightening. I can order a torque limiting screw
      driver to get some control on the screws. Another option, suggested
      by Eric, is to replace the 10-32 X 3/8 SS screws with shoulder
      screws.  This will allow the rings to compress just enough to keep 
      them tight without any distortion.  
      
                                                                     John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.32                           CMP.A2                            32 of 296
ASDG::POIRIER                                        13 lines  14-OCT-1994 23:21
                                  -< CMP.A2 >-
--------------------------------------------------------------------------------
    
    * The tool ran 262 wafers then errored for END EFFECTOR VACUUM. It 
      appears that the air valve is acting up again. Eric has ordered a 
      replacment valve from Strasbaugh. We're  planning to re-start the 1000 
      wafer run on Monday. 
      
    * All monitor wafers used on the A1 & A2 have been cleaned on the
      OnTraks.
    
    * The sytem was left in wet idle mode.
    
                                                                   John.
     

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.33                           CMP.A2                            33 of 296
ASDG::POIRIER                                         6 lines  18-OCT-1994 00:19
                          -< 1000 wafer marathon run >-
--------------------------------------------------------------------------------
    
    * The 1000 wafer marathon run is continuing to run into 3rd shift. 
      A2 has sucessfully completed 372 wafers as of 12:15 a.m. with no
      operator assists. Brian is taking over until 1st shift comes in.
    
                                                                John. 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.36                           CMP.A2                            36 of 296
ASDG::POIRIER                                        13 lines  18-OCT-1994 23:35
                     -< 1000 wafer marathon run complted >-
--------------------------------------------------------------------------------
    
    * The 1000 wafer marathon run was sucessfully completed at 9:40 p.m.
      tonight. There was one operator assist during the run. It happened on
      1st shift today and was caused by the shuttle assembly not completing
      its full travel. We are aware of a slight interference between two
      pieces on sheet metal on the shuttle assembly when it indexes.
      I'm not sure if this contributed to the shuttle failure or not.
      The only thing I noticed about the shuttle tonight was a slight
      hesitation during its travel. We are planning to rework the sheet
      metal covers in the near future.
    
                                                                     John.
    
7.2GEAR BOX PROBLEMSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:4557
ASDG::POIRIER                                        18 lines   8-SEP-1994 23:42
                          -< Table Gear Box  Problem >-
--------------------------------------------------------------------------------
    
    * A problem was encountered by the Strasbaugh rep while performing
      the table alignment procedure today. The table has .018" play
      after tightening all the bolts. We called Strasbaugh and were
      informed that the spec is .002". After a lengthy inspection,
      we determined that the play is in Table gear reduction box.
      
      Strasbaugh instructed us to remove the top plate from the gear box
      and remove one shim. I removed the plate and found no shims present.
      Maybe the contact at Strasbaugh was mis-informed. Anyway, the rep
      will resume work on this issue first thing in the morning. 
    
      We may end up removing the table and the gear box. 
    
    
                                                                     John.
     
      

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.11                           CMP.A2                            11 of 296
ASDG::POIRIER                                        28 lines   9-SEP-1994 23:44
                            -< Start-up Activities >-
--------------------------------------------------------------------------------
    
    * The gear box problem was resolved by the Strasbaugh rep today. Eric
      removed the gearbox top plate and removed some shims. The shaft play
      was reduced from .018" to .0015". This is well within the .002" spec.
    
    * Eric also replaced the Right Spindle Air Cylinder. The cylinders on
      this tool have the same problem that the cylinders on CMP.A1 had. The
      lever arm adjustment was set incorrectly at the factory causing the
      cylinder housing to rub against the bridge box. This induced a
      lateral force on the cylinder piston causing the guide bushing to slide
      inside the cylinder. The cylinders are not permenantly damaged, they
      will be sent back to Control Air co. and reworked.
    
    * Eric brought a new elevator tub with him from conn. This is a
      replacment for the one with a cracked weld.
    
    * I recieved a call from Eric Jacobson (Strasbaugh field service
      manager), he is shipping us a newly designed support bracket to
      stiffin-up the base. This bracket is intended to reduce the
      oscilation problem that the conditioning arm is currently experiencing
      due to the machine base flexing at the point where the conditioning
      arm bolts to the base. The new bracket will be here on Monday. 
    
    * The Strasbaugh rep will be back on Tuesday morning to continue the
      start-up.
    
    
                                                                   John.
7.3VACUUM CABINETSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:4616
ASDG::POIRIER                                        13 lines  13-SEP-1994 23:35
                              -< Vacuum Cabinet >-
--------------------------------------------------------------------------------
    
    * The vacuum cabinet is now functioning and supplying vaccum to the
      tool.
    
    * The vacuum cabinet pressure swith was adjusted to provide the proper
      pressure differential.
    
    * The table was aligned to the spindle.
    
    * The replacement recieve tub was moved into the Service chase. (On
      table)
    
                                                                   John.
7.4Alignments & Leveling bolts SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:4725
ASDG::POIRIER                                        22 lines  20-SEP-1994 23:39
                        -< Alignments & Leveling bolts >-
--------------------------------------------------------------------------------
    
    * Both left and right spindles were aligned today.
    
    * The left and right spindle hubs were also aligned.
    
    * The Strasbaugh rep has a custom indicator support rod that he uses
      to align the spindles. It makes the spindle alignment task much
      easier than the procedure described in the manual. I drew a sketch
      of Eric's indicator support rod and will drop it off at the model
      shop in Shrewsbury tomorrow.
    
    * I looked into upgrading the leveling bolts on the 2 Strasbaughs
      today. The current leveling bolts are 1/2 -13 X 5"grade #2 bolts,
      threaded the entire lenth of the bolt. Grade #3 or better are only
      threaded along the first 2" of the bolt. Therefore, it appears that
      we'e not going to find "off the shelf" grade #5 bolts for this
      problem. I'm going to call Dan Trogen at Strasbaugh tomorrow and 
      ask him to supply us with new bolts of grade #5 quality. He might have 
      special order them.  
        
                                                                     John.
                                                                 
7.5SPINDLE STOPPINGSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:4728
ASDG::POIRIER                                        25 lines  23-SEP-1994 16:56
                           -< start-up activities. >-
--------------------------------------------------------------------------------
    
    * We experienced a problem with the left spindle stopping by itself.
      We swapped cables and the problem persisted. We then switched
      amplifiers and the problem stopped. We're still not positive of
      the cause, but it was most likely a bad connection.
    
    * We completed a 1 wafer run this afternoon. We ran a course cycle
      that still requires much fine tuning.
    
    * The robot lead screw is very noisy when the robot transverses back
      and forth. This is the same problem we encountered on CMP.A1. The
      problem turned out to be a loose lead screw bracket. 
    
    * We discovered a bad Barksdale pressure switch on the right spindle.  
      Eric will bring a replacement when he returns next Wednesday.
    
    * It looks like we'll need a 10-32 heli-coil extractor after all.
      We attempted to work this issue many months ago but Strasbaugh 
      insisted that the silver plated heli-coils in the carriers were
      extreamly reliable and that they never heard of one failing.
      We now have 2 carriers down for stripped heli-coils.
    
    
                                                                   John.
    
7.6Calibration variable for table rotation SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:4718
ASDG::POIRIER                                        15 lines  20-OCT-1994 00:05
                  -< Calibration variable for table rotation >-
--------------------------------------------------------------------------------
    
    * The table would not rotate while the tool was in WET IDLE mode. 
      I called Eric and discovered that there is a calibration variable
      that is not listed on our documentation. (D1 49) will enable or 
      disable the table rotation during wet idle.
    
    * The bracket for the capacitance sensor arrived from Strasbaugh today.
      Eric will be in tomorrow to install it.
    
    * The 90 deg offset tips for the internal retaining ring pliers should
      arrive tomorrow. This will allow us to remove the 4" retaining ring
      on the right spindle. We want to disassemble the spindle and find out
      where the excess play is comming from. 
    
                                                                      John.
7.7ROBOT BINDING UPSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:4820
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"   17 lines  20-OCT-1994 16:44
                            -< Passdown for 10/20 >-
--------------------------------------------------------------------------------
    
    
        Well, we started to level the plate under the robot 
    since it was a little out.  When we tried moving the
    robot back and forth, it kept getting stuck in mid
    travel.  At first, I thought the lead screw got hit
    but that wasn't the case.  This happened to Eric and 
    John before.  They had put some lubricant on the rails
    to fix it and make it run smooth.  This is what happened
    to us.  Eric decided to pull the entire robot and plate
    right off of the machine.  We cleaned up everything and
    removed any debris on the screw.  Before we left for the
    Films meeting, Eric and Bob had the everything back in
    and they were leveling things.
    
    BJ
                                          
7.8ELEVATOR ALIGNMENTSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:4831
ASDG::POIRIER                                        28 lines  21-OCT-1994 00:07
                             -< CMP.A2 activities >-
--------------------------------------------------------------------------------
    
   * The cassette sensor arm on SEND elevator #1 is interfering with the 
     H-bar on the cassettes. This creates a condition where it appears that
     the elevator has initialized but in fact has not because the micro 
     switch on the sensor arm does not make contact. There are no adjustments
     on the cast aluminum sensor arm assembly to rectify this situation.
     However, we could remove a little stock from the teflon locating blocks
     on the elevator platform. I think 1/32" would be enough.
    
   * The elevator heights on all 4 elevators were checked and adjusted as
     needed.
    
   * The pad was removed in preparation for the right spindle alignment 
     procedure. I was about to start the alignment when I noticed that the 
     retaining ring on the right spindle was removed. Eric told me that the
     spindle would be put back together and to re-align the spindle when I
     returned from the meeting. I'm not sure if he changed his mind or
     possibly found a new problem with the spindle. In any case, I didn't
     want to expend all that effort if the spindle was not ready to be
     re-aligned. Eric will be back tomorrow  moring so I guess we'll find
     out whats going on then.
    
    * Please let Eric know that the capacitance sennsor for A2 and the 
      conditioner amplifier for A1 are in the 2nd draw of the remote
      control panel.  
    
     
                                                                  John
7.9Robot lead screw assembly bushings SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:4912
ASDG::POIRIER                                         9 lines   3-NOV-1994 00:28
                    -< Robot lead screw assembly bushings >-
--------------------------------------------------------------------------------
    
    * I replaced the robot lead screw assembly bushings with the new
      ball bearing style tonight. They arrived from Strasbaugh today
      along with new o-rings for the spindle chucks. The old bushings
      are on Jim Tooker's desk.
    
    
                                                              John.
    
7.10ROBOT CLEANED AND ALIGNEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:4912
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"    9 lines   3-NOV-1994 16:19
                        -< 1st shift passdown for 11/3 >-
--------------------------------------------------------------------------------
    
    
          Re-installed the robot and drive assembly.  We
    went through all of the different stations to verify
    calibration.  We made some minor adjustments involving
    positioning of the robot.  We've been running exercise
    wafers all afternoon without any problems.
    
                                             1st shift
7.11#1 Cassette position sensor SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:507
PATE::GEE                                             4 lines  15-NOV-1994 22:55
                        -< #1 Cassette position sensor >-
--------------------------------------------------------------------------------
    The #1 send cassette position sensor wasn't seating properly onto the
    cassette. Had to realigned the sensor bracket.
    
    Bob  
7.12lead screws SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:509
PATE::GEE                                             6 lines  16-NOV-1994 23:55
                                -< lead screws >-
--------------------------------------------------------------------------------
    Receive elevator #2 was binding during its travel. Lubricated all 4
    lead screws on both cmp.a1 and cmp.a2. 
    
    Also, both robot lead screws where lubricated.  
    
    	2nd shift
7.13Robot Z-axis calibration variables SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5116
ASDG::POIRIER                                        13 lines  28-NOV-1994 23:54
                    -< Robot Z-axis calibration variables >-
--------------------------------------------------------------------------------
    
    * It appears that the robot (Z height)is off. I changed the Robot
      Z-axis TOP position calibration variable from .17 to .26 and the 
      robot picked up the wafers from slot #1 thru slot #13 with no
      problems. It starts clipping the wafers from slot 15 thru 25 so it
      looks like the Z-axis MIDDLE and the Z-axis BOTTOM calibration
      variables will have to be changed as well.
     
                                                             John.
    
     
        
    
7.14ROBOT NOT PICKING UP WAFERSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:51108
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"   15 lines  30-NOV-1994 16:25
                            -< Passdown for 11/30 >-
--------------------------------------------------------------------------------
    
    
       Today...  Eric found the end effector vacuum line was 
    hanging down causing the robot to hit the wafer.  Wether
    or not this like was like this all week is questionable.
    A new style arm is coming in next and will be replaced
    on both systems.  We traced all of the input and output 
    sources for the fountains on the LOAD station.  We tweaked 
    the 24V because it was low.  Eric ended up realizing that 
    the current software was never updated to the newer version 
    and he'll do that when the disk comes in.  Eric and John are
    currently tracing a DI water sensor chattering that exists 
    on both machines.
                                            1st shift
                                            

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.60                           CMP.A2                            60 of 296
ASDG::POIRIER                                        34 lines  30-NOV-1994 00:23
                  -< robot Z-height/fountian rinse problems >-
--------------------------------------------------------------------------------
    
    *Load station fountian rinse problem:
     ------------------------------------
     We checked output opto's #158 & #159 during the fountian rinse 
     cycle. The LED's were on.
    
     The connections on the terminal block (next to the solenoid valves),
     were checked. Both #158 & 159 were good.
    
     We checked for 24 V at solenoid valve #72 (right load station fountian
     rinse), 24 V was present.
    
     I operated solenoid valve #72 manually and the fountian rinse sprayed 
     nicely. When the carrier rinse solenoids energized, the fountian rinse
     flow dropped to zero. This may be due to the abnormally low D.I. water 
     pressure of 20 p.s.i. we had during all of 2nd shift. Therefore, the
     valves may be energizing after all, it just doesn't appear so. I
     didn't have time to verify this theory before leaving. Please check it
     out in the morning.
    
     Elevator Z-height problem:
     -------------------------
     The robot continued to give us intermittent problems tonight. We ran
     many auto cycles during the shift and encountered different Z-height
     position problems at various times. The robot picked up wafers from
     both send stations with no problems on some runs and clipped wafers on
     other runs. Rehomming the system didn't appear to have any effect on
     this condition. 
    
     Eric Ronnebaum will be comming in tomorrow morning to take a look
     at this problem.
     
    
                                                                   John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.62                           CMP.A2                            62 of 296
ASDG::POIRIER                                        40 lines  30-NOV-1994 23:56
                            -< 2nd shift passdown >-
--------------------------------------------------------------------------------
    
    * Wafer handling problem:
      ----------------------
    
      Test runs were run using both send stations and both recieve
      stations. Recieve station #1 is still clipping wafers slightly
      in cassette positions 1 & 2. Eric Ronnebaum will be in tomorrow 
      morning to tweak it in. I am leaving the tool DOWN in RTC until
      Eric tweaks it in tomorrow morning.
    
    
    
    * Fountian rinse problem:
      ----------------------
    
      Eric discoverd that CMP.A2 has version 3.0 software loaded instead of
      version 3.6. With 3.0, the fountian rinse shuts off before the
      carriers lift up from the load stations. With this condition, it
      appears that the fountian rinse is not functioning when in fact it
      is. It just shuts off sooner. Strasbaugh is shipping version 3.6 
      to us ASAP. 
       
    
    
      Elevator squeeling/hesitation problem:
      --------------------------------------
      Eric thinks that the problem lies with the elevator assembly
      alignment. He doubts that the lead screw bushing is the problem.
      He is planning to demonstrate the alignment procedure for us  tomorrow.
    
    
      PLEASE NOTE:
      -----------
      A wafer came off the carrier during a test run @ 11:00 p.m. We lost 2
      wafers and the pad. The carriers involved are on the stainless table
      next to A1. The carriers need to be rebuilt and the retaining rings
      are damaged.
    
    
                                                                  John
7.15ELEVATOR HEIGHT NOT REPEATINGSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5112
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"    9 lines   1-DEC-1994 14:36
                             -< Passdown for 12/1 >-
--------------------------------------------------------------------------------
     
       Eric traced the elevator non-repeating problem to the
    pulley that is on the end of the lead screw.  The set screw 
    had loosened up.  He tigtened it back down and tweaked 
    everything back in.  There are harder pulleys that are being
    shipped in and Eric will be replacing all of the pulleys on
    both A1 and A2.
    
                                                 1st shift
7.16ELEVATOR HEIGHT NOT REPEATINGSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5252
ASDG::POIRIER                                        49 lines   1-DEC-1994 23:58
                            -< 2nd shift passdown >-
--------------------------------------------------------------------------------
    
    * Elevator problem:
      ----------------
      Eric found a loose set screw on the pulley that drives the elevator
      on the left RECIEVE station. He also found a crack in the( plastic )
      pulley. The set screw was tightened and the elevator was reinstalled.
      We expect the pulley to function for the short period of time it will
      take for the new one to arive from Strasbaugh.
                                  
    
    
                                                                  
    
    
      Elevator squeeling problem:
      --------------------------
      Eric demonstrated the elevator removal and disassembly procedure
      tonight. It appears that the squeeling is caused by dirty/grimey
      build up on the lead screw. Another contributing factor is the
      small twist we found in the elevator backing plate. This plate 
      has the guide rails and the ball nut bracket bolted to it. Any
      deformation in this plate will cause binding. As usual, Strasbaugh
      used aluminum for the backing plate material and as usual its
      distorted out of tolerance. We place the precision level across it 
      and its twisted about an 1/8 ". We throughly cleaned the lead screw,
      guide rails and the ball nut. It now travels with no sqeeling.
    
    
     Fountian rinse problem:
     ----------------------
     We installed version 3.08 software and the fountian rinse now
     functions in the same manner as CMP.A1.
    
     Carriers:
     --------
     Jim Martin and Jesus Roa trained Bob and myself on carrier rebuilding
     tonight. We rebuilt the two carriers that were involved in last
     night's run that resulted in wafer breakage.
    
     Valve chattering problem:
     ------------------------
     Eric called Strasbaugh about our continuing problems with the D.I.
     water soleniod valves chattering. The people at Strasbaugh informed
     Eric that other 6DS customers have been complaining about the same
     problem. They are looking into this issue and will retrofit the tools
     when they develop a solution. 
    
    
                                                                      John.
7.17Elevators PULLEYSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5217
ASDG::POIRIER                                        14 lines   2-DEC-1994 23:29
                       -< 2nd shift passdown/Elevators >-
--------------------------------------------------------------------------------
    
    * All four elevator assemblies were removed from the tool. They were
      disassembled, cleaned, lubricated and reassembled. We also replaced
      both plastic pulleys on all of the elevators with the new aluminum 
      ones that arrived from Strasbaugh today. We ran test runs utilizing
      both SEND and both RECIEVE stations at the end of the shift. The
      elevators moved smoothly with no sqeeling. The robot loaded and
      unloaded with no clipping in all four cassettes. The rest of the new
      pulleys are on Jim Tooker's desk.
    
                                                     Bob & John.
     
    
    
7.18SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5334
ASDG::POIRIER                                        31 lines  20-DEC-1994 01:28
                            -< servo amp problems >-
--------------------------------------------------------------------------------
    
    * The system went down tonight for a (AXIS #2 SERVO AMPLIFIER FAULT)
      error. The table and spindles were stopping during the auto polish
      cycle. We suspect that the problem was with the servo amp
      connections. The amps had oversized lugs on them when the tool first
      arrived. These oversized lugs came into contact with the inside of
      the electrical box door which in turn bent the connectors on the amps.
      We replaced the lugs with smaller ones 3 months ago but the
      connectors may be permanently damaged causing intermittent failures
      with the servo amps. We repositoned the connector on servo amp #2
      and the problem ceased to exist. 
    
      The repair was delayed by the inability to get 2 good carriers to 
      test the tool. One of the carriers gave us a lot of trouble, we had
      to take it apart 3 times to find the problem. We ended up back
      flushing the 2 check valves in the carrier before it would work 
      properly. Special thanks to Jim Martin for staying late and helping
      us with the carriers.
      
    
      * We also had trouble with the carrier on the left spindle loading
        wafers correctly. We repositioned the left load station to work
        with the carrier on the left spindle and the problem went away.
        I think the problem lies with the carrier clamp rings. They are
        distorted out of shape and tend to cause loading problems in the
        load stations. We can tweak either load station to fit a 
        particular carrier, but the problem will return when a new carrier
        is mounted on the spindle. The clamp rings need to be redesigned 
        so they don't distort or they should be changed more frequently.
    
                                                            John & Bob.
7.19SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:538
ASDG::POIRIER                                         5 lines  21-DEC-1994 00:14
                      -< WAFER DETECTED ON SPINDLE error >-
--------------------------------------------------------------------------------
    
    * The system errored for (WAFER DETECTED ON SPINDLE). The problem 
      turned out to be a malfunctioning carrier.
    
                                                              John
7.20301 = IBM MANUAL ERRORSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5315
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"   12 lines   2-JAN-1995 08:28
                             -< Passdown for 1/2 >-
--------------------------------------------------------------------------------
    
    
       This morning at power-up, the screens kept coming up
    
    with a 301 = IBM MANUAL.  I traced it to the roll around
    
    unit.  After I got A2 running in wet mode, I looked into
    
    the roll around unit and traced it to being the keyboard
    
    being unplugged.  
                                            BJ 
7.21SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5414
ASDG::POIRIER                                        11 lines  10-JAN-1995 01:28
                     -< Conditioner actuator reed switch >-
--------------------------------------------------------------------------------
    
    * PROBLEM: CONDITIONER DISK CONTACTING SIDE OF PARK STATION DISH.
    
       ACTION: The reed switch on the side of the conditioner actuator
               does not appear to be operating properly. The problem
               may be with the switch itself or with the magnet that
               is mounted on the inside of the actuator cylinder 
               housing. Brian will continue to troubleshoot on 3rd 
               shift.
    
                                                              John. 
7.22SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5414
ASDG::POIRIER                                        11 lines  10-JAN-1995 23:56
                         -< conditioning arm problems >-
--------------------------------------------------------------------------------
    
    * Adjusted the mechanical travel in the conditioning arm actuator 
      assembly. The conditioning disk now lands in the center of the
      park station.
    
    * The conditioning arm homing failure problem turned out to 
      be a bad connection on the terminal strip located next
      to the conditoning arm actuator.
    
    
                                                              John & Jim M.
7.23LOSS OF POWERSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5420
SUBPAC::LANDRY                                       17 lines  12-JAN-1995 16:37
                           -< wheres all the juice? >-
--------------------------------------------------------------------------------
    A couple of crazy things with machine power went on today:
    
    *	The system powered down totally today.  Found one leg of the 208V
    	incoming power to be off.  What?  The wall breaker wasn't tripped
    	though.  What?  So, shut said breaker off and turned back on.
    	That brought the "power on" light back, but the system still would
        not power up.  Traced the problem to be an open interlock, located
    	on the back/lower door at the polish table end.  Mechanically 
    	re-positioned the doors bracket that hits the switch.  All set!
    
    *   As Ops Techs returned from dining half-hour, the found the system
    	powered down.  The system power on light was lit though.  Well,
    	back to the interlocks.  Re-adjusted the previous problem bracket
    	as well as the bracket on the door with one missing latching mech-
    	anism.  The system has stayed powered up since.
    
    						1st shifters 
7.24COMPUTER WON'T BOOTSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5616
PATE::GEE                                            13 lines  12-JAN-1995 23:36
                                 -< Computer >-
--------------------------------------------------------------------------------
PROBLEM: System would not boot at initialization

root cause:  board not seated in computer

action: all boards removed and reseated.


Jim Martin taught John P. and myself how to backup all files from the Computer.
John has the floppys.

2nd Shift


7.25POWER PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5631
SUBPAC::LANDRY                                       28 lines  13-JAN-1995 15:12
              -< A Fun and Educational Day, wish you were here! >-
--------------------------------------------------------------------------------
    Arrived this morning to be greeted with a re-occuring problem.  Yup,
    the system has just powered itself down.  OK, let get to it.  As was
    the case from the previous day, one of the legs coming from the wall
    breaker (90 AMP) was not supplying the proper current to the system.
    The day before, we reset the breaker and the system survive about 16
    hours until it happened again.  This time, the breaker wasn't reset.
    A call was logged to have facilities check out the breaker prior  to
    us attempting to track down the problem.  The electrician's first
    reaction was that something on the system was taking it down.  Well,
    I was skeptical cause the power to the system wasn't going to far
    through the system.  We couldn't even get the "power on" light to
    light up.  Thats one of the first things to see juice.  Well, back to
    the wall breaker box we go.  The suspected breaker was removed.  Lo
    and behold, it had signs of mechanical failure.  The electrician
    quickly rounded up an new one, installed it and voila, power!  Joe,
    the electrician, did a great job.  
    
    Next thing was to fire up the system and initialize everything.  The
    system came up with a 301 error.  Whats that?  After a  short delay,
    the screen suddenly showed a cute little hint.  This is what it said:
                  /
    	301 = ( O/K )  >>----->  IBM manual
               /
    Basically, to interpret, it meant 301 = not ok, RTFM (read the manual).
    Found a jack on the back of the system just barely hanging on.  Plugged
    it fully in and rebooted the system.  Alls well at this stage.
    
    							Doc
7.26COMPUTER WON'T BOOTSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5625
ASDG::POIRIER                                        22 lines  13-JAN-1995 23:48
                            -< PC boot-up problems >-
--------------------------------------------------------------------------------
    
    *    Problem: System fails to boot-up approx 30% of the time.
                  ERROR: Board 0 failure.
     
      Root cause: Suspected PC problem.
    
    
          Action: a) Reloaded Initialization files. We thought they might
                     be corrupted.
    
                  b) Opened up PC and removed, inspected and reseated
                     boards. (No schematics available to identify boards.)
    
      
     Current status: a) No improvement. 
    
                     b) Will place call to Strasbaugh field service for
                        board identification and related information.
    
     
                                                             2nd Shift.
    
7.27SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5625
ASDG::POIRIER                                        22 lines  17-JAN-1995 00:00
                    -< conditioner down force calibration >-
--------------------------------------------------------------------------------
    
    Problem: Conditioner Down Force out of calibration.
    
    
      Cause: Needs full calibration. Possible amplifier problem.
    
     
     Action: Followed procedure in Manual. The machine would not respond to
             the commands outlined in the Manual. We called Eric Ronnebaum
             at his home and he informed us that the functions we were
             trying to activate are now contolled by the DCX controller
             not by Analog inputs as stated in the Manual. 
    
             We were able to continue with the procedure until we tried
             to disable the Conditioner Down Force Control with the FLAGG
             function. We cannot locate the Flagg number for this function
             in any of the documentation. Will someone on 1st shift please
             call Strasbaugh for this information.
    
                                                           Bob & John
    
      
7.28ROBOT ARMS NOT ARRIVING AT ZEROSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5712
SUBPAC::BJUBINVILLE "IT'S HOCKEY TIME!!"              9 lines  18-JAN-1995 09:07
                        -< 3RD SHIFT PASSDOWN FOR 1/18 >-
--------------------------------------------------------------------------------
    
        The robot arms were not arriving at the right load station 
    
    all of the way and the machine kept erroring out.  After the
    
    robot arms were referenced, the error went away.  An assist was
    
    logged for this event.
                                                  BJ
7.29SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5827
PATE::GEE                                            24 lines  19-JAN-1995 00:11
                              -< Pad Conditioner >-
--------------------------------------------------------------------------------
    
	PROBLEM: The pad conditioner doesn't submerge into the
                 park station reservoir after it conditions the
                 pad, unless you manually move it. 

          CAUSE: Found the linear bearings need to be replaced.

        ACTION: Disassembled, cleaned and reassembled the pad conditioner 
                assembly. We found that all the bearings in the housing
                are rusted and that the linear bearings need to be replaced. 
                
                There are no linear bearing in-house so we will rush them
                in tomorrow, they should be here Friday.

                After cleaning the bearings, the pad conditioner is 
                submerging about a 1/4" and can be run. Please monitor
                it, and log an assist if you have to manually push it
                down. The system has been put on Warning in RTC.


						John and Bob
    
        
          
7.30SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 04:5916
ASDG::POIRIER                                        13 lines  19-JAN-1995 23:29
                      -< conditioning disk shaft bellows >-
--------------------------------------------------------------------------------
    
    * The bellows on the conditioning disk assembly was twisted around it's
      shaft. The conditioning disk continued to rotate without any problems
      and the neoprene bellows was undamaged. We disassembled the lower
      end of the shaft assembly and found a spacer in upside down. We 
      reassembled it upside down last night on 2nd shift. It's working
      fine now.
                                  
                                                         
    * Bob and I rebuilt 2 carriers while Jim Martin ran product on CMP.A2.
    
    
                                                                John & Bob. 
7.31SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:0115
ASDG::POIRIER                                        12 lines  26-JAN-1995 23:30
                           -< Spindle Vacuum Switch >-
--------------------------------------------------------------------------------
    
    * Recalibrated Barksdale pressure switch on left Spindle. The system
      now detects and alarms for low vacuum.
    
    * Replaced both left and right spindle vacuum filters.
    
    
    * The system will remain in a DOWN status until the cleaning station
      pads are installed. 
    
    
                                                              John & Bob.
7.32DOWNFORCE ISSUESSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:0180
SUBPAC::BJUBINVILLE "IT'S HOCKEY TIME!!"             77 lines  20-JAN-1995 06:34
                             -< PASSDOWN FOR 1/20 >-
--------------------------------------------------------------------------------


   ENG suspected that the spindle down force and aligment might

have problems due to examining some previous lots run on CMP.A2.

Both spindle down forces and were found to be both out of spec.

The left spindle was out 10% and the right was out 3.5%.  Both 

spindles were brought into spec.  The spindle/hub alignment was

checked on both spindles.  The left spindle was within spec.

The right was out by 2 mils.  The right spindle was adjusted 

and was brought into spec.


Listed below is the reading from the down force.  Hit NEXT UNSEEN

if you don't like numbers and charts.......


LEFT SPINDLE DOWN FORCE - FINAL TEST
====================================

LBS.       DISPLAY     ACTUAL     ACTUAL+TEST VALUE   ERROR VALUE 
=================================================================
50         58.5        40.7       58.4                .1
=================================================================
100        99.9        81.9       99.6                .3
================================================================= 
150        149.9       132.8      150.5               -.6
=================================================================
200        199.9       182.4      200.1               -.2
=================================================================
250        249.8       233.3      250.9               -.9     
=================================================================
300        300.1       283.3      301                 -.9
=================================================================
350        350         334.1      351.7               -1.7
=================================================================
400        400         384.6      402.3               -2.3
=================================================================
450        450.2       434.5      453.2               -3
=================================================================
500        499.9       485.6      503.3               -3.1
=================================================================
                                                    = -1.14%


RIGHT SPINDLE DOWN FORCE - FINAL TEST
LBS.       DISPLAY     ACTUAL     ACTUAL+TEST VALUE   ERROR VALUE 
=================================================================
50         64.4        43.7       61.7                2.7
=================================================================
100        100         78.9       96.9                3.1
================================================================= 
150        149.9       129.7      147.7               2.2
=================================================================
200        200         179.5      197.5               2.5
=================================================================
250        249.8       230.6      248.6               1.2     
=================================================================
300        299.9       282.2      300.2               -.3
=================================================================
350        350         333.6      351.6               -1.6
=================================================================
400        400.1       384.8      402.8               -2.7
=================================================================
450        450         435.0      453.0               -3
=================================================================
500        499.8       485.0      503.0               -3.2
=================================================================
                                                    = .09%
                                                             
7.33LOAD CUPSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:0239
SUBPAC::BJUBINVILLE "IT'S HOCKEY TIME!!"             36 lines  27-JAN-1995 08:26
                             -< passdown for 1/27 >-
--------------------------------------------------------------------------------
    
         The brushes came in and were installed.  Once the
    
    MQC got under way, the left spindle kept dropping vacuum
    
    when it had a wafer on it.  After inspection of the spindle,
    
    it was discovered that the left load cup was moving around 
    
    everytime the carrier came down.  The top ring is suppose
    
    to move freely but the hold down was cranked down.  Also,
    
    all four load cup mounting bolts were loose and the shims
    
    were slid out from under the cup.  After re-assembly and 
    
    re-aligned the spindle came down smooth.  As of right now,
    
    the vacuum is great on both spindles.  No pad touch downs
    
    were done as of 8am.  Wether or not the case yesterday with 
    
    the wafers sliding out was tied into what was fixed is 
    
    undetermined at this time.  
    
        The top robot arms still isn't also arriving to the right 
    
    load cup.  I ttied 2nd shifts suggestion with lubing up the 
    
    lead screw a little more but that didn't appear to help.  It's
    
    better but still not quite right.
    
                                                        BJ
7.34SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:0415
ASDG::POIRIER                                        12 lines  26-JAN-1995 23:30
                           -< Spindle Vacuum Switch >-
--------------------------------------------------------------------------------
    
    * Recalibrated Barksdale pressure switch on left Spindle. The system
      now detects and alarms for low vacuum.
    
    * Replaced both left and right spindle vacuum filters.
    
    
    * The system will remain in a DOWN status until the cleaning station
      pads are installed. 
    
    
                                                              John & Bob.
7.35SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1012
PATE::GEE                                             9 lines   6-FEB-1995 23:41
                               -< robot bellows >-
--------------------------------------------------------------------------------

	The robot wouldn't home.

	Found the bellows wasn't seated properly, causing the robot to 
 	stop.

	Lubricated the slide rails and reseated the bellows.

						John & Bob
7.36SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1162
ATE::GEE                                            18 lines  13-FEB-1995 23:49
                             -< Robot & Computer >-
--------------------------------------------------------------------------------
    Problem: The robot was slamming the inside right side of the cassette
             at the send stations.
    
    Cause: Found that the vacuum line had slid down and wouldn't allow the
           lower arm to extend.   
    
    
    Action: Repositioned the lines to allow more slack. Cleaned and
    relubricated the robot leadscrew, it was houling.
    -------------------------------------------------------------------------
    
    Problem: Intermittent computer booting problem.
    
    Action: Called the rep. and he said the problem may be one of the
            modules plugging into the top of board "0" and not board 
            "0" itself. We will look into this when time permits.
    
                                                     John & Bob

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.131                          CMP.A2                           131 of 296
SUBPAC::LANDRY                                       33 lines  13-FEB-1995 17:03
                           -< Robots and switches! >-
--------------------------------------------------------------------------------
    A long long day........
    
    The robot arm has a flaw.  The problem is......the arm enters the
    Send cassette to take a wafer, pauses as should, then just as it's
    about to lower to pickup the wafer, it flings to the side of the
    cassette.  Of course it then errors out.  This happens on both of
    the send cassettes.
    
    As I attempted to trouble shoot the problem, the system started
    exhibiting a "left spindle low vacuum" alarm.  So, this had to be
    repaired before the robot could attempt to work.
    
    Found several things had happened concerning the vacuum.  1)  Swapped
    the carriers to confirm they weren't the problem.  They were alright.
    2) The coellessing filter housing was 2/3rds full of water.  Emptied 
    it and tried to run again.  No luck.  3) Removed all four o-rings 
    from the hub assembly, cleaned them and the grooves they fit into, and 
    then re-installed them.  Again no luck.  4)  You with me here?  Yup,
    the vacuum switch.  I believe that the switch that resided in there was
    the original switch we continued to move from spindle to spindle while
    new ones were ordered.  So, I removed the varment and installed a new
    one.  Success!  Now back to square one....the robot.
    
    Began looking into it again and it wasn't consistant at all times.  Ya,
    it was consistant in it's flinging to the cassette side wall, but it
    was also periodically not returning to the ZERO position.  This is
    where I get off and second shift takes the reigns.
    
    My speculation is that we either have a connector problem or a wire
    either pinched or broken inside the robot housing.  Good luck lads.
    
    						Doc
      
7.37SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1015
PATE::GEE                                            12 lines  17-FEB-1995 23:21
                     -< Spindle back pressure calibration >-
--------------------------------------------------------------------------------
    Calibrated the left and right spindle back pressure. The left spindle
    was within spec. The right spindle was out of spec by a large margin.
    
    Left spindle back pressure was within +/- 1 psi of low setpoint and
    within +/- 1 psi of high setpoint.
    
    Right spindle back pressure was + 10 psi of the low setpoint and
    + 15 psi of the high setpoint.
    
    Spec is +/- 1 psi of both the high and low setpoint.
    
    						Bob, John and Jesus
7.38SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1122
SUBPAC::LANDRY                                       19 lines  23-FEB-1995 16:39
                            -< Force Cal performed >-
--------------------------------------------------------------------------------
    A "concern" was voiced today centering around the Force readout on the
    screen when the spindles were in the up position.  Both of the spindles
    read differently (as did the other systems two spindles, all four read
    something different).  Well, I hadn't the knowledge of why there were 
    that way and no one else was quite sure as to what the reading range
    should be.  The Manual had no information around this question.  So, 
    the spindle down force was performed to assure that the spindles were
    in fact reading right.  Other than the normal tweak here and there, all
    was fine with the spindle and it's respective load cell.  
    
    After performing the above procedure, the "idle" force read outs had
    changed but were still not in sink with each other.  The readings just
    tracked with the amount of adjustment made during calibration.  Eric B
    did all the cals with me.  He wants to run some MQC's to see what
    results the adjustments made.  Time will tell.
    
    		Doc, with early BJ assistance and late Eric assistance  
    
    
7.39SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1118
PATE::GEE                                            15 lines  24-FEB-1995 03:39
                   -< Spindle  air cylinder, E-P Regulator >-
--------------------------------------------------------------------------------
  	* Found the air line going to the left spindle air cylinder badly
          pinched, we repaired it.

	* Calibrated the spindle load current so that it reads "0" amps
          when the spindles are up.

	* Adjusted the zero pod on the Fairchild E-P regulators for the left
	  and right spindles so that the spindle forces displayed on the 
          monitor screen both read the same. This had to be changed back
          becouse the actual down force increased 40 lbs on the left and 
	  decreased the right spindle 20 lbs. 
         
          

						 Bob & John
7.40SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:12209
SUBPAC::LANDRY                                      206 lines  24-FEB-1995 09:00
                             -< Updated Variables >-
--------------------------------------------------------------------------------

        CALIBRATION VARIABLES

  1.    BRIDGE HOME POS.= -33.86       
  2.    BRIDGE LOAD POS.= -33.53       
  3.    BRIDGE POLISH POS.= 0          
  4.    BRIDGE MIDDLE POS.= -11        
  5.    TEMP IN SCALE FACTOR= 1        
  6.    TEMP OUT SCALE= 1
  7.    TEMP IN OFFSET= -.8
  8.    TEMP OUT OFFSET= 1.5
  9.    BRUSH SPEED SCALE FACTOR= .719 
  10.   SLURRY 1 FLOW FACTOR= 1.08
  11.   SLURRY 2 FLOW FACTOR= 1.29
  12.   TABLE SHUNT LOE TEMP= 0
  13.   TABLE SHUNT HIGH TEMP= 0
  14.   BRIDGE LEFT CURTAIN POS.= -12
  15.   BRIDGE RIGHT CURTAIN POS.= -25
  16.   NOT USED
  17.   NOT USED
  18.   NOT USED
  19.   NOT USED
  20.   SLURRY 2 FLOW FACTOR= .96
  21.   NOT USED
  22.   NOT USED
  23.   NOT USED
  24.   NOT USED
  25.   NOT USED
  26.   NOT USED
  27.   NOT USED
  28.   NOT USED
  29.   NOT USED
  30.   LEFT PRESS. OUTPUT FACTOR= 6.1
  31.   LEFT PRESS. OUTPUT OFFSET= 975
  32.   LEFT PRESS. MIM.= 800
  33.   LEFT FORCE SCALE FACTOR= .94
  34.   LEFT SPINDLE DEAD WEIGHT= 144.79
  35.   LEFT SPINDLE REDUCE RATIO= 10
  36.   LEFT SPINDLE CURRENT OFFSET= 0
  37.   LEFT PRESSURE MINIMUM, TBL2= 0
  38.   SPINDLES DOWN SOFT VALUE= 1200
  39.   SPINDLES DOWN HARD VALUE= 1100
  40.   RIGHT PRESS. OUTPUT FACTOR= 6.1
  41.   RIGHT PRESS. OUTPUT OFFSET= 975
  42.   RIGHT PRESS. MIM.= 800
  43.   RIGHT FORCE SCALE FACTOR= .946
  44.   RIGHT SPINDLE DEAD WEIGHT= 142.46
  45.   RIGHT SPINDLE REDUCE RATIO= 10
  46.   RIGHT SPINDLE CURRENT OFFSET= .1
  47.   RIGHT PRESSURE MINIMUM= 0
  48.   SPINDLES DOWN EXTRA SOFT VALUE= 0
  49.   NOT USED
  50.   PRESS. CONTROL LOOP GAIN= .75
  51.   MAX. PRESSURE STEP DIVISOR= 30
  52.   FORCE LOOP DERIVATIVE GAIN= 56
  53.   FORCE LOOP NEG DER GAIN= 0
  54.   DOWN FORCE INTEGRAL GAIN= 0
  55.   LEFT FORCE SCALING= 1
  56.   RIGHT FORCE SCALING= 1
  57.   BACK PRESS. CONTROL GAIN= 4
  58.   NOT USED
  59.   NOT USED
  60.   6DS REPEAT TEST (0-NO, 1-YES)= 0
  61.	TABLE TEMP CONTROL GAIN= 35
  62.	TABLE REDUCER RATIO= 7.5
  63.	TABLE CURRENT OFFSET= .9
  64.	FILTER CONST.(0-UP/FAST-SLOW)= 1
  65.	BACKPRES UNITS(0-PSIA, 1-PSIG)= 0
  66.	LT BACKPRESS OFFSET= -.5
  67.	LT BACKPRESS SCALING= 1
  68.	RT BACKPRESS OFFSET= -.9
  69.	RT BACKPRESS SCALING= 1
  70.	COND FORCE LOOP GAIN= 3
  71. 	COND FORCE DEAD WEIGHT (LBS)= 2.436
  72.	COND FORCE SCALE FACTOR= .8
  73.	COND FORCE 0-LBS, DIA (mm)-PSI= 0
  74.   NOT USED
  75.	PUMP 3 FLOW FACTOR= 0
  76.   NOT USED
  77.   NOT USED
  78.   NOT USED
  79.   NOT USED
  80.	AXIS 5 OFFSET TO SLOT 0= 3.08
  81.	AXIS 6 OFFSET TO SLOT 0= 3.1
  82.	AXIS 7 OFFSET TO SLOT 0= 3
  83.	AXIS 8 OFFSET TO SLOT 0= 2.93
  84.	CASSETTE SLOT SPACING= .25
  85.	LEFT LOAD RINSE, 0-DI, 1-CHEM= 0
  86.	RIGHT LOAD RINSE, 0-DI, 1-CHEM= 0
  87.	LEFT CLEAN RINSE, 0-DI, 1-CHEM= 0
  88.	RIGHT CLEAN RINSE, 0-DI, 1-CHEM= 0
  89.	CURTAIN RINSE, 0-DI, 1-CHEM= 0
  90.	CURTAIN RINSE (1-ENBL, 0-DSBL)= 0
  91.	CURTAIN RINSING TIME (SECONDS)= 0      
  92.	FORCE 0-LBS, WAFER DIA(mm)-PSI= 0
  93.	TEMP UNITS (O-degC, 1-degF)= 0
  94.	SHOW COND FORCE (0-NO,1-YES)= 1
  95.	SHOW PUMP 3 (0-NO, 1-YES)= 1
  96.	SHOW pH (0-NO, 1-YES)= 1
  97.	TBL RNS W/CUR RNS (0-NO, 1-YES)= 0
  98.   NOT USED
  99.   NOT USED
  100. 	SPINDLE FLUSH (0-NO, 1-YES)= 1
  101.	FOUNTAIN RINSE (0-NO,1-YES)= 1
  102.	ROBOT LEADSCREW VALUE= 8000
  103.	SEND 1 ROBOT POSITION= -35.88
  104.	SEND 2 ROBOT POSITION= -24.99
  105.	REC'V 1 ROBOT POSITION= -35.82
  106.	REC'V 2 ROBOT POSITION= -24.95
  107.	LEFT LOAD ROBOT POSITION= -9.7
  108.	RIGHT LOAD ROBOT POSITION= -.36
  109.  NOT USED
  110.	ROBOT ARMS POSITION SEND 1= -3000
  111.	ROBOT ARMS POSITION SEND 2= -3100
  112.	ROBOT ARMS POSITION REC'V 1= 2900
  113.	ROBOT ARMS POSITION REC'V 2= 2980
  114.	ROBOT ARMS POSITION LEFT LOAD= 3000
  115.	ROBOT TOP ARM POSITION RT LD= 4975
  116.	ROBOT Z TOP POSITION= .17
  117.	ROBOT Z MIDDLE POSITION= .255
  118.	ROBOT Z BOTTOM POSITION= .34
  119.  NOT USED
  120.	ROBOT BOTTOM ARM POS'N RT LD= -3000
  121.  NOT USED
  122.  NOT USED
  123.  NOT USED
  124.  PROG CONDITION (0-NO,1-YES)= 1
  125.  PROG COND PARK POSITION= -2500
  126.  PROG COND CENTER POSITION= 109000
  127.  PROG COND EDGE POSITION= 140000
  128.  PROG BRUSH REDUCER RATION= 15
  129.  NOT USED
  130.	ROBOT Z OVER LOAD POSITION= .24
  131.	ROBOT Z LEFT LOAD POSITION= .415
  132.	ROBOT Z RIGHT LOAD POSITION= .415
  133.  NOT USED
  134.  SWEEP START (0-EDGE, 1-CENTER)= 1
  135.	SWEEP STYLE (0-1 WAY, 1-2 WAY)= 0
  136.	IN SITU COND (0-DSBL, 1-ENBL)= 0
  137.  NOT USED
  138.  NOT USED
  139.  NOT USED
  140.	CONTROLLED SWEEP (0-NO, 1-YES)= 1
  141.	COND SCALAR= 0
  142.	COND PERIOD WIDTH= 0
  143.	COND HORIZONTAL SHIFT= 0
  144.	COND MINIMUM COMMAND= 0
  145.	COND VERTICAL OFFSET= 0
  146.  NOT USED
  147.  SEAT WAFERS @ UNLD (0-NO, 1-YES)= 0
  148.  VAC W/FOUNT RINSE (0-NO, 1-YES)= 0
  149.  WET IDLE TABLE SPEED= 10
  150.	AUTO WET IDLE (0-NO, 1-YES)= 1
  151.	WET IDLE TABLE ON (secs)= 30
  152.	WET IDLE TABLE OFF (secs)= 60
  153.	WET IDLE CARRIERS ON (secs)= 5
  154.	WET IDLE CARRIERS OFF (secs)= 60
  155.	CARRIER PAD RINSE TIME (secs)= 5
  156.	LOAD RINSE TIME (secs)= 5
  157.	FOUNTAIN RINSE TIME (secs)= 5 
  158.  NOT USED
  159.  NOT USED
  160.	2nd TABLE (0-NO, 1-YES)= 0
  161.  NOT USED
  162.	2ND TABLE REDUCER RATIO= 0
  163.  NOT USED
  164.	SCRUBBER PRESENT (0-NO, 1-YES)= 0
  165.	USE SCRUBBER (0-NO, 1-YES)= 0
  166.	SCRUBBER LOAD POSITION= 0
  167.	SCRUBBER UNLOAD POSITION= 0
  168.  NOT USED
  169.	WAFER BLOWOFF TIME (secs)= 15
  170.	APPROX WAFR BLOWOFF (psi)= 35	
  171.	MIN LIFT OFF VACUM (psia)= 8
  172.	WAFER PRESENT ERROR VAC (psia)= 8
  173.	APPROX FLUSH BLOWOUT PRES (psi)= 12.5
  174.	WAFER ABSENT ERROR VAC (psia)= 8.5
  175.	COOL WATER DETECT (GPM, 0-DSBL)= 0
  176.	TBL START W/SLRY (RPM, 0-DSBL)= 10 
  177.	ROTATE RESTART (0-NO, 1-YES)= 1
  178.	LOAD RINSE (0-INTRVL, 1-STEADY)= 0
  179.	ELEV TANK FILL TIME (sec)= 0
  180.	DUNK SEND (1-EN,0-DIS) RESET= 0
  181.	FILTER ROUTINE Kf= 10
  182.	FILTER ROUTINE Kl= .1
  183.	SERVICE ALARM (0-DSBL, 1-ENBL)= 0
  184.	CASS RANDOM/MATCH (0,1) RESET= 0
  185.	IDLE DELAY MINUTES (5 MINIMUM)= 0
  186.	CONSERVE LOAD RNS (0-NO, 1-YES)= 0
  187.  AIR WAND 90-OFF, 1-POL, 2-COND)= 0
  188.  NOT USED
  189.  NOT USED
  190.	SECS BAUD RATE= 9600
  191.	SECS DEVICE ID= 1
  192.	SECS T1 TIMEOUT= 5
  193.	SECS T2 TIMEOUT= 5
  194.	SECS T3 TIMEOUT= 15
  195.	SECS T4 TIMEOUT= 15
  196.	SECS RETRY LIMIT= 3
  197.	SECS COMM PORT (1 OR 2)= 1
  198.	REPORT INTERVAL (SECONDS)= 10
  199.	REPORT MODE (1-ENBL, 0-DSBL)= 1
  200.	PASSWORD (4 DIGIT NUMBER)= 



7.41SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1123
ASDG::POIRIER                                        20 lines  26-FEB-1995 22:35
                           -< Rotory Joint Problems >-
--------------------------------------------------------------------------------
    
    Problem: New Right Spindle Rotory Joint Is Wobbling Badly.
    
    
    
    History: The new Joint is higher that the old one. This made it
             necessary to fabricate new brackets to prevent the entire
             assembly from rotating. DONE
    
             The vacuum tube hole that is drilled through the hub mount
             is off center. This is causing the tube to RUN OUT and the 
             Rotory Joint, which is 28" higher, to wobble.  The old joint
             was a much different design and didn't seem to be affected
             by this condition. Therefore, I'm recommending that we re-
             install the old joint until we can work out the problems
             with the new one. I've asked Brian to get started on 3rd
             shift tonight.
    
    
                                                            John.
7.42SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1229
SUBPAC::LANDRY                                       26 lines  28-FEB-1995 16:49
                            -< load cells checked >-
--------------------------------------------------------------------------------
    Today..........
    
    Eric and I went over a couple of items concerning the spindles:
    
    	*  We checked the spindle dead idle weight.  By this we mean that
           the locking pin was removed from  the spindle assembly where it
    	   mates with the air cylinder, hence at an idle state.  Using the
    	   diagnostics screen, the reading for the spindles should be at
           2048 +/- 5.  The right spindle was fine.  The left spindle was
    	   at 2060.  So, we adjusted the load cell "zero" potentiometer to
    	   read 2048.
    
    	*  We then checked the display monitor for the force readings. 
           They still were not matching.  We tried swapping the load cells
           for the spindles to see if the readings went with the load cells
           but they didn't.  This tells us the load cells are fine.
    
    	*  Since one load cell was tweeked, we then performed down force
           calibration.  As expected, the left spindle reading was out (due
           to it being tweeked) and the right spindle was fine (due to no
           load cell tweeking).
    
    	*  The system will now cycle wafers to see if there is any improve-
           ment or degregation.
    
    						Doc  /  Eric
7.43SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1219
ASDG::POIRIER                                        16 lines   1-MAR-1995 00:17
                      -< "Spindles not Lifting" problem >-
--------------------------------------------------------------------------------
    
    * Checked the eccentric cam adjustment on both spindles. Both
      adjustments were correct. (checking for excessive friction)
    
    * Checked the span adjustment on the Fairchild E-P regulators.
    
    * I spoke with Strasbaugh field service about the (surface tension/
      spindles not lifting) problem. They suggested that we make sure the
      vacuum is off until the end of step #3 in the auto polish cycle.
      They also suggested we check the spindle lever arm linkage for
      excess friction and make sure the spindle air cylnder needle valves
      are open. 
    
    
    
                                                               John.
7.44SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1327
ASDG::POIRIER                                        24 lines   3-MAR-1995 23:56
                      -< Spindle lift-off problem solved >-
--------------------------------------------------------------------------------
    
     * PROBLEM:   Spindles not lifting carriers off polishing table.
    
    
    ROOT CAUSE:   Conditioning disk was worn out. It appears that the pad
                  was not being completely conditioned. This caused excess
                  surface tension at the wafer/pad interface and prevented
                  the spindles from lifting up.  
    
    
      SOLUTION:   Eric Bodensieck replaced the conditioning disk with a 
                  new one. We left the old pad on the table, conditioned
                  it with the new disk and ran 25 monitor wafers. There
                  were no fly-outs or lifting problems during the run.
    
    
    
        
             *   All pivot points on both spindle assemblies were
                 lubricated. 
    
    
    
                                                           John & Bob.
7.45SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1226
ASDG::POIRIER                                        23 lines   7-MAR-1995 18:01
           -< Back Pressure E-P Regulator/Conditioner Air Cylinder >-
--------------------------------------------------------------------------------
    
    
    * The Spindle Down Force E-P Regulator Calibration was performed today.
      
    
    * The Back Pressure Calibration procedure was also completed today.
      It was discovered during the calibration procedure that the Right 
      Spindle E-P Back Pressure Regulator was not repeating. Strasbaugh
      was called and agreed to a warranty replacement. They shipped a 
      replacement today via UPS Red. We should have it tomorrow.
    
    
    * Conditoner: The conditioning disk was not completely retracting into
                  the Park Station Resivoir. This is the same problem that
                  led to the Ball Spline Assembly repacement 2 months ago.
                  The Ball Spline Assembly checked out O.K. The conditioner
                  down force(air cylinder) turned out to be the cause this
                  time. Strasbaugh was called and agreed to a warranty
                  replacement. The shipped it today along with the
                  regulator.
    
                                                       John, Tim & Jim.
                                                                       
7.46SPINDLE REGULATOR INSTALLEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1331
SUBPAC::LANDRY                                       28 lines   8-MAR-1995 17:53
                  -< A busy afternoon!  New parts installed. >-
--------------------------------------------------------------------------------
    Todays activities.............
    
    *  The new regulator for the right spindle back pressure arrived late
       this afternoon.  It has been installed.  Tomorrow morning we will
       calibrate it per the new procedure being tested.  Please leave it
       as it is.
    
    *  The new air cylinder for the conditioning arm's verticle disk move-
       ment arrived late today as well.  It was installed and tested for 
       proper movement.  All is working fine.  This will also be calibrated
       in the morning per the new procedure just developed.  Please leave
       it as it is.
    
    *  The system had many problems with the computer today.  The problems
       centered around system re-powerup.  Kept getting the infamous "bad
       command" right after the initial setup checks.  Took several retries
       each time re-powering.  Computer connections / board connections 
       may possibly be the culprit.  Tried repositioning the cables in the
       back of the computer and booted up fine the first time.  Future work
       may be necessary to improve connection situation.
    
    *  The top slurry pump hose has a minor leak at this time.  The hose
       itself may need replacing.  If the time permits tonight, feel free
       to acquire the proper hosing and replace.  Be a good idea to change
       out the bottom pumps hose as well.
    
    						JP / Jim / Bobby / Doc
                                         
7.47SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1318
ASDG::POIRIER                                        15 lines  14-MAR-1995 20:29
              -< Left Load Station / Load ring out of tolerance. >-
--------------------------------------------------------------------------------
    
    
    * The load stations were realigned on 1st shift today because of the
      misloading errors we've been experiencing lately.
    
      Jim Tooker identified a problem with the Load Ring in the Left Load
      Station. The I.D. of the Left Load Ring is .005" less than the Right
      Load Ring. This is causing a .001" interference fit between the Ring
      and the Carrier. The result of this condition is consistantly mis-
      loaded wafers on the Left Load Station. Strasbaugh was contacted and
      has agreed to ship us a new Load Ring for the Left Load Station via
      UPS RED. We expect it to arrive tomorrow afternoon.
       
                                                    John, Jim, Doc & Roger.
                                                                           
7.48SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1333
ASDG::POIRIER                                        30 lines  15-MAR-1995 20:33
              -< Load Ring replacement/Conditioner Disk Assembly >-
--------------------------------------------------------------------------------
    
    * Left Load Station:
    
      The (Load Ring) on the Left Load Station was replaced on first shift
      today. The Station was realigned after the ring was replaced and a 50
      wafer test run was completed. 1 load failure was observed during the
      50 wafer run. We'll repeat the alignment procedure tomorrow and look
      for improvement. 
    
    
    
    * Conditioner Head Assembly:
    
      The conditioner was partially disassembled due to the reoccuring
      problem with the "disk not completely retracting into the park
      station resevoir." The Ball Spline Assembly in the Conditioning
      Head was replaced a few months ago because of this problem. We
      knew the Ball Spline was defective because the ball bearings were
      falling out of it's linear bearings. Last week the Conditioner Air
      Cylinder was replaced because it's seals were swelled and it had
      excessive friction. Now the 2 lb disk retracts only 50% of the time.
      Since the system is now free of excessive friction, we're of the
      opinion that the ball bearing in the ball spline don't move freely
      on every cycle. When they do, the disk retracts and when they don't,
      the disk retracts only partially. We called Strasbaugh and they
      suggest we bias the air pressure on the Conditioner E-P reguator
      by 1 or 2 p.s.i. to ensure that the disk retracts. 
    
    
                                               John, Doc, Jim & Roger.
7.49SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1314
PATE::GEE                                            11 lines  22-MAR-1995 23:48
                             -< Conditioning arm >-
--------------------------------------------------------------------------------
         * When the conditioning arm homed, the pad conditioner was resting
           on the side of the clean station bowl. 
       
              - Repositioned the conditioned arm home position Reed 
                switch 1/16" to the right.
   		
              - Ran some homing tests and the pad conditioning program.

         Back up for production.
                
				Bob
7.50SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1411
SUBPAC::BJUBINVILLE "Brooks, Tritt and Diffie"        8 lines   3-APR-1995 08:02
                              -< roller replaced >-
--------------------------------------------------------------------------------
    
        Replaced the new style rollers with the old style.  The
     new style caused the retaining ring to scrape up everywhere. 
     the worse ring is on public display in the chase for your veiwing
     pleasure.  John will be contacting Strasbaugh regarding this
     problem.  Longer shoulder screws would be the best solution...
    
                                                   BJ
7.51SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1317
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              14 lines   4-APR-1995 06:27
                 -< roller problems, actuator being replaced >-
--------------------------------------------------------------------------------
    
          Problems with the rollers were focused on the flow controllers
    for the load chucks.  The load chucks were moving to fast and hitting
    the end of travel.  Using solenoid #4 and #6, they were turned down.
    The Load Station Air Cylinders didn't seem to go down to hard or fast.  
    In fact, they they came down rather soft so I didn't adjust them at
    all.  The value for #38, SOFT DOWN was changed to 1100 lbs.
    
        Jim had also wanted me to swap out the conditioner arm
    actuator with a new style that was in his office.  Still 
    working on it...
    
                                                        BJ
    
7.52SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1451
ASDG::POIRIER                                        48 lines   4-APR-1995 19:53
                -< Load Chuck/Soft Down/Hard Down adjustments >-
--------------------------------------------------------------------------------
    
    * Conditioner Assembly Actuator replacement:
    
      Brian Jubinville replaced the actuator on 3rd shift last night. He
      asked me to double check the edge, center & park positions. They
      all checked-out fine.
    
    
    
    * Load Chuck adjustments:
      
      I tweaked the Load Chuck "travel speed" and "supply pressure"
      this morning per Strasbaughs instructions. The jarring motion in
      the Spindles is now gone. It appears that the supply pressure was
      indeed too high as Strasbaugh suspected. This excess pressure was
      over-powering the Spindles during the load sequence.
    
    
    
      Spindles "down soft" & "down hard" calibration variables:
    
      I reduced the down soft value to 1000 and the hard down to 1200.
      The carriers no longer put excessive force on the Cleaning Station
      Rollers during the cleaning sequence or during the Wet Idle mode.
    
      FYI: Annie has placed an order for replacement rollers and has added
           them to the Fab 6 stockroom BOM for the Strasbaughs.
    
    
     
    * Load Rings:
    
      Annnie has placed an order with Strasbaugh for 2 additional Load
      Rings. This will allow us to cycle them out for replating 2 at a
      time.
    
    
    * Conditioning Disk:
    
      It appeared that the carriers were experiencing the same excessive
      suction condition that was discovered during the last rash of 
      wafer breakage. This was detected by visual and audible observation
      by Bob Gee and myself. Bob and I replaced the conditioning disk as a
      precaution. It has been 32 days since the last disk replacement. We
      will enter this info in the EST notes file for tracking purposes. 
    
    
                                                             John & Bob.
7.53SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1424
SUBPAC::LANDRY                                       21 lines  13-APR-1995 19:18
                         -< Slurry issues of the day >-
--------------------------------------------------------------------------------
    The system experienced a couple of issues today..........
    
    *  The slurry loop kept kicking off every time ANY VMB box was touched.
       Ashland Chemical attributed this to the frozen ball valve in the
       last slurry box.  They came up an installed a new one.  
    
    *  Next, we experienced problems with slurry getting past the pumps and
       eventually dripping onto the pad.  This was attributed to the new
       ball valve installed an hour earlier.  We were having a hard time 
       maintaining a steady pressure.  As the pressure rose, the leak would
       increase.  It took awhile, but we got it under control.
    
    *  Due to the pressure change, the slurry pumps were calibrated.  They
       were way off as expected....but much more than expected.  We
       therefore checked out the in-line filters.  You could see through
       them that they were "jell-in up".  Changed them out and reran the
       calibration.  Made the adjustment and turned the system back over
       to manufacturing.  (Kudo's to Earl for his help in the matter).
    
    						Doc  /  JP
    
7.54SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1512
ASDG::POIRIER                                         9 lines  17-APR-1995 23:37
                   -< DC Servo Module for Axis #4 Replaced. >-
--------------------------------------------------------------------------------
    
    * Replaced the DC Servo Module for Axis #4 on the DCX Motion Control
      Board. This was the Module that the rep brought on Friday. We
      couldn't get access to the tool on Friday do to the amount of
      work in the area. We're not planning to run a verification test
      on A2 since the results of the 500 wafer run on A1 were good.
      (The DC Servo Module was replaced on A1 early last week.)
    
                                                               John.
7.55SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1421
ASDG::POIRIER                                        18 lines  24-APR-1995 21:23
              -< Send Elevator problems/Cleaning Station rollers >-
--------------------------------------------------------------------------------
    
    * The "Right Send" elevator is experiencing problems transfering 
      wafers to the robot. It appears that the Z-height is incorrect
      even after the axis is rehomed. Because this repair will require
      removal of the Elevator assembly (a time consuming task), we
      decided to put the system back UP with a WARNING on the Elevator.
      This allowed Manufacturing to perform MQC's and prepare the system
      to run product tonight. The system can still function using the 
      Left Send Elevator. Please do not use the Right Send Elevator until
      futher notice.
    
     * New rollers have been installed on both Cleaning Stations. These
       rollers spin on the shoulder screws unlike the old ones that were
       dragging. We expect the new rollers to remain functional as long
       as the SPINDLES SOFT DOWN VALUE calibration variable remains under 1200.
      
                                                                              
                                                                   John.
7.56SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1550
ASDG::POIRIER                                        47 lines  25-APR-1995 19:56
                        -< Particle problem activities >-
--------------------------------------------------------------------------------
    
    * Particle problem activity:
    
      The system was experiencing particles in the 20,000 range this 
      morning. A thorough cleaning was performed and MQC run. The 
      particles came down somewhat but were still over 10,000. A
      second cleaning was performed and a "pad break-in" run was
      completed. Another particle check indicated particles still
      in the 10,000 range.
    
    
      The following plan of action was taken this afternoon:
    
      1) Clean oxide monitors (<50 particles) were used. They were
         premeasured and run thru scrubber A2 and remeasured on the
         surfscan. They came within spec (<50 adders). CLEAN SCRUBBER
    
      2) The system was cleaned once again with extra attention paid to 
         polishing pad. The pad was thoroughly rinsed with D.I. water
         for 15 min using the spray gun.
    
      3) Eight wafers were run thru CMP.A2. Four were run thru the Load
         Station & Cleaning Station with NO touch down on the pad. The
         remaining four wafers ran thru the same process AND a 10 sec
         touch down on the polishing pad.  
     
         
      4) The four wafers that touched down on the pad averaged 650 adders. 
         The spec is (<800).
    
         The four wafers that did NOT touch down on the pad averaged 175
         adders.
    
       CONCLUSION: It appears that the thorough cleaning of the pad made
                   a significant difference in the amout of particles added.
                   Also, there is a big difference in the amount of
                   particles added between the wafers that touched down
                   on the pad and the wafers that didn't.
    
                   This may be due to the fact that we're presently
                   adapting Speedfam pads to run on the Strasbaughs.
                   For example, cutting off the outer 2" of the pad
                   & plugging the center hole. (exposed fibers &
                   exposed PSA adhesive). Just a theory, no data to verify.
          
    
                                                     John, Jim S. &  Jesus.
7.57SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:148
PATE::GEE                                             5 lines  28-APR-1995 21:13
                    -< Right send elevator would not home >-
--------------------------------------------------------------------------------
    	The Right send elevator wouldn't home, it would go up looking for
        home instead of down. Powered down the system and re-homed, it's
    	working good now. I removed the Warning in RTC.
    
    Bob
7.58SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1516
FABSIX::R_POIRIER                                    13 lines   1-MAY-1995 23:46
                              -< Spindle problem >-
--------------------------------------------------------------------------------
 

Had a problem with Left Spindle. Error message
said Left Spindle Vacuum not found. Unloaded left 
carrier and cleaned O-rings and carrier. Problem
still the same. Swapped left and right carriers
and problem went away. Put carriers back on the
original spindles and both carriers worked fine.


Jesus Roa running MQC's on CMP.A2
			
		2nd Shift
7.59SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1510
PATE::GEE                                             7 lines  16-MAY-1995 23:58
                           -< Robot Travel Problems >-
--------------------------------------------------------------------------------
	Problem: Robot would not arrive at left send position.

	Fix: Removed and re-aligned the bellows. Cleaned and lubricated 
             the rails, lead screw and robot bearings. Re-aligned the 
	     slide rails. We removed the Warning from RTC.
 
						2nd Shift
7.60ROBOT TRAVEL PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1612
SUBPAC::LANDRY                                        9 lines  17-MAY-1995 18:58
                                 -< wired!!! >-
--------------------------------------------------------------------------------
    This morning...........
    
    The robot Z movement was not completing it's full travel.  The movement
    was also very jerky.  Took the front cover of robot assembly off. 
    Found some wires/connectors in a stressed situation.  Relievd the
    stress and tested results.  All was working fine.  I tried to place the
    wires/connectors in a place that would be more user friendly.
    
    						Doc 
7.61SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:16214
PATE::GEE                                           211 lines  17-MAY-1995 23:53
                          -< Updated Cal. Variables >-
--------------------------------------------------------------------------------

        CALIBRATION VARIABLES
	
	Change : #103 before -35.88 after -35.85
	----------------------------------------


  1.    BRIDGE HOME POS.= -33.86       
  2.    BRIDGE LOAD POS.= -33.53       
  3.    BRIDGE POLISH POS.= 0          
  4.    BRIDGE MIDDLE POS.= -11        
  5.    TEMP IN SCALE FACTOR= 1        
  6.    TEMP OUT SCALE= 1
  7.    TEMP IN OFFSET= -.8
  8.    TEMP OUT OFFSET= 1.5
  9.    BRUSH SPEED SCALE FACTOR= .719 
  10.   SLURRY 1 FLOW FACTOR= 1.08
  11.   SLURRY 2 FLOW FACTOR= 1.29
  12.   TABLE SHUNT LOE TEMP= 0
  13.   TABLE SHUNT HIGH TEMP= 0
  14.   BRIDGE LEFT CURTAIN POS.= -12
  15.   BRIDGE RIGHT CURTAIN POS.= -25
  16.   NOT USED
  17.   NOT USED
  18.   NOT USED
  19.   NOT USED
  20.   SLURRY 2 FLOW FACTOR= .96
  21.   NOT USED
  22.   NOT USED
  23.   NOT USED
  24.   NOT USED
  25.   NOT USED
  26.   NOT USED
  27.   NOT USED
  28.   NOT USED
  29.   NOT USED
  30.   LEFT PRESS. OUTPUT FACTOR= 6.1
  31.   LEFT PRESS. OUTPUT OFFSET= 975
  32.   LEFT PRESS. MIM.= 800
  33.   LEFT FORCE SCALE FACTOR= .94
  34.   LEFT SPINDLE DEAD WEIGHT= 144.79
  35.   LEFT SPINDLE REDUCE RATIO= 10
  36.   LEFT SPINDLE CURRENT OFFSET= 0
  37.   LEFT PRESSURE MINIMUM, TBL2= 0
  38.   SPINDLES DOWN SOFT VALUE= 1200
  39.   SPINDLES DOWN HARD VALUE= 1100
  40.   RIGHT PRESS. OUTPUT FACTOR= 6.1
  41.   RIGHT PRESS. OUTPUT OFFSET= 975
  42.   RIGHT PRESS. MIM.= 800
  43.   RIGHT FORCE SCALE FACTOR= .946
  44.   RIGHT SPINDLE DEAD WEIGHT= 142.46
  45.   RIGHT SPINDLE REDUCE RATIO= 10
  46.   RIGHT SPINDLE CURRENT OFFSET= .1
  47.   RIGHT PRESSURE MINIMUM= 0
  48.   SPINDLES DOWN EXTRA SOFT VALUE= 0
  49.   NOT USED
  50.   PRESS. CONTROL LOOP GAIN= .75
  51.   MAX. PRESSURE STEP DIVISOR= 30
  52.   FORCE LOOP DERIVATIVE GAIN= 56
  53.   FORCE LOOP NEG DER GAIN= 0
  54.   DOWN FORCE INTEGRAL GAIN= 0
  55.   LEFT FORCE SCALING= 1
  56.   RIGHT FORCE SCALING= 1
  57.   BACK PRESS. CONTROL GAIN= 4
  58.   NOT USED
  59.   NOT USED
  60.   6DS REPEAT TEST (0-NO, 1-YES)= 0
  61.	TABLE TEMP CONTROL GAIN= 35
  62.	TABLE REDUCER RATIO= 7.5
  63.	TABLE CURRENT OFFSET= .9
  64.	FILTER CONST.(0-UP/FAST-SLOW)= 1
  65.	BACKPRES UNITS(0-PSIA, 1-PSIG)= 0
  66.	LT BACKPRESS OFFSET= -.5
  67.	LT BACKPRESS SCALING= 1
  68.	RT BACKPRESS OFFSET= -.9
  69.	RT BACKPRESS SCALING= 1
  70.	COND FORCE LOOP GAIN= 3
  71. 	COND FORCE DEAD WEIGHT (LBS)= 2.436
  72.	COND FORCE SCALE FACTOR= .8
  73.	COND FORCE 0-LBS, DIA (mm)-PSI= 0
  74.   NOT USED
  75.	PUMP 3 FLOW FACTOR= 0
  76.   NOT USED
  77.   NOT USED
  78.   NOT USED
  79.   NOT USED
  80.	AXIS 5 OFFSET TO SLOT 0= 3.08
  81.	AXIS 6 OFFSET TO SLOT 0= 3.1
  82.	AXIS 7 OFFSET TO SLOT 0= 3
  83.	AXIS 8 OFFSET TO SLOT 0= 2.93
  84.	CASSETTE SLOT SPACING= .25
  85.	LEFT LOAD RINSE, 0-DI, 1-CHEM= 0
  86.	RIGHT LOAD RINSE, 0-DI, 1-CHEM= 0
  87.	LEFT CLEAN RINSE, 0-DI, 1-CHEM= 0
  88.	RIGHT CLEAN RINSE, 0-DI, 1-CHEM= 0
  89.	CURTAIN RINSE, 0-DI, 1-CHEM= 0
  90.	CURTAIN RINSE (1-ENBL, 0-DSBL)= 0
  91.	CURTAIN RINSING TIME (SECONDS)= 0      
  92.	FORCE 0-LBS, WAFER DIA(mm)-PSI= 0
  93.	TEMP UNITS (O-degC, 1-degF)= 0
  94.	SHOW COND FORCE (0-NO,1-YES)= 1
  95.	SHOW PUMP 3 (0-NO, 1-YES)= 1
  96.	SHOW pH (0-NO, 1-YES)= 1
  97.	TBL RNS W/CUR RNS (0-NO, 1-YES)= 0
  98.   NOT USED
  99.   NOT USED
  100. 	SPINDLE FLUSH (0-NO, 1-YES)= 1
  101.	FOUNTAIN RINSE (0-NO,1-YES)= 1
  102.	ROBOT LEADSCREW VALUE= 8000
  103.	SEND 1 ROBOT POSITION= -35.85
  104.	SEND 2 ROBOT POSITION= -24.99
  105.	REC'V 1 ROBOT POSITION= -35.82
  106.	REC'V 2 ROBOT POSITION= -24.95
  107.	LEFT LOAD ROBOT POSITION= -9.7
  108.	RIGHT LOAD ROBOT POSITION= -.36
  109.  NOT USED
  110.	ROBOT ARMS POSITION SEND 1= -3000
  111.	ROBOT ARMS POSITION SEND 2= -3100
  112.	ROBOT ARMS POSITION REC'V 1= 2900
  113.	ROBOT ARMS POSITION REC'V 2= 2980
  114.	ROBOT ARMS POSITION LEFT LOAD= 3000
  115.	ROBOT TOP ARM POSITION RT LD= 4975
  116.	ROBOT Z TOP POSITION= .17
  117.	ROBOT Z MIDDLE POSITION= .255
  118.	ROBOT Z BOTTOM POSITION= .34
  119.  NOT USED
  120.	ROBOT BOTTOM ARM POS'N RT LD= -3000
  121.  NOT USED
  122.  NOT USED
  123.  NOT USED
  124.  PROG CONDITION (0-NO,1-YES)= 1
  125.  PROG COND PARK POSITION= -2500
  126.  PROG COND CENTER POSITION= 109000
  127.  PROG COND EDGE POSITION= 140000
  128.  PROG BRUSH REDUCER RATION= 15
  129.  NOT USED
  130.	ROBOT Z OVER LOAD POSITION= .24
  131.	ROBOT Z LEFT LOAD POSITION= .415
  132.	ROBOT Z RIGHT LOAD POSITION= .415
  133.  NOT USED
  134.  SWEEP START (0-EDGE, 1-CENTER)= 1
  135.	SWEEP STYLE (0-1 WAY, 1-2 WAY)= 0
  136.	IN SITU COND (0-DSBL, 1-ENBL)= 0
  137.  NOT USED
  138.  NOT USED
  139.  NOT USED
  140.	CONTROLLED SWEEP (0-NO, 1-YES)= 1
  141.	COND SCALAR= 0
  142.	COND PERIOD WIDTH= 0
  143.	COND HORIZONTAL SHIFT= 0
  144.	COND MINIMUM COMMAND= 0
  145.	COND VERTICAL OFFSET= 0
  146.  NOT USED
  147.  SEAT WAFERS @ UNLD (0-NO, 1-YES)= 0
  148.  VAC W/FOUNT RINSE (0-NO, 1-YES)= 0
  149.  WET IDLE TABLE SPEED= 10
  150.	AUTO WET IDLE (0-NO, 1-YES)= 1
  151.	WET IDLE TABLE ON (secs)= 30
  152.	WET IDLE TABLE OFF (secs)= 60
  153.	WET IDLE CARRIERS ON (secs)= 5
  154.	WET IDLE CARRIERS OFF (secs)= 60
  155.	CARRIER PAD RINSE TIME (secs)= 5
  156.	LOAD RINSE TIME (secs)= 5
  157.	FOUNTAIN RINSE TIME (secs)= 5 
  158.  NOT USED
  159.  NOT USED
  160.	2nd TABLE (0-NO, 1-YES)= 0
  161.  NOT USED
  162.	2ND TABLE REDUCER RATIO= 0
  163.  NOT USED
  164.	SCRUBBER PRESENT (0-NO, 1-YES)= 0
  165.	USE SCRUBBER (0-NO, 1-YES)= 0
  166.	SCRUBBER LOAD POSITION= 0
  167.	SCRUBBER UNLOAD POSITION= 0
  168.  NOT USED
  169.	WAFER BLOWOFF TIME (secs)= 15
  170.	APPROX WAFR BLOWOFF (psi)= 35	
  171.	MIN LIFT OFF VACUM (psia)= 8
  172.	WAFER PRESENT ERROR VAC (psia)= 8
  173.	APPROX FLUSH BLOWOUT PRES (psi)= 12.5
  174.	WAFER ABSENT ERROR VAC (psia)= 8.5
  175.	COOL WATER DETECT (GPM, 0-DSBL)= 0
  176.	TBL START W/SLRY (RPM, 0-DSBL)= 10 
  177.	ROTATE RESTART (0-NO, 1-YES)= 1
  178.	LOAD RINSE (0-INTRVL, 1-STEADY)= 0
  179.	ELEV TANK FILL TIME (sec)= 0
  180.	DUNK SEND (1-EN,0-DIS) RESET= 0
  181.	FILTER ROUTINE Kf= 10
  182.	FILTER ROUTINE Kl= .1
  183.	SERVICE ALARM (0-DSBL, 1-ENBL)= 0
  184.	CASS RANDOM/MATCH (0,1) RESET= 0
  185.	IDLE DELAY MINUTES (5 MINIMUM)= 0
  186.	CONSERVE LOAD RNS (0-NO, 1-YES)= 0
  187.  AIR WAND 90-OFF, 1-POL, 2-COND)= 0
  188.  NOT USED
  189.  NOT USED
  190.	SECS BAUD RATE= 9600
  191.	SECS DEVICE ID= 1
  192.	SECS T1 TIMEOUT= 5
  193.	SECS T2 TIMEOUT= 5
  194.	SECS T3 TIMEOUT= 15
  195.	SECS T4 TIMEOUT= 15
  196.	SECS RETRY LIMIT= 3
  197.	SECS COMM PORT (1 OR 2)= 1
  198.	REPORT INTERVAL (SECONDS)= 10
  199.	REPORT MODE (1-ENBL, 0-DSBL)= 1
  200.	PASSWORD (4 DIGIT NUMBER)= 




7.62SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1620
ASDG::POIRIER                                        17 lines  18-MAY-1995 19:27
              -< Air motors / D.I. gauge replacement / D.I. leak >-
--------------------------------------------------------------------------------
    
    
    * Air Motors:
    
      Two replacement Air Motors arrived for the Cleaning Stations.
      The system was not available today due to production needs. 
      2nd shift will replace them if the tool becomes available. 
    
      
    * D.I. pressure gauge:
    
      The gauge was not replaced today for the same reason as above.
      Also, a small D.I. leak was noticed just below gauge diaphram.
    
    
                                                                John.
                                                                     
7.63SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1622
FABSIX::R_POIRIER                                    19 lines  22-MAY-1995 23:58
                         -< Wafer break/Brush motors >-
--------------------------------------------------------------------------------

  There was a wafer break on 1st shift today. We rebuilt
  two carriers, replaced the pad with a Speedfam pad, 
  and cleaned the tool. We have started wafer break-in. 
  MQC's will have to be run after pad break-in. Per a 
  request from 1st shift we also checked out both the 
  left and right send elevators while running break-in 
  wafers. Both seem to be working fine.

  We also replaced the Cleaning brush motors.

  We replaced the DI water pressure gauge but the problem 
  does not appear to be the gauge. Bob suspects the problem 
  is with the gauge guard. A new gauge guard will have to
  be ordered from Facilities.
	

		
				2nd Shift
7.64SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:17151
PATE::GEE                                            37 lines  26-MAY-1995 05:00
                                   -< Robot >-
--------------------------------------------------------------------------------
	Problem: Robot bottom arm not moving to home. 

	Action:  Checked the the top and bottom arm home sencors
		 in diagnostics and they looked good Input 41 and 42.

		* Removed the robot and repaired many water damaged  wires.
                  Reinstalled the robot and tried to home. We found that
                  everything would home but, when going to pickup a wafer
                  the top arm would not move into the cassette.

 	        * Checked all connections again and found a pin socket pushed
                  out. Tried to rehome and found that the bottom arm would
                  not move at all

		* After checking every connection again we decided to 
                  replace the bottom arm encoder. Reinstalled the robot and
                  found the bottom arm still would not move. Swapped the 
                  top and bottom arm encoders everything moved, so we put
                  them back into there proper positions and reinstalled the
                  robot.

	----->  The Robot condition now is:

		It's in the system again the top arm homes, the up/down travel
		homes, but the bottom arm shots out to the right then the left
		and will not home. I suspect that the connector for the bottom
                arm was damaged on the last move.

		Note; We need to replace all these wire connectors with
		      water resistant and stronger connectors as soon as
                      possible if we expect the robot to run. I like the 
                      machine, but these connectors are not made to be 
		      submerged in water or even handled.

 

						Second Shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.226                          CMP.A2                           226 of 296
SUBPAC::LANDRY                                       20 lines  26-MAY-1995 18:07
                  -< Wires/connectors and skeet shooting!!! >-
--------------------------------------------------------------------------------
    This must be National Robot Week!  
    
    *  The robot bottom arm isn't homing properly.  It's doing the skeet
       shooting thing.  I didn't spend alot of time on this system as I
       felt the other two polishers would be an easier time of repairing.
       Only God knows the real truth to that as B1 got partial success and
       A1 has continual failure.
    
       John Barbarito was kind enough to come in and give it a run for it's
       money though.  He did find a couple of problems.  Yup, one was at-
       tributed to a broken wire on a connector and another suspected cul-
       prit was a couple of cold solder joints.  After repairing both of
       the items, we tried it out.  No luck!  The arm was still "skeetin'".
    
       John checked out all the connectors/pins and all seemed ok.  This is
       where we stand.  Thanks to John for his gallant effort in attacking
       the robot problems.
    
    							Doc
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.228                          CMP.A2                           228 of 296
ASDG::POIRIER                                        17 lines  30-MAY-1995 20:48
                      -< Robot RS-232 Connector Problems >-
--------------------------------------------------------------------------------
    
    
    Robot RS-232 Connector Problems:
    
    Although the current Plan Of Action is to upgrade the Robot connectors
    to a watertight/mil-spec type tomorrow, some new RS-232 connectors and
    pins were located tonight. 2nd shift is planning to replace the conn-
    ectors tonight and resume verifiying the cause of the Homing problems
    with the Robot Bottom Arm. 
    
    Note: Please use extra caution when handling the connector that con-
          tains the pnuematic sensor wires. The output transistor in this
          sensor will blow if the (white) output wire contacts the 24v 
          (red) input wire.
    
    
                                                                John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.229                          CMP.A2                           229 of 296
ASDG::POIRIER                                        10 lines  31-MAY-1995 21:23
                      -< Robot Connector Problem update >-
--------------------------------------------------------------------------------
    
    Robot Connector Problem:
    
    The new connectors arrived this morning and were installed on both
    the Robot and the machine harness. 2nd shift will be re-installing
    the Robot and testing the system for functionality. If the tool
    performs normally, a verification test run will be started and
    will continue into 3rd shift. 
    
                                                                John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.230                          CMP.A2                           230 of 296
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              16 lines   1-JUN-1995 10:01
                   -< Robot cables replaced and now working >-
--------------------------------------------------------------------------------
    
    
         Came in to help Roger get the robot in.  The upper and
     lower arms on the robot were instant discus throwers.  We
     pulled the robot back off and traced the problem to the
     encoders and the need to have their respectiove leads swapped.
     The lower arm was now fixed.  The orignal problem from last 
     week was the lower encoder which was now appearing on upper
     robot....  thus...  the encoder was bad.  The encoder was swapped
     out and now it worked fine except... the home sensor wasn't working.
     That was traced to being a broken shield wire.  The sensor was
     replaced and now everything works.  John and the Rep are re-routing
     some cable for more slack for the installation of the covers.
     After that, break-in wafers look like they will be run shortly.
    
                                              BJ  

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.231                          CMP.A2                           231 of 296
ASDG::POIRIER                                         8 lines   1-JUN-1995 20:06
                         -< Connector Problem Update >-
--------------------------------------------------------------------------------
    
    * Robot Update:
    
      The connectors on the machine cables and Robot wires were completed
      last night on 3rd shift. The machine passed the verify test this
      afternoon and is now running MQC's.
    
                                                                 John.
7.65SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1611
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               8 lines   3-JUN-1995 05:55
                       -< Bridge stopped during process >-
--------------------------------------------------------------------------------
         
          During a run both spindles had wafers on them leaving the LOAD
       table but the bridge stopped half way to the polish table.  The
       bridge would not move so it had to be re-homed.  The only thing
       that makes sense is that the back pressure must of gotten just
       out of spec during travel.  This is the first time we've seen this
       happen.
                                                            BJ  
7.66SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1754
ASDG::POIRIER                                        51 lines  13-JUN-1995 20:30
                  -< Cable Carrier Track/Slurry Pump Tubing >-
--------------------------------------------------------------------------------
    
    
    * Robot Z-Axis Bellows:
    
      The Bellows detached itself from the Robot Base this morning.
      The screws evidently backed out and fell off. Replacement
      10 - 32 X 3/4" SS Socket Head screws were used. The old ones
      were slotted round heads 1/2" long. (A real pain to work with).
    
    
    
    * Lead Screw/Rail Assembly:
    
      The Robot X-Axis (Lead Screw/Guide Rail Assembly) was cleaned and
      lubricated while the bellows were off. The motion is greatly 
      improved.
    
    
    
    
    * Cable Carrier Track:
    
      The bottom link of the Cable Carrier Track broke-off this morning.
      This caused the entire track to float around unsecured. A temp-
      orary fix was initiated using 3 tie wraps. It has held-up fine
      thus far, please keep a close watch on it for the next 2 shifts.
     
      The replacement Cable Carrier Mounts (END LINKS) have been order-
      ed from Strasbaugh and added to the Fab6 Stockroom BOM. Also, Cable
      Carrier Links, (REGULAR LINK AND COVER) have been ordered and
      stocked. The order is being FEDEX'ed PRIORITY #1 and will arrive
      tomorrow morning at 9:00 a.m.
    
    
                                                                 
     * Slurry Pump #1 Tubing:
    
       The Tygon Tubing on Slurry Pump #1 blew-off around 11:00 this 
       morning. The tubing was replaced and the flow recalibrated.
    
       NOTE: Don't use the tubing in the top draw of the Vidmar cabinet
       for the Slurry pumps. It has the wrong wall thickness and will not
       work in the Peristalic pumps. (I found out the hard way) 
    
       The correct size is 3/16" I.D. X 7/16" O.D. X 1/8" Wall
       I placed a req for this tubing today. In the mean time, there 
       is a small quantity of this tubing in the draw under the Stainless
       sink.
       
                                                                 John.
                                                                      
7.67SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1730
ASDG::POIRIER                                        27 lines  20-JUN-1995 19:29
              -< Cabel Carrier Track Replaced/Annual P.M. Begun >-
--------------------------------------------------------------------------------
    
    * Cable Carrier Track:
    
      The Cable Carrier Track was replaced and the slack in the cables
      was adjusted. A little extra slack was allowed inside the Robot
      base to ensure easy removal of the connectors. At the other end
      of the track, the cable bundle was fastened to the rear deck 
      with a tie wrap to prevent any contact with the Shuttle. This
      routing is identical to the way the cables were routed at the 
      factory.
    
    
    
    
    * Annual P.M.:
    
      The Annual P.M. was started this afternoon. Roger has removed the 
      pad, rinsed-out the Carriers and placed them in a tray to keep the
      films wet. These Carriers have had only 135 wafers polished on them.
      Lets try to reuse them. Please change the water a few times per shift.
    
                                           
                                                                    John.
    
    
                                                                  
       
7.68SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1711
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               8 lines   8-JUL-1995 07:55
                             -< noise from within >-
--------------------------------------------------------------------------------
    
         A loud metal noise was being heard when the shuttle was moving.
      It ended up being the metal strip that secures the rubber skirt
      on the shuttle access door making contact with the door lock.
      There was several nuts missing thoughout the door area.  All 
      have been replaced and secured.
    
                                                        BJ
7.69SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1711
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               8 lines  15-JUL-1995 06:14
                       -< Bridge stopped during travel >-
--------------------------------------------------------------------------------
    
      The bridge was on the way to the polish table to begin 
    polishing when it stopped half way.  After it was homed,
    it moved fine.  It is suspected that one of the carriers
    lost its back pressure for a second and caused the bridge 
    to halt.
    
                                              BJ
7.70SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:186
PATE::GEE                                             3 lines  17-JUL-1995 07:44
                       -< Robot to left send position. >-
--------------------------------------------------------------------------------
    	Had problems with the robot traveling to the left send position
    	a few times tonight. Cleaned and lubricated the slide rails and
    	lead screw, also cleaned the front and rear robot bearings.
7.71SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:188
PATE::GEE                                             5 lines  25-JUL-1995 08:55
                                -< Water flow  >-
--------------------------------------------------------------------------------
	Both carriers were sticking to the table after the polish cycle
	was completed. Rich swapped the conditioner disk from cmp.a1, but 
	they still were sticking. I checked the water flow during step 3
	and found it set at 4 gpm. Checked cmp.a1 and found that system 
	set at 10 gpm. Adjusted cmp.a2 to 10 gpm and the problem went away.
7.72SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1920
PATE::GEE                                            17 lines  26-JUL-1995 08:52
                           -< Bridge / Robot travel >-
--------------------------------------------------------------------------------
c
    	* The Bridge wouldn't move from the polish table.The error
	  was "Bridge not arriving at load". We couldn't find anything 
          wrong, vacuum was in the fours. Tried to move it manually, it
          would not move. Checked the lubrication on the rails and lead
          screw both looked good, oil level was good also. Powered down
          the system, moved the bridge manually with no problem, very
          smooth. Powered up the system, homed the bridge, ran it back
          and forth, it looked smooth and sounded good. Ran the rest of
	  the night with no problems. Need to contact the field rep. 
	  for a possible cause.
	  
	* Robot travel problems, the robot wouldn't make it to the left 
	  receive position. Cleaned and lubricated rails, lead screw 
	  and robot bearings.

	
7.73SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1818
SUBPAC::LANDRY                                       15 lines   3-AUG-1995 10:28
                         -< conditioner air cylinder  >-
--------------------------------------------------------------------------------
    A day late, but non_the_less data to enter..........
    
    Yesterday, the conditioning arm had a problem where the air cylinder
    would not raise the diamond disk from the dish.  The cylinder was
    definetly getting the air input to raise the cylinder, but mechanically
    it would move.  Disassembled the air cylinder asswmbly.  Found the
    cylinder disconnected from the ball shaft.  It had unscrewed from it's
    normal mounting position.  screwed it back on and re-assembled.  Tested
    fine.
    
    This fix required two people do to space constraints and a poorly
    designed assembly that utilizes a retaining ring.  
    
    							Doc  /  JP
    
7.74SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1921
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              18 lines   4-AUG-1995 07:06
                      -< pad change / robot cal. changed >-
--------------------------------------------------------------------------------
    
        completed pad change as pad count reached 600.  
     However.. some good news and bad news.  The bad
     news is that it failed MQC's.  The good news is 
     that it might just need another boat of wafers.
     That's going on right now and then we'll give the
     MQC's another shot.
    
        As far as the robot is concerned.  It was not
    arriving at the the front receive almost everytime
    the robot was told to move.  I ended up tricking the
    the machine into think that the receive elevator was
    closer than it was.  The value for SEND-1 was -35.83
    and it noe -35.01.  RCV-1 was -35.82 and it now -35.
    We haven't had a hang up with the robot for over 50
    wafers now.
     
                                               BJ
7.75SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1828
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              25 lines   9-AUG-1995 19:51
                              -< Robot repaired >-
--------------------------------------------------------------------------------
    Robot activities continue...........
    
    The robot / rail assembly was removed again at the end of 1st shift.
    The bearings, lead screw, rail heights/widths, and the general assembly
    were all thoroughly checked.  The bearings looked good.  The rails were
    a little pitted.  Used crocas paper to smoothen out best we could. 
    Lubricated the rails and tried moving around by hand turning the lead
    screw.  Could not initially get both ends of the rails to have the same
    torque feeling.  Adjusted a couple of componants, they being the set
    screws on the pillow blocks, the two bolts on the assembly that
    supports the robot, and lastly, the lead screw mounting block
    assemblies on each end.  The robot seemed to be performing much better.
    Re-installed the assembly, bellows, and robot.  Hooked up the
    electronics to the robot and tested results.  The robot repeatably made
    it to it's proper destination EVERY time.  No failures what_so_ever.
    
    JP has begun building carriers as I type.  Also, a pad will need to be
    installed.  Normal recovery procedures from a pad change/carrier change
    need to be performed.
    
    Two more items need to be verified to finish off the PM.  See the
    check sheet for details.  
    
    					JP  /  Doc
    
7.76SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1910
SUBPAC::TOOKER                                        7 lines  12-AUG-1995 10:53
                      -< spindle down force calibrations >-
--------------------------------------------------------------------------------
    Checked left and right spindle Fairchild E-P regulator calibration.  No 
    adjustments necessary.  Re-checked left and right spindle downforce cal-
    ibration.  All displayed and actual downforce readings are well within
    +/- 2.0% of each other as per Strasbaugh spec.  Downforce calibration
    does not appear to be the cause of left/right removal rate differences. 
 
    Jim T.
7.77SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1917
PATE::GEE                                            14 lines  22-AUG-1995 07:27
                 -< 4" Conditioner / variable change / Robot >-
--------------------------------------------------------------------------------
      * We had to replace the pad because of poor removal rates and 
	uniformity problems.	

      * We replaced the conditioner disk assembly with the 4" type
	(no 6" in stock), do to poor performance. 
	There has been a rush order placed on the 6" conditioner
	disks since last week.
    
      * We had to change calibration variable #126 (center position)
 	from 108000 to 104000.

      * Robot end effector was detecting a wafer when no wafer was 
	present. Had to clean out the end effector air blow off
	port.  
7.78SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:1936
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               9 lines  31-AUG-1995 07:07
                          -< Solenoid #23 inspected >-
--------------------------------------------------------------------------------
    
    
         During the pad change, we took apart solenoid #23 to see
     if anything could be done with it.  It had been leaking for
     quite some time judging from the rust stains inside of the 
     valve.  We re-assembled the valve and it still leaked but
     not as much.  A new one is coming in today, tomorrow at the
     latest.
                                                Brian x 2

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.264                          CMP.A2                           264 of 296
ASDG::POIRIER                                        16 lines  30-AUG-1995 19:55
                         -< Solenoid #23 is leaking. >-
--------------------------------------------------------------------------------
    
    * Sol #23 leaking:
    
      Sol #23 has developed a leak which triggered the leak detection
      system. Since the leak is a small steady drip and a replacment
      solenoid will not arrive until Friday morning, we have put the
      system back UP and will continually replace the water soaked
      wipes every few hours. Strasbaugh will try to catch the UPS
      truck tonight. I they do, the solenoid will arrive tomorrow
      afternoon. If they can't, it will arrive at the Fab6 recieving 
      dock Friday morning.
    
      "D" shift is looking into improved methods of containing the water.
      Possibly a small trough to direct the excess water into a container.
    
                                                         John, Roger & Doc.
7.79DI WATER VALVE REPLACEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:2017
SUBPAC::LANDRY                                       14 lines   1-SEP-1995 18:19
                               -< valve......pm >-
--------------------------------------------------------------------------------
    Today, during a pad change, the new DI water valve was installed.
    Quite the task as panels, many DI connections, and wiring were needed
    to be removed.  The valve mounting screws were under the entire DI
    solenoid manifold.  Anyway, the job got done with the assistance of
    Earl D and Barry C's help.  They were quite instrumental!
    
    The defective valve needs to be sent back to Strasbaugh within 30 days
    for financial credit.  They have sent an RMA # for this purpose.  I
    have the number elsewhere at the moment.
    
    The Weekly PM was also completed at this time.
    
    							Doc
    
7.80SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:2026
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              23 lines   7-SEP-1995 06:21
                -< Contactor and monitor replaced, PM finished >-
--------------------------------------------------------------------------------
    
    
        The down force was completed on both spindle as part of
      the monthly PM.  The data can be verified in the PC under
      A2_L906.XLS and A2_R906.XLS.  The remaining items need to
      have the rest of the data retrieved by C shift in the morning.
    
        The monitor and keypad that were shipped in have been re-
     installed on the swivel mount in front.  The front right panel
     latch was re-adjusted so the panel doesn't pinch the wires when 
     the cover in locked into place.  Everything test fine.  The 
     remote monitor is back home in the maintenance chase between 
     CMP.A1 and CMP.A2.
    
        The EMO contactors that were shipped in with the wrong coils
    in them.  We needed a 24V coil and we got 208V coils.  We installed
    a new contactor with the orginal 24V coil from the machine.  The
    contactor section is on the tool box marked N.G.  Inspection clearly
    shows cooked contacts.  The system was EMO'ed numerously without
    any hesitations or delays.       All set.... case closed...
    
                                                D shift
    
7.81SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:2141
ASDG::POIRIER                                        38 lines  16-SEP-1995 19:43
             -< Pack Press Verification / Hub O-rings / Vac Leak >-
--------------------------------------------------------------------------------
                                                                       
    
    * Back pressure verification:
      --------------------------
     
      The Hubs were removed from both Spindles and presence of back pressure 
      was verified at the Feed Tubes on both Spindles.
    
      Conclusion: Back pressure is being applied thru both Spindles to the 
                  Carriers.
      
    
    
    
      Left Spindle Hub O-rings:
      -------------------------
    
      The silicone o-rings in the Left Hub are in bad shape. "B" shift,
      please replace them tonight and mount both Hubs back on the tool.
    
    
    
      Right Spindle Rotory Joint:
      ---------------------------
      
      The Right Spindle Rotory Joint is experiencing a vacuum leak.
      The leak is in the area of the Joint that supplies vacuum to
      the Spindle Quill chamber.(outside the Feed Tube)
    
      You can test the leak by attaching a vacuum gage to the barbred
      fitting protruding from the Hub Mount. Put the tool in manual mode
      and turn the Right Spindle Vacuum ON. The gage will read approx
      20" of vac. (It should be at 24") Now push-up on the Rotory Joint 
      and the vacuum will drop to 12" and the alarm will sound. (The
      press switch is set to alarm at 15")
    
      Since it's possible to drop a Carrier off the Spindle during a Polish
      Cycle, disable the Right Spindle until the Joint can be replaced. 
7.82SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:2037
SUBPAC::LANDRY                                       34 lines  17-SEP-1995 19:52
                 -< rotary union/spindle assembly vacuum leak >-
--------------------------------------------------------------------------------
What a day..........
    
The right spindle was looked at today for a problem with it losing 
vacuum when up pressure is applied to the rotary union.  The carrier
vacuum would only pull down to @20 Hg when the vacuum was applied to
secure the carrier.  When physical pressure was exerted to the rotary 
union, the pressure would drop below the 15 Hg set point, hence alarming
for low vacuum.

We dismantled the spindle assembly.  We found the top o-ring to be 
sheared severely and the bottom o-ring to be in bad shape as well.  So,
both o-rings were replaced and re-assembly of the spindle commenced.
It was then tested out.  Initial vacuum went to @21 Hg (not much improve-
ment) and it would drop to about 12 Hg when up pressure was applied.
So, disd-assembled again to find the top o-ring damaged.  The threads on
the nut were messed up a bit, so a new rotary union was installed, using
the original shaft.  After re-assembly, the initial vacuum was @24" Hg.
As up pressure was performed, the drop was still poor, but better than
before.  Out we come again.  The o-ring was bad again.  Upon closer
inspection of the parts, we found that the shaft was bent enough to cause
the threaded nut to pinch the o-ring as it attempted to seal properly.
With no shafts to be found, we tried to straighten the shaft bend out.
Surprisingly, it came out pretty good, enough to try a re-assembly.
The initial vacuum was now at @25" HG and the drop after up pressure was
minimal (@20" Hg).  After doing up pressure for 5-10 minutes, the leak 
gets bad again.

Looks like we'll have to order a shaft or two and a spare rotary union
on Monday morning.  In the mean time, the B shifters will take a shot 
at it tonight.  Good Luck.

      					the A shifters


7.83SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:219
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               6 lines  13-OCT-1995 05:17
                           -< End effector adjusted >-
--------------------------------------------------------------------------------
    
         The end effector switch was detecting a wafer when
      there was none present.  The switch had to be adjusted.
      It was tested and continued on its way without any problems.
                                            
                                            D shift
7.84SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:2025
ASDG::POIRIER                                        22 lines  19-OCT-1995 19:28
                      -< Wafer Breakage/Shuttle Movement >-
--------------------------------------------------------------------------------
    
    System Performance/ Wafer Breakage:
    
    A2 broke wafers on the Pad during step 2 of the Auto Cycle early
    in the shift. Brad and I built-up 2 more Carriers and the system ran
    for the rest of the shift wihout any wafer breakage.
    
    
    
    
    Shuttle Movement Problem:
    
    The Shuttle failed to move this afternoon. The Shuttle rails and
    Pillow Bock Bushings were cleaned and relubricated. The Shuttle
    translates back and forth now without any problems.
    
    
                                                       
     
    
    
                                                             Brad & John
7.85SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 17 1995 05:2153
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              13 lines  20-OCT-1995 09:04
                               -< table stalled >-
--------------------------------------------------------------------------------
         The pad was replaced earlier in the shift and was broken
      in using the BREAKIN recipe.  As soon as wafers were started
      on the tool for MQC's, the table stalled out during polishing 
      after the amps got over 40.  The system was EMO'ed before 
      anything crashed.  Larry and I monitored the table with 
      different downforces applied the table.  The amps went to 
      40 several more times but it did not stall the table.  At 
      this time, MQC's are being run and the machine is going to be 
      closely watched.
     
                                                    D shift
       
                                         

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.294                          CMP.A2                           294 of 296
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              16 lines  21-OCT-1995 08:12
                             -< Another break.... >-
--------------------------------------------------------------------------------
    
           stall...   crash...  clean-up...
    
           Another stall more breaks.  The Table AMPs went over
      45 and stoppped without any ewrrors.  We lost an exercise
      wafer and a product wafer.  The pad was removed and the 
      machine was cleaned up.  New carreiers and pad have been
      installed.
                             
           The mystery of all of the wafer breaks on A1 and A2
      have been narrowed down to the table stalling out.  The
      gear box is being pushed to its limit and we are now seeing
      the end results.  John and Brad now the proud owners of this 
      issue.
                                                   BJ
                             

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 56.296                          CMP.A2                           296 of 296
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               5 lines   2-NOV-1995 06:07
                             -< Stall speed error >-
--------------------------------------------------------------------------------
         
        The table appeared to have stalled out and alarmed for table
      speed too low.  This error occured with exercise wafers where on
      the tool.  When product was run, it gave the same error.  The 
      computer was softbooted and ran fine for the rest of the monring.
7.94CMP.A2 Status - 11/18/95ASDG::POIRIERSat Nov 18 1995 19:279
    
    
    SyStem Status:
    -------------
    
    * Pad changed & MQC's completed. System ready for Super Hot Lot.
    
    
                                                               John.
7.95broke wafers / replaced carrier platesFABSIX::R_GEESun Nov 19 1995 07:1423
	* Broke product wafers on the pad.

	      System status at the time:

		- 406 wafers on the pad (2 suba IV, 1 IC1000)
		- 203 wafers on the carriers.
		- Down force was 313 lbs
		- Slurry flow set at 500 ml/min
		- 180 sec polish.
		- Back pressure 15.8
		- 10 wafers into a 21 wafer lot.

	   Replaced the pad and carriers. Began running a wafer slid out
	   from under the left carrier twice, the emo was hit and it 
	   didn't crash. Dis-assembled the carrier and replaced the backing
	   plate with the older style threaded standoffs. We encountered
	   the same problem with the right carrier and replaced that one
	   also.   

					B shift

	   

7.96High particles / brush changeFABSIX::R_GEESun Nov 19 1995 07:1519
	* Failed particles and uniformity. 

	 - Ran a particle check using the "Transfer" program, this
	   cycles the wafers thru the system without touching the
	   pad. Particle check showed 1 adder, this shows that the
	   particle problem has to be coming from the polish table 
	   area. 
	 - Removed and cleaned the conditioner.

	 - Cleaned the index table and polish table area.

	 - Ran a second particle check, no improvement.

	* Broke 2 wafers in the clean station.

	 - Found a wafer stuck in the paddle rails. This stopped the 
	   paddle assembly and caused the mis-load at the scrub station.
	  
					B shift
7.97DI Water leakFABSIX::R_GEETue Nov 21 1995 07:1714
    
	* Errored for DI water leak detected.

	   - Found about 1 inch of water in the bottom of the system.

	   - Cleaned up the water, and found valve SOL31 had a small 
	     leak coming from the top. Looked in stockroom 6 with no luck.
	     A shift will need to order one.
	   
	      
					B shift

	   

7.98valve on the waySUBPAC::LANDRYTue Nov 21 1995 18:588
    
    	We contacted Strasbaugh today concerning the DI solenoid valve.
    They are sending one to us for tomorrow morning.  It's shipping by
    Fed-Ex Priority 1.  We'll try and schedule some time with manufac-
    turing to replace it.  
    
    					A shift
    
7.99slurry lines changedSUBPAC::LANDRYWed Nov 22 1995 19:0517
    Today, we had a hose split on the slurry line to pump #1.  The mess was
    cleaned up and the tubing was replaced on both pumps.  Once completed,
    slurry pump flow calibration was performed on both pumps.  Both needed
    new calibration variable numbers as they were flowing on the low side.  
    
    The infamous DI valve didn't make it through shipment today.  Checked
    both docks (morning and afternoon delivery) with no luck.  Therefore,
    D shift will have to make a decision by the end of their shift.  The
    water will either need to be shut off on the system OR a union can be
    put in-line instead of the valve.  A union has been procured and it
    currently resides in the top draw of the tool box.  If the pad life is
    bad/old, maybe best just to shut off the incoming water and put the two
    carriers in a basin of water for the day.  If the pad is good/new, it
    may be best to put in the union.  Do what you think is best!
    
    						Doc  /  JP
    
7.100DI attack!!!SUBPAC::BJUBINVILLENFD Fire - RescueThu Nov 23 1995 06:308
    
    
       The right SEND elevator wasn't repeating.  We pulled
    out the elevator and for several wire corroded and broken.
    We cleaned up the connectors and rewired the broken ones.
    Everything ran fine after that.
     
                                                D shift
7.101Conditioner Arm Actuator "Rod End" ReplacedASDG::POIRIERFri Nov 24 1995 19:4013
    
    Conditoner Arm Failure:
    ----------------------
    
    The Conditioning Arm failed to rise out of the Park Station this 
    afternoon. The Whole Conditioning Assembly had excessive backlash.
    Further inspection revealed a worn-out Rod End on the Actuator.
    The Rod End was replaced and the Actuator Reed Switch was adjusted
    for proper Arm location at the Park Station. The Flow Controllers for
    the Conditioner Up/Down motion were also adjusted to eliminate the
    cluncking on the down stroke. The motion is much smoother now.
    
                                                               John.
7.102low slurry flowSTRATA::FINNMon Nov 27 1995 19:0523
	No slurry flow , Ashland's pump had shut off. Contacted ashland
to restart pump.  Pressure to all systems was still high, but flow through 
system was minimal. Bypassed system using external loop at A2 and slurry flow 
was still low.  Check B1 with same results.  Ashland Chemical changed the input 
and output (return) valve in valve manifold box, hence referred to as VMB.  
Return was clogged solid. Slurry filter was changed, it was clogged.
	
	We turned systems back on with great expectations.  BUT,  same problem 
remained.  Took this opportunity to modify A2 slurry manifold so pump two is in 
series series with pump one, as it on CMP.A1. To make a long story short 
it turns out turns out that  in addition to the valves in the VMB, the last 12"
of the return line was clogged with slurry.  Cut and flared and installed a new 
return line back to the manifold.   Slurry calibration was verified.
During breakin, after 10 successful wafers,  one wafer flew out of the carrier 
and broke on the tape.  Pad and carries have been changed. System is in MQC.

	Water valve has not arrived from Strasbaugh,  and due to the hectic 
nature of the day, we have not found time to follow up on this shipment.. 
Although the bypass seems to be working fine.


					Tim/Doc/Brendan
7.103changed v31STRATA::FINNTue Nov 28 1995 19:197
	Picked up valve at Fab 6 dock today.  
        Replaced water valve 31. Tested aok.  
        System up and running product.

					Brendan/Jim T.

7.104Uniformity issuesSUBPAC::BJUBINVILLENFD Fire - RescueThu Nov 30 1995 08:4214
    
            Since early Wed. morning, the left carrier showed
        a 6% uniformity difference from the right.  A pad and
        carriers change was done due to life.  All day Wed. it 
        still showed signs of problems with the left.  At 8:10pm,
        the tool crashed a wafer from the left spindle.  It was
        a reclaim wafer to boot.  After the tool was cleaned up,
        I checked both hubs due to the uniformity.  I Found the left
        hub out .08 mils and the right was at .0025.  They were both
        brought in within .0005 mils. 
      
                                                 BJ
            
                                                              
7.105Semi annual completedFABSIX::B_FINNWed Dec 06 1995 19:1213
	Completed semi annual pm.  Pad and carrier changed.  System
turned over to manufacturing for breakin and MQC. Left carrier passed Right 
carrier had edged heavy removal.  Carrier was changed.  Wafer broke on pad fro 
new carrier.  Pad and carrier were changed. 

The last MQC has heavy center removal, shapes looks good,  operation tech is 
running more wafers to bring Sdev into spec.

	


					Brendan
7.106Robot vacuum lineFABSIX::R_GEEMon Dec 11 1995 06:516
	Replaced the 1/8" robot end effector vacuum line.
	      
					B shift

	   

7.107broken waferFABSIX::R_GEETue Dec 12 1995 07:0214
	* Broke a wafer on the pad.

		476 wafers on the carriers.
		232 wafers on the pad.
		1st 2 wafers of the 2nd 3 min. run.
		looks like wafer broke inside the carrier with 1 piece
		that had slid out.

		Replaced the pad and carriers, the system is now in mqc's.

					B shift

	   

7.108LUDWIG::FINNTue Dec 12 1995 19:287


	Wafer broke on pad. Pad and carrier changed, Breakin and MQC to start.


						Brendan/John
7.109Wafer BreakageASDG::POIRIERWed Dec 13 1995 19:167
     
    Wafer breakage:
    --------------
    
    The system experienced a wafer fly-out and breakage at 7:00 p.m. 
    D shift, please start the clean-up activity and replace the Pad & Carriers.
    
7.110WAFER BREAKAGESUBPAC::BJUBINVILLENFD Fire - RescueThu Dec 21 1995 06:3810
    
         A wafer slid out of the right carrier and crashed.
      The wafer was an exercise oxide wafer.  After the carrier
      was torn down.  A faulty check valve seemed to be the 
      culprit.  The tool was cleaned up and new carriers and
      pad was installed.  After the first round of MQC's, it 
      failed for unifomity.  OPS are running more wafers to 
      try and bring it in.
                                               D shift
    
7.111Start-up for 12/29SUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 29 1995 15:1012
    
    
        The DI came on but is still out of spec so we didn't want
      to keep it to the tool so we shut off the DI feed and opened the
      by-pass.  Facilities still have some by-passes opened, that's
      why we don't have very high pressure.
    
        The slurry lines are all back to normal and slurry is flowing.
      We put the filter on but but we didn't run the slurry through
      the machine because of the DI situation.
     
                                                       Mitch,John,Leon,BJ
7.112Shuttle Adjustments / Air Motor PerformanceASDG::POIRIERSat Dec 30 1995 15:3927
    
    Shuttle:
    --------
    The Shuttle was not moving smoothly so Mitch and I decided to lube &
    readjust the assembly. The rails were cleaned and relubed. We opened-up
    the Needle Valve Flow Controllers to increase the Shuttle speed. After
    we reached the desired speed, (same speed as CMP.A1's Shuttle) the 
    Shuttle was banging hard against the front stop. Mitch readjusted the
    Stop dampner but found it difficult adjust for a smooth deceleration.
    It appears that the unit is failing. We'll have to call Strasbaugh 
    this week for a replacement.
                                                   
    
    Cleaning Station Air Motors:
    ----------------------------
    Since the "RIGHT" Air Motor an CMP.A1 was not performing normally, we
    decided to start-up the Air Motors on CMP.A2 and check them out. The
    "Right" Air Motor on CMP.A2 would not rotate. We let the Motor sit
    for 5 min with air pressure applied and it started to turn slowly. 
    After 20 min, the motor appeared to rotate at full speed although
    somewhat noisy. We should probably replace all 4 motors at a time
    since they don't seem to have a long life. Approx 9 months.
    
    Note: The Left Motor on CMP.A1 was repaced in November 95, the other
          3 have been on the tools approx 9 months.   
    
    
7.113CMP.A2 START-UP COMPLETESUBPAC::BJUBINVILLENFD Fire - RescueSun Dec 31 1995 14:3243

                    CMP.A2 STARTUP CHECKSHEET
                    -------------------------

                                                       Shift       Date
                                                       -------     -------
  1.  Ensure that all necessary facilities are         B, C + D     12/31/95  
      available and ready to use.                      

  2.  Power up each system.                            B, C + D     12/31/95  
                                                       
  3.  Power up each vacuum pump cabinet.               B, C + D     12/31/95  

  4.  Power up each Neslab TCU.                        B, C + D     12/31/95  

  5.  Open incoming DI Water Valve to each             B, C + D     12/31/95
      system.                                       

  6.  Close UDR valve at each system.                  B, C + D     12/31/95

  7.  Install wafer carriers with new films.           B, C + D     12/31/95

  8.  Place cassettes in left and right send
      and receive stations to fill DI Water
      tanks.                                       

  9.  Re-connect slurry inlet and outlet lines         B, C + D     12/31/95
      to filter housings and install new filters.    

 10.  Change RTC status to SCHEDULED DOWN and
      perform a Weekly PM on each system.           
  
 11.  Install new polishing pads.               

 12.  Re-install PVA scrub brushes.                    B, C + D     12/31/95

 13.  Place systems in wet idle mode.                  B, C + D     12/31/95

 14.  Change RTC status to VERIFY and notify
      Operations that the systems are ready
      for MQC.                                        
                                                        
7.114slurry line / broken wafersFABSIX::R_GEETue Jan 02 1996 07:4613
	 The slurry line for pump #1 was leaking, replaced pump #1 
	 slurry line. Calibrated the slurry flow for both pump #1 
	 and #2. 

	 * After the system passed mqc's 2 breakin wafers broke on the pad.
           the pad was changed and carriers from cmp.a1 installed.

	   The system is now in verify_u and running breakin now.
	  	
					B shift

	   

7.115Wafer breakage....SUBPAC::BJUBINVILLENFD Fire - RescueThu Jan 04 1996 08:258
    
           A break-in wafer slid out for the left carrier and
         blew away.  Everything was checked out and we came up 
         empty.  The tool was cleaned up and brought back up for
         production with 45 minutes or so.  The wafer count for
         both carriers was around 250 and the pad count was almost
         400.
                                                 D shift
7.116wafer slippageSTRATA::RIDLONEliminate the obviousSun Jan 07 1996 06:324
    Product wafer slipped out of right carrier. Had to cut a 32" inch pad
    down to fit. Cleaned up the mess and changed the carriers. 
    
    					B shift
7.117Wafer breakage, pmSUBPAC::BJUBINVILLENFD Fire - RescueWed Jan 10 1996 06:114
    

           A wafer slid out of the right carrier and blew apart.
       We took this opportunity to complete the monthly PM.  
7.118Bridge not arriving at polishFABSIX::R_GEEWed Jan 17 1996 07:0411
	The bridge stopped midway between the load cups and the polish
	table., before polish.  No reason was found for the malfunction, 
	the error was, "Bridge not arriving at polish". 

	Had to unload the wafers and home the bridge to recover, ran
	fine the rest of the night.
	  	
					B shift

	   

7.119LEFT SEND LEFT ELEVATOR NOT REPEATINGSUBPAC::BJUBINVILLENFD Fire - RescueSat Jan 20 1996 05:2410
    
        The robot was hitting the wafers past the 15th slot
     in the left send elevator.  After homing, cycling the 
     power and swapping cassettes, the problem stayed with
     the left elevator.  Due to a hot lot and A1 going down
     in a few minutes, we are unable to take the elevator out
     to have a look.  A2 is still with the right on-line but 
     there is a note on the machine as well as a WARNING on
     in.  The elevator is not repeating so it could be anything 
     from the lead screw to the possible corroded wiring,
7.120New airmotors have arrivedSUBPAC::TOOKERMon Jan 22 1996 15:068
CMP Equip. Techs,

I have just received 4 new air motors for the cmp.a1/.a2 clean stations.  These
need to be installed as soon as possible.  I will leave them on my desk in a
plastic bag.

Thanks,
Jim Tooker
7.121Slurry #1 line leak.FABSIX::R_GEETue Jan 23 1996 06:577
	Slurry pump #1 line was leaking. We replaced the line and 
	calibrated the slurry flow for pump #1.
	  	
					B shift

	   

7.122Air brush changeFABSIX::B_FINNTue Jan 23 1996 18:416
    
    Changed both air motors at scrub/clean station, per Jim Tooker request.
    MQC's pasted fine, system was up but will be going down for a normal
    pad change @ 272 wafers. 
    
    					A shift 
7.123pad change, left send elevatorSUBPAC::BJUBINVILLENFD Fire - RescueThu Jan 25 1996 06:507

      Bubbles we found on pad.  Replaced pad and carriers.

      OPS used the LEFT SEND elevator with no problems this morning.

                                          BJ
7.124bubbles between the pads/ broke waferFABSIX::R_GEESun Jan 28 1996 08:1912
	Had to remove one pad do to bubbles between the pads after 114
	wafers. Did a second pad change do to poor uniformity
	after only 34 wafers, it failed 2 mqc attempts. The system
	is still down for poor uniformity 12% and may need a third
	pad change.
	At the end of our shift 2 wafers broke on the pad, it needs
	a pad and carriers.
	
					B shift

	   

7.125robot upgradeSTRATA::FINNTue Feb 06 1996 19:1013
	System is down for the scheduled robot upgrade. Robot rails and 
bearings have been removed.  New rail assembly has been installed 
with new hardware and aligned to motor.  Boot has been reinstalled over rails.
Wires have been run and spliced, pins have been installed. New pressure gauge 
and bracket have been installed.

	Wire connectors will be installed tomorrow.  Robot and Software are 
still to go. Project is on Schedule.
 

			                	A shift
                          
7.126robot upgrade statusSUBPAC::LANDRYWed Feb 07 1996 19:0321
    More good progress with the upgrade today..........
    
    -  ALL wiring/connectors/hookups completed.
    -  Software was sucessfully installed.
    -  Robot functionality has been verified.
    -  ALL homing sequences have been verified.
    -  Some robot alignments were performed.  Will need to check again
       in the morning.  Will also have to perform other alignments/
       calibrations.
    
    -  Two new carriers have been loaded up.
    -  Both brushes have been re-installed.
    -  The pad will need to be changed PRIOR to any wafers touching down.
       PLEASE DO NOT put a new one on until Todd has finished with the
       remaining alignments/cal's/verifications.
    
    ***  Please leave the system as is.  Todd has several things to
         do before we try any polishing.
    
    					Doc  /  Todd F.
    
7.127Strasbaugh Robot UpgradeASDG::POIRIERThu Feb 08 1996 20:299
    
     * Robot Upgrade:
       -------------
       The Strasbaugh Rep ran into problems with the new Robot. A 2nd unit
       will arrive at 10:00 p.m. tonight. He may resume the upgrade work
       tonight. If not, he will be back in the morning with a second 
       Strasbaugh field service tech.
        
      
7.128Robot upgradeSUBPAC::BJUBINVILLENFD Fire - RescueFri Feb 09 1996 05:177

     The problem with the first new robot was either the motor or the
  encoder for the in/out axis.  The 2nd robot checked out ok.  Todd
  went home and will finish the upgrade in the morning.

                                                 BJ
7.129cmp.a2 robot upgrade updateSUBPAC::TOOKERFri Feb 09 1996 16:5913
                                                                     Jim Tooker
                                                                     2/09/96



The new new robot in cmp.a2 is now operational.  At ~5:15 pm we will be 
ready to return the tool to Operations for a new pad, carriers, and MQC.
The Strasbaugh field service engineer will be staying late to monitor
the tool's peformance.  Please call me at home, 508-874-0839, if there
are any problems or questions.

Thanks,
Jim
7.130Robot Upgrade completedASDG::POIRIERFri Feb 09 1996 19:275
    
    * Robot upgrade:
      -------------
      The robot upgrade has been completed. The system is currently running
      MQC's. 
7.131Robot Vacuum Sensor needs recalibrationASDG::POIRIERSat Feb 10 1996 19:4610
    
    * Robot failing to pick-up wafers:
      -------------------------------
      It appears that the vacuum sensor is in need of adjustment. The same
      condition developed on cmp.A1's upgraded robot last week and a 
      sensor recalibration solved the problem. The procedure for cali-
      brating the (new) robot vacuum sensor is in Vol #2 of the Stras-
      baugh maintenance manual.
    
              
7.132right carrier stickingSTRATA::RIDLONEliminate the obviousSun Feb 11 1996 01:019
    Right carrier sticking to pad on lift off. Pad was changed on previous
    shift and operator was trying to qualify tool. Ran pad condition and 
    retested with no improvement. Removed right carrier and checked
    tolerance of retainer ring. Result was -.0055in to -.0070in. There were
    three rebuilt carriers so one of these was installed. Ran some more
    breakins and the carrier lifted off as it should repeatedly. 
    
    
    					B shift
7.133ScratchesFABSIX::R_GEESun Feb 11 1996 07:0212
	* The system failed for scratches 2 times. Traced the
	  scratches to the polish table area. 

		Replaced the pad and conditioner disk, cleaned 
		the system, it still failed for scratches. Replaced
    		the pad and conditioner a second time. "A" shift will 
    		need to keep working on it.
	
					B shift

	   

7.134MQC's are fineSTRATA::FINNSun Feb 11 1996 19:236
	Jesus and Kim ran breakin and MQC's after B shift pad change.
        No scratches and particles were in spec. Fix belongs to B shift.


					A shift
7.135Slurry Pump hoses replacedASDG::POIRIERThu Feb 15 1996 19:276
    
    * Slurry lines:
      ------------
      The hose on Slurry Pump #1 ruptured this afternoon. Brad replaced
      the hoses on both pumps and returned the tool to Manufacturing.
    
7.136Slurry lines adjusted.FABSIX::B_WESTPHALThu Feb 15 1996 19:403
CMP.A2 - Down for slurry pump #1 leaking.  The line was cracked so replaced
	 both #1 & #2's slurry lines.  Also changed the pad as requested by
	 production.
7.137Blown slurry line on pump #1SUBPAC::BJUBINVILLENFD Fire - RescueFri Feb 16 1996 07:289

      The slurry line on pump #1 was too long and pinched off causing 
   a tear and causing slurry to spray everywhere.  It also caused the
   tool to stop during polishing.  No wafers were damaged.  The line 
   was replaced and both pumps were recalibrated.
      
                                                  D shift
                                                 
7.138Softt ware bugSUBPAC::TOOKERFri Feb 16 1996 14:3315
On cmp.a1 and cmp.a2 there is a bug in the software that will occasionally
cause the autocycle to abort and go to idle.  This occurs after the robot
has unloaded wafers and before trying to load more wafers.  At this point,
the system checks to see if there are any wafers present on the robot before
trying to take any wafers from a load cassette.  Strasbaugh says that the
time delay in the software is too short and this is causing the system to
eroneously sense the presence of a wafer.  Next week, Strasbaugh is sending
a new software Rev. to us that will correct this problem.  If the auto cycle
does abort to idle, hit "start" to resume the autocycle.  Maintenance techs.,
please do not adjust the vacuum switch as a fix for this problem.  Only adjust
the vacuum switch when the robot misses picking up a wafer from a cassette.
(i.e. the robot moves into the cassette and comes back out without a wafer)

Thanks,
Jim T.
7.139No response from the Computer BezilSUBPAC::LANDRYSun Feb 18 1996 14:449
    This afternoon, the opts people were unable to activate any touch pads
    on the front computer.  It originally appeared that the system had
    locked up.  Further checking proved otherwise.  The remote terminal was
    having no problems with keyboard use.  So, back to the front computer
    for a look_see.  Found the keyboard cable just barely inserted into the
    interface module (which is mounted on the back wall directly in back of
    the computer).  Re-inserted said cable.  Computer bezil working fine.
    
    						A shift
7.140Robot/Cassette problemFABSIX::R_POIRIERWed Feb 28 1996 05:0615

	Operations had a problem with robot crashing into and/or
	not picking up wafers from cassette. We found that a wrong 
	cassette - Empak P/N XT204-02 was being used. Should be 
	XT200-02E. 

	XT204-02 has a wafer gap that is about 1/16th of an inch 
	bigger than the XT200-02E.




					B Shift

7.141Conditioner arm up sensor bracketFABSIX::R_GEEWed Feb 28 1996 07:2714
	* The conditioner was erroring for "conditioner not arrive up"
	  it was up.

		Found a broke off screw that holds the up sensor
	        on. There are 2 screws holding it on. We tightened
		up the one remaining. Production requested that they 
		beable to finish the lots running. After these have   
		been completed it will need to be drilled out and 
		retaped.
	
					B shift

	   

7.142passdownFABSIX::B_FINNWed Feb 28 1996 19:348
	One of the screws holding the Pad Conditioner up sensor 
        snapped off in it hole. Removed bracket and sensor, drilled out
        old screw and re taped to 10-32. Reinstalled and tested aok.  
        Completed weekly pm. 


					A shift
7.143New wafer push bars installedSUBPAC::TOOKERFri Mar 01 1996 15:399
Today we installed new style wafer push bars in cmp.a1 and cmp.a2.  These
are the bars located inside each elevator rinse tank to help keep the wafers
from floating out of the cassettes.  The new bars are thicker than the old
ones and should do a much better job of keeping the wafers in the cassettes.
Please ensure that the wafers are completely in the cassette when placing a
cassette of wafers onto a cassette station.

Jim T.

7.144pads...........SUBPAC::LANDRYTue Mar 05 1996 19:338
      The system went down late morning for scratches.  Process Eng
      decided to try an expiremental pad (spiral) at this time.  It
      was installed, broken in and MQC's run.  Poor uniformity was
      the result.  The pad is currently being removed in favor of a
      normal pad.  It will need the usual break_in and MQC's run.
    
      					A Shift
7.145Wafer breakageFABSIX::B_JUBINVILLEFri Mar 08 1996 06:5612
    

           During the running of the MQC wafers after the 
       monthly PM was completed, a ~crash~ consisting of 
       two wafers and their respective carriers took place.
       The small pieces of film that was as a result of not
       being blown out during assembly was evident on the
       brushes, the wafers, pad and check valves.  After
       everything was cleaned up, new carriers and pad  
       were put on the tool.

                                      BJ
7.146wafer breakage, brushesSUBPAC::BJUBINVILLENFD Fire - RescueSat Mar 09 1996 04:528

        While running pre-break-in wafers, the tool crashed.  The 
    tool was cleaned up and wiped down.  The brushes were removed 
    to bring up A1.  This system will have to stay down until
    Monday when VSR-1 is repaired.

                                   BJ
7.147brushes foundSTRATA::RIDLONEliminate the obviousSun Mar 10 1996 00:537
    Found the keys to open the lower door of VSR-1 and as luck would have
    it thats where draw 14 was. Removed the brushes needed to bring A2 back
    up. New pad and carriers being installed now. 
    
    The stock room was notified about the removal of the two brushes.
    
    						B shift
7.148Robot End EffectorFABSIX::R_GEEMon Mar 11 1996 03:3112
	* After swapping the end effector from CMP.A1 the problem
	  with the robot not detecting the wafer on the robot moved
	  to CMP.A2. We found a small burr on the end effector created
	  by the end effector hitting wafers in the past. We scrapped
	  the burr off and siliconed the stainless steel tube where
	  it meets the end effector, looked like it could have been 
	  leaking vacuum.  No more robot problems.

					B shift

	   

7.149robot to left send station positionFABSIX::R_GEEWed Mar 13 1996 02:3112
	The robot wouldn't pickup the wafer out of the left send
	station.

	We re-homed the robot and elevator, still wouldn't work. The
	robot end effector wasn't low enough. We adjusted calibration
	variable # 80 "left send elevator offset to slot 0" from 3.02 
	to 3.05, no more problems.

					B shift

	   

7.150BROKEN WAFERSFABSIX::B_WESTPHALThu Mar 14 1996 19:066
    
CMP.A1 - Down for broken wafers on the polish table.  Could not determine any
	 definite cause.  Pulled the pad and cleaned up the process area.  The 
	 carriers were replaced and the conditioning disc was checked and 
	 cleaned.

7.151Sslipping wafersFABSIX::B_WESTPHALFri Mar 15 1996 19:4110
CMP.A2 - Down for breaking wafers during polish.  Originally Suspected poor slurry flow to be
	 the problem.  The filter was checked and it was heavily clogged with 
	 slurry.  The filter was Changed out with a new one that was sent in 
	 FEDEX.  The polish area was cleaned and pad and carriers were changed
	 by production. The problem reoccurred while ops was tring to run there 
	 breakins.  Pulled the pad and cleaned up the mess.  Checked the
	 carriers for any obvious defects, none were found.  O-rings between 
	 the spindle and carrier.  The same seemed good.  Ran out of time.  The
	 tool presently sets with 2 new carriers and no pad on the table.

7.152MQC failuresSUBPAC::BJUBINVILLENFD Fire - RescueSat Mar 16 1996 07:0012

            The tool was hurtin' all night running wafers
        and MQC's.  The hub alignments were checked and
        they were ok.  We swapped the carriers with (2)
        new carriers that were sitting on A1 and they
        re-ran MQC's.  This time the left carrier came
        right into spec at 6%.  The right was at 12%.
        Another new carrier was placed on the right spindle 
        and as of 7am, they are running MQC's once again.

                                             BJ
7.153scratchesSTRATA::FINNMon Mar 18 1996 19:309


	System down for scratches, wafer were broke on an engineering
        experimental run.  Table Amps increased and wafers broke. Pad and 
        carrier changed.  Mqc showed scratched.  Pad about to be changed.
                                                                        

			                	A shift
7.154System was in a funk for a couple hoursSUBPAC::LANDRYTue Mar 19 1996 18:2433
    What a fiasco on this baby today.............
    
    The system was crankin' along fine (running product) until about 2:00
    PM, when all of a sudden, it unloaded both carriers in the middle of
    the product run.  It dropped them off in the clean station as if it
    told to do it.  Of course I was in the process of a Semi-Annual PM on
    A1 at the time.  Don't it figure?  So, the vacuum system was checked 
    and it was fine.  The carriers were loaded up and we attempted to get 
    back on-line.
    
    Well, first attempt, it dropped a wafer from the left carrier.  A small
    piece of material (looked like a shaving off the film) was with the 
    wafer on the table.  Kim was quick to EMO, hence preventing any crash.
    Checked the carrier and it looked alright.  Retried again.  This time,
    there was no rotation at the left carrier, EMO'd and saved a potential
    crash.  Rebooted and manually tried rotating the left carrier.  Worked
    fine.
    
    OK, lets try running again.  First try, the Amps on the table rose 
    dramatically, emo'd again.  Rebooted, checked the carriers again.  The
    left carrier was showing signs of being a problem.  Looked like small
    shavings of the retaining ring were now on the table.  Alright, get
    this carrier outta here.  Put a new one on the left spindle, ran the
    system as if breaking in a new pad (low down force to begin; gradually
    increasing down force each run).  The system seemed to be back in
    proper working order dispite the multitude of inadequacies the previous
    couple hours.  They ran MQC's with good results.  Wish I could tell
    you what this fiasco was all about, but NONE of it made any sense.
    
    Get me away from this tool......please!  I've had enough.
    
    			Doc......with help and amazement from Kyle
    
7.155System upgraded with new load and clean station ringsSUBPAC::TOOKERWed Mar 20 1996 16:278
CMP.A1 and CMP.A2 have been upgraded with new load station rings and new
6 roller clean station rings.  I have checked the alignment between the
spindles and the new rings, and cycled 50 wafers through each system with
the handling recipe without any problems.  Please be aware that if either
system starts to have misloads, that a re-alignment may be necessary.

thanks,
Jim T.
7.156Crash... again...SUBPAC::BJUBINVILLENFD Fire - RescueThu Mar 21 1996 05:5128

         the latest definition of CMP:  
 
              Continuously Mauling Pads

            Came to a tool that was just re-assembled
         as a result of yet another crash.  It just
         needed some carriers built.  After I built up
         some carriers and got A2, running it crashed
         as it got to 20 wafers on the pad and at 313lb
         downforce.  The wafer in the left carriers was
         the culprit.  I checked the downforce on both
         spindle and found the right just a bit out
         of spec and brought it in.  The left spindle 
         was a different case.  On my first pass, I 
         got a force error of -21, it's suppose to be 
         ~+/- 2.  I was able to bring it in according to
         spec.  I decided to check the EP regulators but
         I was following the spec and "it ain't happenin'!"
         It is in reference to the old software but I don't
         think that mattered much.  I called the 24-hotline
         and got no response.   They don't need the tool so 
         I'm leaving it down and I'm going to speak to EE
         at 8am.
                                            BJ
         
      
7.157Performed left & right spindle downforce and backpressure calibrationSUBPAC::TOOKERThu Mar 21 1996 16:1213
Today the left and right spindle downforce and backpressure was calibrated on cmp.a2.
The Strasbaugh manual is a little ambiguous when it talks about the different Fairchild
E-P regulators.  Please remember that the three small Fairchilds bundled together in the
lower right hand corner of the electronics cabinet are for left spindle backpressure,
right spindle backpressure, and pad conditioning downforce.  The two large Fairchilds
located above this are for left and right spindle downforce.  The calibration procedures
for downforce and backpressure are located in the back section of the vol. 2 Strasbaugh
manual.  I will be ordering two new sets of manuals from Strasbaugh.  This should help
some of the confusion created by the new software with changed calibration variable
numbers and analog output nomenclature.

Thanks,
Jim T.
7.158More crashes and ENG experimentsSUBPAC::BJUBINVILLENFD Fire - RescueFri Mar 22 1996 05:0018

          Another night more slides and crashes.  The carriers
      on the tool were the new style without the check valves
      and no o-rings.  The back pressure is not being used. 
      After the second crash on the night, removed the center
      plug as well as the first inner ring of plugs in the
      right carrier and left the left carrier as is with just
      the outer plugs out.  We had over currents and two more
      slide out near crashes.  The tool has a new pad and has
      been conditioned.  The pad was used in the last experiment.
      The current condition of the carriers have the left being
      the new ENG way with just the outer holes open and the right 
      has the outer ring, inner ring and center hole opened.  
      I changed the tool to status DOWN/PROCESS.
 
 
                                                     BJ
7.159Wafer Fly-Outs - Cut Down Speedfam 32" IC1000 Pads ASDG::POIRIERFri Mar 22 1996 19:3921
    
    * Wafer Fly-out Problem:
      ---------------------
    
      The system was flying-out wafers all day during the break-in recipe.
      At Break-in recipe #3, the Table Controller Amperage would Spike-up
      to 35 AMPs is Step #3. (@313 lbs of down force)
    
      We even went to the stockroom and tried a different lot # of 28" pad.
      The problem persisted!
    
      At 6:00p.m., we cut down a 32" Speedfam IC1000 pad. The Amperage
      spiked-up to 32, dangerously close to the 35 amp fly-out value.
      However, NO wafers flew-out. On the next two wafers, the Amperage 
      spiked to 27. It spiked lower and lowaer thru consecutive runs.
      It appears that the Speedfam pads are able to break-in nicely. 
    
      We recommend using only Speedfam 32" IC 1000 pads on the Strasbaugh's
      for the next shift. Process Engingeering will be in tomorrow morning
      to continue with analysis of this problem.
    
7.160rotation axis errorSTRATA::RIDLONEliminate the obviousMon Mar 25 1996 05:1910
    It was noticed by the operator that the right spindle was going down
    slower than the left and that removal rates were 1000 angstroms
    different from left to right. The tool then errored for rotation axis
    error. The right and left ep regulators were checked and both were ok.
    Down force was also checked and again both were ok. During our
    investigation we found that the right air cylinder had a leak at the
    clippard fitting. AFter the leak was repaired both spindles went down
    at the same rate and force. 
    
    						B shift
7.161Wafer CrashFABSIX::R_POIRIERWed Mar 27 1996 07:0510

A2 crashed last night. Wafer came out of left carrier.
The carriers were new as was the pad. Replaced pad and
carriers and Ops. ran break-in and MQC's. Passed MQC's.



				B Shift

7.162CMP.A2 shutdown for facilities tie-in completedSUBPAC::BJUBINVILLENFD Fire - RescueThu Mar 28 1996 06:0757

  SHUTDOWN:  5:00 AM ON 3/28/96

  STARTUP:   9:00 PM ON 2/28/96


                    CMP.A2 SHUTDOWN CHECKSHEET
                    --------------------------


                                                        Tech.       Date
                                                       -------     -------


  1.  Change RTC status to SCHEDULED DOWN.               BJ        3/28/96

  2.  Remove all wafer carriers.  If possible
      keep them wet.  If not ensure that there
      are enough carriers with new films for
      startup.                                           BJ        3/28/96

  3.  Remove all wafer cassettes and drain Send
      and Receive DI Water tanks.                        BJ        3/28/96

  4.  Purge both slurry lines with DI Water.             BJ        3/28/96

  5.  Remove and discard all polishing pads.             BJ        3/28/96

  6.  Remove all PVA scrub brushes and submerge         
      in DI water.                                       BJ        3/28/96

  7.  Perform a complete clean of each system
      to remove dried slurry.  Wipe down and dry
      all wet surfaces.                                  BJ        3/28/96

  8.  Remove and discard all slurry filters.             BJ        3/28/96
                                                         
  9.  Disconnect inlet and outlet lines on each
      filter housing and remove housing.  Connect
      inlet and outlet lines together with a 3/8"
      union.  This will completely bypass the 
      filter housings.                                   BJ        3/28/96

 10.  Close main incoming DI Water valve to each
      system and open the DI Water return valve
      (UDR).  This will allow all incoming DI Water
      to flow to UDR.                                    BJ        3/28/96

 11.  Power down each system.                            BJ        3/28/96

 12.  Power down each vacuum pump cabinet.               BJ        3/28/96

 13.  Power down each Neslab TCU.                        BJ        3/28/96
    
    
                                                                       
7.163carrier issuesSUBPAC::BJUBINVILLENFD Fire - RescueFri Mar 29 1996 06:3414

       After just getting the tool up from the shut-down,
    both carriers kept sticking to the table.  When they 
    did finally come up, the wafers were being left on the
    table.  I took the right carrier apart only find that
    one of the plate screws snapped off upon dis-assembly.
    New carrier for right...  There appeared to be nothing
    wrong with the left upon inspection so it went back
    on the tool.  The right carrier was now ok but the left
    still left wafers on the table.  I put a new carrier on
    the left and the tool ran fine after that.

                                          BJ
7.164software upgrade completedSTRATA::FINNSun Mar 31 1996 19:3711
	Pad was changed due to unifomity and wafers sticking to pad 
        at lift off. 

        Performed weekly pm. 
    
        Upgraded Software to Verersion. 4.16.3.  No problems.

	Mqc's passed first time.

			                	A shift
7.165Wafer crashFABSIX::R_POIRIERMon Apr 01 1996 08:167
         We lost a product wafer out of the left carrier at lift-off.
    	 It slid under the right carrier and both were destroyed. 
    	 We replaced the pad and carriers and tool is running break-in
    	 and MQC's
    
    						B Shift
    
7.166new water nozzles addedSTRATA::FINNMon Apr 01 1996 19:207
	Added new additional spray nozzles to the left and right clean 
        stations. Spliced into the nozzles for the load stations.

	Tested aok.

			                	A shift
7.167breakin wafers brokenFABSIX::B_FINNSun Apr 07 1996 20:169
	Ran hot lot fine, after an idle period breakin wafers were 
        run, prior to next lot. Breakin wafer from right spindle flew out 
        at lift off.  Monitor wafers looked a little bare. Cleaned system 
        and replaced pads and both carriers. Mqc's being run.


						A shift

7.168New Strasbaugh maintenance manuals are hereSUBPAC::TOOKERMon Apr 08 1996 13:4212
We have received 2 new sets of maintenance manuals for cmp.a1 & cmp.a2.  These
new manuals are marked ver. 3.  One set has been placed in the gray cabinet
in the CMP service aisle and the other set is in my office.  These new manuals
reflect all the changes brought about by the new software rev. and the new
style robot.  The old manual containing the schematics has been left in the
service aisle.  You may find that in some cases these will need to be 
referenced.  Please be gentle with these manuals and ensure that pages are not
ripped out or lost.  These manuals cost $1200.00 per set.


Thanks,
Jim
7.169Left spindle air leakFABSIX::R_GEEWed Apr 10 1996 08:1011
	When we were performing the left spindle down force part of the 
	Monthly p.m. we noticed the down force readings were bouncing
	30 lbs. at the 300 lb setting. 
	We found 2 air leaks at the air cylinder, one at the clippered 
	valve and one at the connector next to the air cylinder. We 
	tightened the clippered valve and replaced the connector, all set.
				
					B shift

	   

7.170SLIPPING WAFERSLUDWIG::WESTPHALThu Apr 11 1996 20:336
CMP.A2 - Slipping wafers.  Changed the pad and ran Barry's modified breakin 
	 method.  where 200 lbs. of down force was applied for 4 wafers than 
	 250, 260, 270, 280, 290 300, 313.  This was done because the fly outs
	 were thought to be caused by the pad not breaking in.   Process also 
	 extended the rate of ramp to help aliminate friction on the pad.

7.171broken wafer / poor uniformity / poor rangeFABSIX::R_GEETue Apr 16 1996 08:4516
	The system failed MQC's for poor uniformity 21% and
	removal rate between the spindles of 1000 ang. We
	found that the left carrier exposure was only .004 to
	.0045.
	We replaced the pad and carriers then began running breakin
	wafers. On the second set of breakin wafers, before
	the second MQC check a wafer slid out from under the right
	carrier. 
	The system has been cleaned and will need a down force check,
	carriers and a pad. Carriers are being built as I type this.

	
					B shift

	   

7.172Down Force checkedSUBPAC::LANDRYTue Apr 16 1996 20:3614
Today......

We arrived to find the system had just broken a wafer.  B Shift found
a huge difference in removal between the two spindles.  We, along with
Bobby G, performed down force calibration on both spindles.  Both were 
out of spec a tad bit, but consistant between each.  We had expected one 
spindle to differ from the other in a major way, but no, they were
close to each other.

One of the carriers was found to be built with an extention ring gap
out of spec.  The system (at 7:30PM) is being "broken_in".  MQC's to
follow.  

      						A Shift
7.173scheduled DI work doneSUBPAC::LANDRYTue Apr 23 1996 19:4810
      The system went down today for scheduled work to be performed
      by the contractors on the DI water line.  The pad was removed
      at this time and the carriers were also changed out.  The 
      brushes were removed and placed in water to remain intact.
      They were re-installed once the system was returned to us.
      A new pad and two new carriers were installed.  Breakin wafers
      and MQC's followed with positive results.  The system is
      cranking product.

      						A Shift
7.174Broken waferFABSIX::R_GEESun Apr 28 1996 07:5110
	At the beginning of the shift we broke a product wafer on 
	the polish table. A wafer slid out from under the right 
	carrier on the first run, in the middle of a 120 second polish. 
	We suspect the fly-out was caused by a bad carrier, it had 
	high wafer count, 387 and was due to be changed.
	
					B shift

	   

7.175vac sensorLUDWIG::FINNTue Apr 30 1996 20:3510

System would not move robot back to unload cassette from load cups, although 
end effect had a wafer.  Vacuum switch was not seeing the sensor, checked air 
pressure, check hose  then end effect, then  adjusted vacuum sensor. AOK 
gave back to ops for continuation of MQC.

						A shift 

 
7.176Down force calibration performed todaySUBPAC::TOOKERWed May 01 1996 16:1710
This morning cmp.a2 broke wafers on the pad and it was reported that there was
a difference of ~800 Ang/min in removal rates between the left and right
spindles.  John Swenson and I did a complete down force calibration including
E-P regulators.  We found that both E-P regulators were OK, but both spindles
were out of calibration by around 10%.  After the calibration, a new pad and
new carriers were installed.  The MQC wafers that were polished after breakin
had only a 10 Ang/min difference in the left and right removal rate
and the removal rate mean was very close to the mean of the SPC chart.

Jim T. & John S.
7.177SLURRY LEAKSTRATA::WESTPHALSat May 04 1996 20:145
CMP.A2 - Down for a slurrry leak.  Pump #1 slurry line had a crack in it.  
	 Removed and cleaned both pump heads and replaced the lines.  Tool
	 was returned to production.

7.178Broken waferFABSIX::R_GEESun May 05 1996 08:159
	Broke a wafer on the polish table. The wafer slid out
	from under the right carrier. Checked out the machine
	and found nothing. There was 234 wafers on the pad and 
	283 on the carriers. The system is now running MQC's.	

					B shift

	   

7.179Broken waferFABSIX::R_GEEMon May 06 1996 08:1312
	Broke a wafer on the polish table. 

	The wafer slid out from under the right carrier at lift off. 
	It was running a 10 sec touchup, both carriers stuck to the 
	pad. When the right carrier lifted off the pad the wafer stayed
	and hit the left. 77 wafers on the carriers and 154 on the pad.
	Suspect the crash was a pad or conditioner problem. 	

					B shift

	   

7.180New E-P regulator installed on left spindleSUBPAC::TOOKERThu May 09 1996 17:2110
Today a new E-P regulator was installed on the left spindle of cmp.a1.  After
calibration the +/- 10 lbs of downforce fluctuation previously seen at 
downforces greater than 350 lbs. on the old E-P regulator is now at +/- .2 lb.
Strasbaugh now uses the Fairchild T6000 E-P'S (same model that is used on
backpressure) in place of the big Fairchild E-P'S that we usually use for
downforce.  The T6000's require different analog counter settings for the
calibration.  I will publish these new analog counter settings tommorow.

Jim T.

7.181Fairchild E-P analog counter settings for calibrationsSUBPAC::TOOKERFri May 10 1996 13:2342
To all CMP ET's;

Here are the proper analog counter settings for calibrating the Fairchild E-P
transducers for left and right spindle downforce on cmp.a1/.a2.  These analog
counter settings apply only to the spindle downforce E-P's and do not apply
to the left and right backpressure or conditioning downforce E-P's.  Please
refer to the Strasbaugh manual for the proper counter settings for those E-P's.
Prior to any spindle downforce E-P calibrations, please check the E-P to see 
which model number it is and then use the appropriate analog counter setting
listed below.  The adjustments work the same on both models.  Adjust the
"ZERO" pot for 10.0 PSI and the "SPAN" pot for 80 PSI.  When either the "ZERO"
or "SPAN" is radically adjusted, it is sure to through out the other one, so
be sure to go back and forth between the two until the proper PSI is achieved.

You may have noted that the Strasbaugh manual calls for the "ZERO" pot to be 
adjusted to 5.0 PSI.  We adjust our tools to 10.0 PSI.  We do this (with Stras-
baughs blessing) to bias the spindles with a few pounds of downforce, so that 
when the spindles are manually lowered over the load station or cleaning sta-
tion they don't float upwards over time.

Please contact me with any questions or if anyone needs more one on one training
for this calibration.
================================================================================

      CMP.A1/.A2 LEFT OR RIGHT SPINDLE DOWNFORCE E-P TRANDUCER CALIBRATION
      --------------------------------------------------------------------

        FAIRCHILD E-P                    ANALOG COUNTER SETTINGS
                                         -----------------------
           MODEL #                10.0 +/- .2 PSI        80.0 +/- .2 PSI
        -------------             ---------------        ---------------

         Z18109-3                      580                    3410
        (old style)

         TT6000-06                     601                    3481
        (new style)

    NOTE:  If any of the E-P's are re-calibrated, then the downforce must
           also be re-calibrated.

================================================================================
7.182wafer breakageSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat May 11 1996 06:1813


         Tool crashed at beginning of shift.  A vacuum problem 
     with the right spindle is suspected.  We pulled the hub 
     off and found lots of debris inside, a lot...  We cleaned
     everything and inspected the tubing and o-rings inside.
     The hub was re-installed and re-aligned.  A new pad, carriers
     and brushes were installed.  MQC's were run and tool failed
     uniformity twice with no end in sight.  Another pad from another
     lot was installed.  We are awaiting MQC results....

                                                D shift    
7.183pad changeFABSIX::B_FINNSat May 18 1996 20:436
	Changed pad due to wafer count approaching upper limit.
        Wafer count 392. System is in MQC.


						Brendan
7.184Carrier problemFABSIX::R_GEESun May 19 1996 08:099
	The left carrier left a wafer on the pad, the wafer
	broke while placing it back onto the carrier. 
	Found a defect on the carrier film, replaced the carrier
	and conditioner disk. The system is running MQC's now.

					B shift

	   

7.185lift off delaySTRATA::FINNSun May 19 1996 20:0413

	The Right Carrier was significantly slower in lifting off from the pad 
        than the left carrier,after polish. Carrier was swapped with the left 
        carrier and problem stayed with the carrier. We replaced both the 
        white oring and the check valves on the suspected carrier. Tested aok.  
        All work was accomplished in between lots to reduce downtime, (your 
        welcome Jim).

	
						A shift 

 
7.186Condition Arm Down Force CalSUBPAC::LANDRYWed May 22 1996 17:5019
    The system had experienced much trouble trying to pass MQC's.  The
    first issue was concerning uniformity.  The second issue concerned
    the removal rate from spindle to spindle, which had a variation of
    ~400 � a minute.
    
    Issue one was resolved by performing down force calibration on the 
    Conditioning Arm, as well as installing a new diamond conditioning
    disk.  The down force was found to be 5 pounds off, which is quite
    substantial when you consider that we condition at 13 pounds DF.  
    Calibration procedure was thoroughly performed.
    
    Issue two still remains at this time.  The problem has been traced to
    a carrier.  After the early morning MQC attempt (and the arm cal),
    the carriers were swapped.  The problem shifted to the other spindle,
    hence identifying the carrier as the culprit.  Due to a vast workload,
    Opts people will run as is.  At the next MQC check, the problem carrier
    (or both carriers, depending on the wafer count) should be changed out.
    
    						Doc  /  Jim T
7.187robot sensor calibration.FABSIX::L_SHKOLNIKOVThu May 23 1996 06:537
    
    * Robot Sensor Adjustment:
    
      The Robot experienced wafer pick-up problems tonight.     
      The sensor was readjusted and It seems to be working fine now. 
    
      Leon                                  
7.188Slurry on Pad during Wet Idle ModeASDG::POIRIERFri May 24 1996 20:2510
    
    Slurry On Pad During Wet Idle Mode:
    ----------------------------------
    We just noticed that slurry is making its way past the rollers in the
    pump heads and building up on the pad during Wet Idle Mode.
    "D" Shift, please take a look at this problem as we are running out of
    time.
                                                                  thanks,
                                                                  
                                                                   C Shift.
7.189passdownSTRATA::FINNTue May 28 1996 20:2010
	Completed weekly pm and slurry cal, installed new carriers from rebuild 
cabinet. MQC's were started but plagued by wafers fly. Found check valves in 
backwards. Reinstalled. Mqc's passed aok.



						A shift 

 
7.190particlesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Jun 01 1996 05:278
       Tool keep failing for high particles.  The transfer program
    looked good.  The conditioner disk was changed again.  We looked
    at the rinse cycle and it was fine.  It was discovered that the
    slurry psi was up around 45 psi and slurry was leaking onto the
    table during wet idle.

                                                     D shift
7.191Robot vacuum problemFABSIX::R_GEEMon Jun 03 1996 08:1610
	The robot went down for failing to pickup wafers. We cleaned
	the robot end effector, check all lines for clogs & leaks, 
	replaced the 1/8" end effector line and adjusted the vacuum 
	switch. The problem has been narrow down to the vacuum switch
	itself, it looks like it will needs to be replaced. We couldn't 
	find one in stock.
					B shift

	   

7.192robot/wiring problemsSUBPAC::LANDRYMon Jun 03 1996 20:2137
      This morning, the system was not responding properly concerning
      the robot wafer sensor circuit.  It was first believed to have
      been a faulty vacuum switch.  Began looking at the problem.  
      Opened the electronics cabinet and then rebotted.  Could not
      home the system.  It would do the bridge ok, but the next axis
      (robot arm position) wasn't working properly.  The arm was trying
      to home moving in the opposite direction.  When the arm had
      extended out to it's physical limit, you could still see it
      attempting to go further, as it pulsed repeatably.  ESC'd out
      and retried several times.  The same result was consistant every
      time.  So, homing each axis individually was tried.  All homed
      fine except for the robot arm and robot rotation.

      Started tracing prints/wires/connectors.  Everything was looking
      fine.  Decided to take apart the connector (RS232) in the elec-
      tronics cabinet. The connector goes to the cable which terminates
      at the top connector on the robot.  Well, we've seen this before
      folks, water was in the connector.  The wires were corroded pretty
      bad.  Water had "wicked" it way from the robot all the way back
      to the electronics cabinet.  Same situation we saw with the 
      original robot assembly.  Of course we don't stock this cable, so
      the connections were cleaned in an attempt to temporarily get 
      around the problem.  Success!  All axis' were homing as should.
      My next concern, could we be so lucky that the vacuum sensor would
      now function.  Praise GOD!  It was responding to changes in 
      vacuum state.  Re-calibrated the switch, check results in manual
      and all was fine.  Ran two dozen wafers in AUTO CYCLE using the
      HANDLING recipe, excellent results.

      A cable will need to be ordered (and an RS 232 connector acquired
      as well) to replace the one that is in the system now, as it still 
      remains moist.  The corrosion WILL be back.  But, I feel we can 
      resume running for the time being.  The system was logged as an
      end_repair and logged down for a pad and carriers change.  There
      are currently NO MQC wafers available in the room.   
    
    						Doc
7.193robot missed againSTRATA::RIDLONEliminate the obviousWed Jun 05 1996 03:1610
    Robot crashed into a wafer at send cassette. Same problem as yesterday.
    Went immediately to the connector and it was soaking wet. Removed the
    metal connector jacket and stripped back the cable housing about 4
    inches. After drying and cleaning the connector a couple of times input
    21 finally toggled when a wafer was picked up in manual. Cycled 25
    wafers using no touch with out a mishap. The pad was replaced due to a
    larger than acceptable bubble in the center. 
    
    
    						b-shift
7.194Robot Cable Replaced / Cable not the same.ASDG::POIRIERSat Jun 08 1996 10:1916
    
    Robot Cable Replacement:
    -----------------------
    The Robot Cable has been replaced and tested. All Robot axes homed
    correctly. D Shift will reassemble the Cable Track and replace the
    covers.
    
    NOTE: The replacement cable has 3 out of the 12 wires that do
          not match the old cable. (color & pin location)
    
          Watch out for this situation if you ever have the pleasure of
          changing a Strasbaugh Robot Cable.
     
                                                              C Shift.
                                                                      
                                                  
7.195head to head removal / carrier problemFABSIX::R_GEESun Jun 09 1996 08:0611
	The system was down for head to head removal out of spec.
	it was +800 ang. and the left carrier was sticking to the
	pad, it had a new pad and carriers. We checked down force 
	per operations request, it looked good. We removed the 
	carriers, replaced the o-rings and the backing plate o-ring 
	bolts. The removal is now within 100 ang. between the heads.
  
					B shift

	   

7.196Particles - Cal Var #158 changed from 0 to 1ASDG::POIRIERWed Jun 12 1996 20:3812
    
    * Load Station Pedestals:
      ----------------------
      Gwen noticed scratches in the center of both wafers consistent with
      the round shape of the Load Pedestals. We found silicon debris
      embeded in the top surfaces of both Pedestals. They were cleaned
      MQC's were run. Particles were decreased but were still out of spec.
      
      Jim T. found calibration variable #158 set to 0 instead of 1. This 
      caused the vacuum to remain on during the unload segment of the auto
      cycle. The paramenter was changed to 1 (no vacuum) and MQC's were run
      again. This time the particles were in spec. 
7.197RobotFABSIX::R_GEETue Jun 18 1996 08:286
	Robot wouldn't pickup wafers, cleaned the end effector.
  
					B shift

	   

7.199bubble on pad/ weekly pmLUDWIG::RIDLONEliminate the obviousSun Jun 23 1996 06:374
    Bubble found on pad. Did the weekly while it was down.
    
    
    					B-shift
7.200particlesFABSIX::J_SWENSONFri Jun 28 1996 20:142
    particles failed.
    Adjusted scrub brush heights and rinsed system.
7.201Start up A2LUDWIG::SWENSONFri Jul 05 1996 20:383
    powered tool up with no problems.
    ordered new scrub brushes and retrieved.
    obtained new carriers for installation.
7.202Crash and PMSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed Jul 10 1996 08:1611
    
            Tool crashed a Metal-1 lot during polishing.
          Found pad barely sticking to base SUBA pad. 
          Basically, it looked like one huge bubble.
          The SUBA pad was completely soaked through.
          Not too normal from what I've seen in the 
          past.  Cleaned up tool and completed monthly
          while I was at it.  There are no complete carriers
          built-up so it's waiting for carriers and a pad.

                                            BJ
7.203system not detecting vacuum lossFABSIX::J_SWENSONThu Jul 11 1996 20:254
    wafer flew off right carrier and system did not detect vacuum loss.
    measured vacuum at both hubs and  found them to be @23in/hg.
    suspect clog in within carrier so replaced carrier.
    
7.204SCRATCHESFABSIX::J_SWENSONFri Jul 12 1996 20:1311
    DUE TO GLITCH TODAY DI WATER WAS LOST.
    REQUIRED PAD CHANGE AND CLEAN OF SYSTEM.
    PERFORMED PAD CHANGE AND CONDITIONING DISC.
    MQC FOUND TO HAVE MAJOR SCRATCHES.
    REMOVED SCRUB BRUSHES, RAN MQC STILL SCRATCHES.
    INSTALLED NEW PAD AND REPLACED CONDITIONING DISC WITH
    WITH ORIGINAL. MQC'S PASSED.
    
    NOTE:: THESE WERE NEW DISCS FROM THE STOCKROOM LOOKS TO BE A PROBLEM.
    
                       C-SHIFT
7.205Failed MQC'sFABSIX::R_GEEWed Jul 17 1996 08:0417
	Failed MQC's for uniformity and head to head removal.

	We did the weekly p.m., replaced the pad, carriers and 
	condition disk. The head to head removal came in tight
	but we found bad scratches coming from the pad. We felt
	that the scratches were coming from the conditioning disk.
	
	Replaced the pad and conditioning disk, (we filed the
	holes with burrs and wire brushed for loose diamonds).
	Ran MQC's again, the scratches are gone and particles
	passed, but the uniformity failed at 16 %. More breakin
	wafers are being run alone with another set of MQC's. 	

				B shift

	   

7.206right spindle uniformitySUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Jul 18 1996 07:4110

        A shift OPS put a WARNING on the right spindle because
    the left was in spec and the right was way out, ~22.  We
    swapped the carriers so MQC's will have to be run to see
    if the problem follows the carrier.


 
                                                     D shift
7.207particlesSTRATA::SWENSONFri Jul 19 1996 20:385
     bubble in pad and had particle problems on right carrier.
     replaced right carrier.
     replaced pad.
     cleaned system.
     put in wet idle.
7.208Wafers stickingFABSIX::R_GEEMon Jul 22 1996 08:209
	After a pad and carrier change both carriers had problems 
	with sticking to the pad and leaving wafers on the pad. 
	We replaced the black o-rings, screws and 1 check valve. The
	problem went away.
			
					B shift

	   

7.209PARTICLESSTRATA::WESTPHALWed Jul 24 1996 20:393
CMP.A2 - Down for scratches.  Suspect the conditioner is loosing diamonds.  A
	 different old disk was already installed and the tool was cleaned up. 
	 Ran out of time. The tool presently sets needing a pad.
7.210blankSUBPAC::LANDRYMon Jul 29 1996 19:5211
    Today, the system failed MQC's for uniformity.  The pad and the
    carriers were changed out to correct this problem.  MQC's passed
    after this activity was completed.
    
    Late in the shift, the left carrier was having problems picking up the
    wafer from the polish table.  The viton o-ring for the plate circum-
    ference and the three screws/o-rings were replaced.  Positive results
    followed the activity.
    
    						A shift
    
7.211bubblesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Aug 02 1996 06:436


            the pad was replaced due to a large bubble in the 
         middle.
                                         D shift
7.212broken wafer FABSIX::R_GEEMon Aug 19 1996 07:5713
	A product wafer slid out from under the left carrier during
 	the last part of a 130 sec PMD polish. 102 wafer on the pad 
	and 51 wafers on the carriers. 
	We checked the system out and found nothing wrong. The carriers
	had slight damage so we could still check the wafer exposure height. 
	The right carrier was set between .005 & .006, the left carrier was 
	set between .007 & .0085, to high. We believe this was the reason 
	for the fly out. Replaced the pad, carriers and conditioner disk.

					B shift

	   

7.213Dropping Wafers At Lift Off / Right Hub O-Ring ReplacedASDG::POIRIERSat Aug 24 1996 16:5310
    
    Right Spindle Leaving Wafers On Pad:
    -----------------------------------
    The right spindle has been repeatedly leaving wafers on the pad at 
    lift off today. This condition eventually ended in wafer breakage.
    The 2nd o-ring (counting from the center outward) on the right hub
    was replaced. The system is now running MQC's.
    
    
                                                    C Shift.
7.214two vacuum problemsSUBPAC::LANDRYSun Aug 25 1996 20:2910
    A couple of vacuum problems today......
    	
    *  The robot vacuum was inconsistant.  Replaced the small tubing thats
       attached to the end effector.  All set.
    
    *  The robot missed picking up a wafer from the right load/unload cup.
       Found the vacuum to the pedistal was releasing slower that the left.
       The system was rebooted and a dozen wafers were cycled.  All set.
    
    						A shift
7.215Left Spindle Vacuum ErrorsASDG::POIRIERFri Sep 06 1996 19:538
    
    Left Spindle Vacuum Trouble:
    ---------------------------
    The system experienced repeated vacuum errors on the left spindle.
    We're going to take a look at the left carrier & left hub o-rings
    shortly.
    
                                                             C Shift.
7.216STRATA::WESTPHALSat Sep 07 1996 20:2317
 
    
    
    
    
================================================================================

EQUIPMENT HIGHLIGHTS:

CMP.A1 - Down for poor uniformity.  Replaced the pad.  

CMP.A2 - Down for slipping wafers from the right carrier on every lift off.  
	 replaced the carrier plate O-ring and ran 6 wafers with no further 
	 problems. 


CMP.B2 - Weekly PM is in the incomplete note file.
7.217LOOSING WAFERSSTRATA::WESTPHALSat Sep 07 1996 20:246
CMP.A2 - Down for slipping wafers from the right carrier on every lift off.  
	 replaced the carrier plate O-ring and ran 6 wafers with no further 
	 problems. 


7.218Carrier sticking to the tableFABSIX::R_GEETue Sep 10 1996 08:402
    	The right carrier was sticking to the table. Replaced the
    	black o-ring and the backing plate screws, allset.
7.219new load cups bridge locationSTRATA::FINNTue Sep 10 1996 20:3810
	Installed new  left and right load station chucks. Retaught robot.
	Upon reboot bridge drove to left travel limit switch. Found both
	Home and Load software locations at 0. Retaught bridge home, load 
        and polish positions and realigned carrier to load stations and clean 
        station positions. Ran test wafers in HANDLING program

					Jim and A shift

	
7.220Load adjustments/Slurry linesFABSIX::J_SWENSONFri Sep 13 1996 15:117
    Adjusted left and right load buckets. Retaught robot for left and right
    load positions. Ran verification wafers. Replaced tygon slurry lines
    with Masterflex tubing. Checked slurry flow. Returned system to 
    production.
    
    
                                         Jim/C-shift 
7.221SYSTEM CRASH/LEAD SCREW LUBESTRATA::SWENSONFri Sep 13 1996 20:038
         System crashed, left carrier flew a wafer and made quite the mess.
         Cleaned tool. Replaced both carriers. Replaced pad and conditioned.
         Lubricated lead screw. Adjusted oil flow, marked off oil level. 
         25 wafers no touch were run, handling looked fine.
         Returned system back to production.

                       C-Shift

7.222re-constructed a carrierSUBPAC::LANDRYSat Sep 14 1996 18:267
    We had some problems today with a carrier.  The left carrier would NOT
    come up off the table when it's time was due.  Had to halt the process
    to get the spindle to raise.  Changed out the 3 o-rings on the plate
    retaining screws.  Also, the white o-ring was repositioned as well. 
    All is fine.........
    
    					Hawk  &  Doc show
7.223PARTICLESLUDWIG::WESTPHALWed Sep 18 1996 20:224
CMP.A2 - Down for failing MQCs for particles. Found the brush heights to be too
	 low.  Adjusted the same and cleaned the process area and conditioner.
	 Tool failed again for high particles.  Ran out of time.

7.224Brushes changedSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Sep 19 1996 08:035
    
       tool down at shift change for partticles.  Brad asked to 
    change the brushes.  Tool came up after the brushes were changed.
    
                                             D shift
7.225BRIBGE NOT OSCILLATING?STRATA::WESTPHALFri Sep 20 1996 20:114
    
CMP.A2 - FYI it was noticed that the bridge was not oscillating during polish.
	 Was a process change made?
7.226carriers stickingSTRATA::RIDLONEliminate the obviousSun Sep 22 1996 07:3214
    Carriers sticking to pad. Changed cond disk with no change. Checked
    carriers and replaced the 3 plate screws (orings shot) with no change. 
    Changed carriers but had trouble loading. The right one wouldn't load.
    Swapped carriers and now neither would load. Swapped them back and
    neither would load. Checked vacuum going to spindles and both seemed to
    be low (but now sure what they should be so it could have been ok).
    Removed the vacuum lines from solenoids 16 and 17 and vacuum was good
    out of solenoids. Reinstalled lines and carriers loaded. Good enough
    for now. Ran a couple of wafers through to check for carrier sticking
    to pad and that looked good as well. Over to production for mqc's.
    
    
    						B-shift
    
7.227ParticlesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Sep 26 1996 07:4918
 
         Tool down all night for high particles.  Tried the following:
  
         changed pad
         cleaned tool
         changed brushes
         checked brush heights
         changed and wire brushed conditioner disk
         checked recipes and programs
         checked slurry flow
         checked surfscan recipe
         checked particles with B2, B2 below a 1000, A2 ~4500
         
         latest move: changed pad with one from new lot...  awaiting
         MQC results.....
                                    D shift
         
7.228CMPA-TOOL particle updateSTRATA::WOOLDRIDGEPleasure, Spiked With PainThu Sep 26 1996 16:5878
     Strasbaugh particle problem update: 9/26
-Kyle

Today:

o  Repeated no touch recipe (eliminating IC pad) and particle counts were good.

o  Particle checks on both machines were in the 4000-6000 range

o  Because of low pad count on A1, pad was conditioned for an extended 
   time.  New MQC was run- no change.

o  Changed conditioning disk on A1- counts went up? - put old disk back on

o  Pad was scrubbed down with a brush and rinsed repeatedly on A1. No 
   change.

o  Noticed that the MQC wafers after polish looked like the polish pad?  
   This is the first time I have seen this.  The wafers show the pad 
   perforation marks perfectly.

o  Looked at pad expiration dates of the IC lots we have in the 
   drawer- one expired Aug. 27th and the other expires Oct 1st.
   (Rodel claims pads are good 6 months after expiration date)

o  Decided to change pad on A1 and install a pad with expiration date 
   of Jan 97'.  Conditioned & broke-in the pad with NEWPAD recipe.
   Ran MQC's and the counts were 4288 & 3323.  Interestingly, no pad 
   perforation marks were seen on these wafers however.
   Followed this by running a modified MQRINSE recipe: Defects went 
   down to 200.

o  Ran a modified MQC recipe with extended rinse at the end of the 
   recipe on CMP.A2.  DEFECTS WERE REDUCED TO 267 on one head and 245 on the 
   head. (no pre-measurements were taken so in fact the defects were 
   actually lower) Specifically changed was:
                                   Step 3: 20 seconds polish time
                                           40 rpm spindle speed
                                           40 rpm table speed
                                   
                                   Step 4:
                                           60 seconds polish time
                                           20 rpm spindle speed
                                           60 rpm table speed
                                  
The difference here is simply the addition of a step 4 rinse time with 
different spindle/table speeds & changing the final rinse step from 
10 seconds to 60 seconds. The recipe name is MQRINSE.  

o  John P. has cut a speedfam pad down to try on A2.  There is also 
   another pad from the batch now on A1 on top on the pad cabinet. 
   (good through Jan '97) if we want to look at that.

o  Speedfam tools defectivity OK?


     Musings:
              
- defects are hard to remove, but an extended rinse brings defects 
into a very low level.

-Pads may be partially the problem- but this is not fully 
understood since the new pad on A1 did not qual.

-It appears for what ever reason, the slurry is NOT rinsing off the 
wafers. This could be related to the pad/polishing OR to something 
going on with the slurry/delivery system.

-It appears we can run the process currently, but the process recipes 
will need to be modified to reflect steps 3 & 4 in the MQRINSE recipe.
A TDCO/edco will need to be put in place tonight (pss) in order to 
reflect this change until the process/hardware fix is fully 
understood.  This is simply an extension & addition to the rinse step.

If needed, I can be reached at home @ 464-5721

/Kyle
7.229broke wafersSTRATA::FINNMon Sep 30 1996 21:179
System crashed, operations heard alarm and found right head stuck to table and
left head raised with broken wafer on pad, cleaned system, replaced pad, 
installed one carrier and built and installed second carrier, ran pad cond, and 
breakins, MQC in progress.


				A shift
	
7.230crashedSTRATA::RIDLONEliminate the obliviousTue Oct 01 1996 05:278
    Tool crashed while breaking in. The breakin wafers in the fab are
    pretty thin. No carriers for this tool yet. Other priorities have
    prevented their assembly. Will try to get to but we have a fire drill
    today so doubt if we we'll have the time. This crash was not induced by
    the slurry problem we had tonight. 
    
    
    					B-shift
7.231Wafer BreakageSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed Oct 02 1996 07:4219
        Three instances in 45 minutes relating to vacuum.
     First, a wafer on the right carrier wasn't seeing
     good vacuum and needed a personal touch.  The next was
     the left carrier leaving a wafer on the table. Finally,
     the robot failed to pick up a wafer in the left send
     elevator.  During polish, the final two wafers were on the
     table and the wafer on the right spindle slid out.  I was
     looking at the shuttle area when it happened and I tried
     stopping the tool quickly by opening the polish door BUT
     SOMEONE USED NUTS TO DEFEAT THE MICROSWITCHES ON BOTH POLISH
     DOORS!!!!  Fortunately, Barry saw it and hit the EMO as I was
     in motion for my button.  The tool was cleaned up and new
     pad and carriers were installed.  Upon install, I found a 
     a small slice on the 1/4" vacuum line on the outside of the left
     spindle, repaired same....  Tool was turned over to OPS for
     break-in and MQC's.
                                    D shift
         
7.232NO WAFER ERROR/ SLIPPED WAFERSTRATA::WESTPHALWed Oct 02 1996 20:5511
CMP.A2 - Down for "NO WAFER AT LEFT UNLOAD" errors.  Ran 14 wafers using the 
	 handling recipe. Watched the chuck vacuum sensors and for any obvious
	 problems.  Didn't see anything wrong.  Ran 14 more wafers on a recipe 
	 using an abbreviated polish time.  On the last run it brokes a wafer.
         It came out of the left carrier during polish.  Cleaned up the process
	 area, conditioner, and load stations.  Changed all consumables
	 conditioned the pad and returned it to production for BREAKINS. 
	 Because no problems with the original error were found.  Engineering 
	 requested that the a dozen wafers be run and the tool observed.  If 
	 no problems occur, return it for MQCs.

7.233Removal rates misc. repairsSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Oct 03 1996 07:2031

         The tool was down when we came it and it was in the
      middle of VERIFY.  It failed three times for poor vacuum
      on the right carrier.  The right load cup rinse jet wasn't 
      aiming at the cup so the wafers were kind of dry at load 
      and had slurry after polish.  The robot was also having 
      problems trying to pick up wafers at SEND and LOAD.  Found
      the 1/8" line on the end effector hanging out and catching
      The jet was repositioned and 1/8" repositioned.  The tool
      then ran ~36 wafers for verify and had no problems.   The 
      tool failed MQC's for low removal rate.   After looking at 
      at the rate chart for A2 with OPS it shows a decrease trending
      since last week which indicated something was heading south.
      We checked the fairchild transducers and the left was coming
      in at 70psi.  It was tweaked up to 80psi.  The right was fine.
      The downforce was checked on both heads and both had force errors
      of ~4.  Downforce calibrations were done on both and brought 
      in at .41 on the left and -.1.  The left send elevator wouldn't
      home.  We pulled that out and found two things....  The drain
      for the dump solenoid on the RIGHT SEND fell off and the water was 
      flowing right onto the LEFT SEND elevator wiring connections, 
      repaired same.  We also found the home sensor for the LEFT SEND was
      pinched under the elevator and was shorting out.  These two 
      items are most likely the cause of the load problems we've seen 
      lately.  All this work and I don't think the removal problems
      have been corrected.  A new pad had been installed and is ready
      for break-in and MQC's.
                                                  D shift

        
7.234Wafer load problemsSUBPAC::TOOKERThu Oct 03 1996 17:0018
Today cmp.a2 went down due to wafer loading problems.  Wafers were not 
properly seating in either the right or left carrier.  After checking the
spindle to load station alignmnent, it was discovered that both spindles
were about 1/2 inch too far to the left of the load station.  After re-
homing the bridge and re-checking the spindle to load station alignment, 
the alignment was perfect.  After running the bridge back and forth several
times from the load station to the polish table, the bridge had again
drifted about a 1/2 inch to the left of the load station.  The cause of the
bridge drifting turned out to be an extremely loose bridge drive belt.
After tightening the drive belt, the bridge drifting went away.  The load
stations were also re-aligned and the hard down software variable was in-
creased from 1000 to 1200 counts to give the spindles a little for force
when loading.  As a part of troubleshooting this problem, it was necessary
to defeat both of the side door interlocks with tie wraps.  After completing
the work, the tie wraps were removed and the interlocks were once again
operational.

     Jim Tooker
7.235low polish rate/slurry statusLUDWIG::WOOLDRIDGEPleasure, Spiked With PainThu Oct 03 1996 17:4719
Low CMP polish rate- UPDATE:

     A new batch of slurry (lot #2h-16-96-1) was manually pumped to 
CMP.A2 due to the problems we are currently having with low rates.
The results were:
                 head 1-1300
                 head 2-1115
Directly after this, the slurry from the day tank was hooked back up 
and the results were as follows:
                head 1-1132
                head 2-1054
Based on the above, I have asked Ashland Chemical to dump the current 
day tank and replenish the slurry with the new batch.  We should have 
new slurry by 6pm or 7pm tonight.  The change over should be transparent
to operations.  Unfortunately, there is only enough slurry of this 
batch to tide us over through the weekend.  Then a new batch will have 
to come on line.
                   - Kyle   
7.236slipped waferSTRATA::WESTPHALFri Oct 04 1996 20:4410


CMP.A2 - Down for slipping a wafer.  Tool was cleaned up and consumables
	 replaced.  Tool was returned to production.





7.237crashSTRATA::WESTPHALSat Oct 05 1996 20:103
CMP.A2 - Down for crashing 2 product wafers.  Process area was cleaned up and
	 all consumables were changed.  Conditioner was cleaned.  Tool sets
	 needing only MQCs.
7.238slurry pmp 1 STRATA::RIDLONEliminate the obliviousSun Oct 06 1996 00:4110
    While running NEWPAD recipe noticed that the pad was kinda dry. My
    technician instincts kicking in I quickly halted the machine. Ahhhhh
    no crash....amazing considering that the pad was completely dry!!!
    Ran slurry flush and pump 1 wasn't rotating. Isolated problem to the DC
    motor control card in the electronics cabinet (P/N 06-12023-00, VP/N 
    106668). Replaced and tested/cal'd fine. INspected pad and no damage
    could be seen. Only 197 wafers on pad so it was not changed.
    
    
    						B-shift 
7.239CRASH.STRATA::RIDLONEliminate the obliviousMon Oct 07 1996 08:088
    CRASH, and just like that two product wafers turned to sand. Root cause
    for crash suspected to be slurry flow. Air bubbles found in supply line
    in rear vmb. CDU in slurry room pumping very slowly. Called ashland.
    Filters switched and flow came back to normal. 
    New pad, carriers, cond disk. Needs mqc's.
    
    
    					B-shift
7.240Furon valves orderedSUBPAC::LANDRYMon Oct 07 1996 20:3715
Per Engineerings request, the 3-way valve for the slurry distribution
was checked out.  We were checking to see if DI water was leaking by
the 3-way valve, hence diluting the slurry and/or supply.  This work
was done in leu of the low removal rate being seen on all the tools.
What we did find was that both 3-way Furon valves were leaking by.  
We tried re-building a valve, as we don't have any in stock.  The per-
formance was better, but still not good enough.

Jim T has ordered several of these valves.  They are being rushed in
for tomorrow.  The 1/4 turn slurry valve shut-offs to pumps 1 & 2 are
currently in the "off" position.  Please leave the valves off until we
install the new valves when they arrive.

      					A shift
  
7.241weekly and monthlySTRATA::RIDLONEliminate the obliviousTue Oct 08 1996 06:016
    Started the weekly and monthly pm's. Most of both have been completed. 
    The check sheets have been forwarded to Doc and Bob for A-shift
    completion.
    
    
    						B-shift
7.2423 way valves installedSUBPAC::LANDRYTue Oct 08 1996 20:3318
      The Furon valves for the slurry distribution/pump flush arrived
      today.  They were installed and tested with positive results.

      We also picked up where Mitch left off on the Monthly PM.  The
      PM is now completed.  

      New software was also installed today.  The current Rev is now 
      version 4.0.  After the software was installed, we came across
      a couple of problems.  We needed to change the state of some of 
      the calibration variables to correct.  There currently exists
      two problems which will need to be addressed.  The Send tank dump
      valves are not responding correctly all the time and the "dunk"
      performance of the send cassettes is off.  We are currently working
      the issue with Strasbaugh.  These problems do NOT necessitate that
      the system remain down.  A warning has been logged in RTC.

      						A shift

7.243right side high stddev.LUDWIG::RIDLONEliminate the obliviousWed Oct 09 1996 04:158
    Failed mqc's X5 for right carrier high stddev. After the first 2
    A-shift changed the right carrier. MQC's still came out high. B-shift
    swapped the left and right carriers. MQC still came out high. Checked
    the run out on the right carrier and it was great. Pulled the pad and
    it came off way too easily. New pad was installed and conditioned. 
    Ready for mqc's.
    
    						B-shift
7.244Send cassette water dunk featuer now operationalSUBPAC::TOOKERThu Oct 10 1996 16:487
Today the old revision of the Strasbaugh software (4.16.3) was reloaded onto
cmp.a1 & cmp.a2.  This software version will allow us to use the send cassette
water dunk feature.  Strasbaugh is implementing a fix to software version
4.20 that will give all the new features of 4.19 plus the send cassette water
dunk feature.  All warnings have been removed from cmp.a1 & cmp.a2.

Jim T.
7.245HIGH PARTICLE / POOR UNIFORMITYLUDWIG::WESTPHALSat Oct 12 1996 20:597
CMP.A2 - Down for high particles and poor uniformity.  For uniformity changed
	 the pad.  The high particles were on the wafer that came off the right 
         head.  The right scrub brush intermittently does not turn unless you 
	 physically help it.  The motor appears to be bad.  Ordered the part
	 from stock but could never get a hold of anyone.  Tool presently sets 
	 in wet idle still needing break ins and MQCs.

7.246brush motorLUDWIG::RIDLONEliminate the obliviousSun Oct 13 1996 03:116
    Removed the brushes and ran both in diagnostics. The right was actually
    fine and the left one bad. REplaced the left one and installed/adjusted
    the brushes. Waiting on MQC's.
    
    
    							B-shift
7.247passedLUDWIG::RIDLONEliminate the obliviousSun Oct 13 1996 22:012
    MQC's passed. The left remained good at less than 500 and the right
    came down to ~1000.
7.248robot crashLUDWIG::RIDLONEliminate the obliviousMon Oct 14 1996 08:146
    Robot crashed into a wafer on the right send. Cleaned up mess. cycled
    some wafers with the handling recipe. MQC's were ran using the handling
    recipe, which passed. 
    
    
    					B-shift
7.249TDCO on lineSTRATA::WOOLDRIDGEPleasure, Spiked With PainTue Oct 15 1996 16:523
    I put a TDCO on line for an added rinse step to reduce particle levels. 
    Assuming this works out ok, I'll spec it.
                                              -Kyle
7.250Dropping wafersFABSIX::O_WHITFORDWed Oct 16 1996 23:555
    Right carrier dropping wafers on lift-off after polish cycle. Noted pad
    count approaching max, replaced pad and both left & right carriers, ran
    10 pad condition cycles and turned over for MQC's.
    
    Dave W.
7.251Carrier staying on table, wafer not blowing offSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Oct 17 1996 07:178
    
         During pad break-in, the left carrier kept staying
      on the table as well as the wafers weren't blowing off 
      into the load cup.  Took apart carrier and found both 
      valves installed backwards.  Tool finished break-in
      without any problems and needs full MQC's.
     
                                       BJ
7.252WAFER BREAKAGESUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Oct 18 1996 05:068
    
           Tool crashed two product wafers.  Problems found
        were the right carrier had exposure out of range and
        the pad was practically floating on the SUBA pad.  See
        note 9.11 for more details...
    
                                    D shift
    
7.253PVA BRUSH FELL OFFLUDWIG::WESTPHALFri Oct 18 1996 17:192
CMP.A2 - Down for the PVA brush falling off.  Ordered a new one from stock and
	 installed it.
7.254air motor replacedFABSIX::J_SADINFreedom isn&#039;t free.Sun Oct 20 1996 11:547
    
    
    	cmp.a2 - replaced right air brush motor (old one squeaking loudly).
    Checked brush height, ok.
    
    
    
7.255uniformitySTRATA::RIDLONEliminate the obliviousTue Oct 22 1996 06:267
    AFter b3 and a1 crashed we decided to increase step 3 of the mqc recipe
    by 5 seconds. Particles passed (barely) but uniformity failed. Pad
    count at 386 so it was changed. It will be conditioned for breakins by
    7:30. 
    
    
    						B-shift
7.256Wafer breakageSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Oct 24 1996 08:126
           
    
           The tool crashed to no reason.  Tool was cleaned up and
       the weekly PM was completed.
    
                                                  D shift
7.257A2 CrashFABSIX::J_SWENSONThu Oct 24 1996 19:3116
     system crashed taking out both carriers.
    
    1. cleaned tool and replaced pad and suba
    
    2. replaced carriers 1 and 2
    
    3. ashland chemical called and said that they had a float sensor 
       problem which in turn shut the slurry flow off. ( phone call came in
       5 min after A2 crash, see Earls crash report.
    
    4. Ashland has requested time to investigate this problem further.
       ie. try to recreate error, But due to excessive work load (32 lots)
       request was denied.
    
    
                                         C-shift
7.258Robot loading problemsSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Oct 25 1996 05:4910

           The robot kept having loading problems. Several homing
       sequences and a power down couldn't correct it.  Leon tweaked
       the sensor switch and it ran fine after that.


                                                  D shift

        
7.259Wafer Crash - No Slurry FlowingASDG::POIRIERFri Oct 25 1996 20:1410
    
    Wafer Crash:
    -----------
    The system broke 2 product wafers at 6:30 tonight. Ashland called
    to say there was no slurry flowing when we heard the machine groan.
    Earl ran to the tool but he was a few seconds too late. This is
    the same senario as yesterdays crash on A2. See Earl's wafer crash
    report for more details.
    
                                                               C Shift.
7.260No vacuum on left carrierSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Oct 26 1996 04:5213



            Installed new carriers after C shift crash.  New 
         carriers installed but the left failed for vacuum.
         After opening up the carrier, no o-ring was found inside.
         Installed o-ring and the tool passed MQC's and is now up and 
         running.

                                                  D shift

        
7.261Crash updateLUDWIG::RIDLONEliminate the obliviousTue Oct 29 1996 05:438
    Had a couple of crashes tonight. The weekly was performed 
    after the first crash. Both crashes were dummies (with
    plenty of oxide). The first was unexplainable, the second had 
    some robot problems, which were taken care of, but when the wafers got
    to the table it hadn't conditioned in awhile so there probably wasn't
    enough slurry (the pad was kept wet though).
    
    						B-shift
7.262POOR UNIFORMITYSTRATA::WESTPHALThu Oct 31 1996 19:385
CMP.A2 - Down for poor uniformity.  There were 230 wafers on the pad so it was
	 changed.  The conditioner was changed out due several spots showing 
	 wear.  The process area was cleaned up.  Tool passed MQCs and was
	 returned to production.

7.263Slurry delivery and bubblesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Nov 01 1996 06:0013

              The tool was given to us around 03:30am for the
          slurry change over to barrel feed.  The tool was 
          converted over and slurry flow was adjusted in software.
  
              The 3-way slurry valves were replaced as well.

              When OPS went to use the tool, several bubbles had
          formed so we had to change the pad.  The tool is turned
          back of to OPS for MQC's.

                                          D shift
7.264Robot Vac Switch / Left Spindle Vac Error / D.I.- SLurry ValvesASDG::POIRIERFri Nov 01 1996 19:3130
    
    Robot pick-up problems:
    ----------------------
    The robot vacuum switch was recalibrated this morning due to pick-up
    problems. Jim T. increased the end effector blow off pressure slightly.
    Jim thinks the puddle of D.I. water on top of the robot may be getting
    drawn into the end effector tubing. This condition might be the cause
    of the vacuum switch calibration drift we've been seeing.
    
    
    Left Spindle Vacuum Errors:
    --------------------------
    The systen had numerous errors for "Left Spindle Vacuum Not Found".
    The Barksdale pressure switch on the left spindle was adjusted to
    13 in/hg. No Spindle Vac Errors for the last 4 hours.
    
    
    Slurry/ D.I. 3 Way Valve Problems:
    ---------------------------------
    The Furon valves on the back of the tool were getting a burst of air
    during every conditioning cyle. This condition was causing D.I. water
    to get drawn into the house slurry system.  Jim T. called Strasbaugh
    and was told to remove the exhaust lines from the manifold in the
    pnuematic cabinet. After the lines were removed, the D.I. back flow
    problem ceased. Thhe Strasbaughs can now be connected back to the house 
    slurry system.
    
    
                                                                C Shift.
    
7.265Barksdale ain't the problemSUBPAC::LANDRYMon Nov 04 1996 19:358
    While performing the Monthly pm today, we installed a new Barksdale
    pressure switch on the left spindle assembly.  Unfortunetly, the new
    switch is responding the same way as the previous one.  And yes, we
    did setup the switch per the Strasbaugh manual procedure.  The system
    is still running under a warning statu for this issue.
    
    						A shift 
    
7.266robotLUDWIG::RIDLONEliminate the obliviousTue Nov 05 1996 04:245
    Robot started missing wafers in the left send cassette. Ended up having
    to recalibrate the left send 0 position to correct problem. 
    
    
    					B-shift
7.267Right carrier leaving wafers on pad at lift-offFABSIX::R_POIRIERTue Nov 05 1996 06:5710
	We had a problem with the right carrier. It was leaving
	wafers on the pad at lift off. We took the carrier apart. 
	Replaced check valves, O-ring, and carrier screws. 
	No problems since then.



					B Shift

7.268Left & Right Spindle Rotory Unions ReplacedASDG::POIRIERSat Nov 09 1996 19:0825
    
    Left Spindle Vacuum Problems:
    ----------------------------
    The system has been intermittently erroring for various Left Spindle 
    vacuum errors all week. This morning the Bridge deservoed during an
    auto cycle for a "no vacuum found on left spindle" error.
    
    1) The left & right spindle vacuum was checked at the hubs. (24-26"/hg)
    
    2) The Barksdale Vacuum switches in the Bridge were checked and
       readjusted. The left alarmed at 12"/ hg and the right at 8"/ hg. 
       Both were set to 15"/hg per the Strasbaugh manual.
    
    3) With the vacuum ON and reading 26"/hg at the hub, a slight lifting of
       the vacuum supply tube in the Bridge caused the vacuum to drop to
       12"/hg on the left spindle and 0"/hg on the right spindle. 
       At this point we concluded that both spindle rotory unions were BAD. 
       Both rotory unions were replaced and 4 new o-rings were installed on
       both the left and right hubs. At the present time, we're mounting
       the hubs and starting the hub alignment procedure. "B" shift will
       be finishing up the hub alignments, installing a new pad and
       turning the tool over to Mfg for MQC's.
     
                                                                C Shift.
      
7.269Hub AlignmentsFABSIX::R_POIRIERSun Nov 10 1996 21:016
    On Sat. nite we did the Hub alignments for both spindles, after
    the Rotary Union changes were done on C shift. Turned tool back
    over to OPS.
    
    						B shift
    
7.270Vacuum/FABSIX::R_POIRIERMon Nov 11 1996 04:0810
    We were having two problems. Right carrier was leaving wafers on the
    pad at lift off, and "left/right spindle off" errors. 
    We found that the rotary union for the right spindle needed to be
    tightened. This took care of the wafer vacuum problems.
    Did slight adjustments to the Barksdale switches and this got rid
    of the spindle off errors.
    
    
    						B Shift
    
7.271more on hooverLUDWIG::RIDLONEliminate the obliviousMon Nov 11 1996 07:068
    Just an fyi to let everyone know that we put a vacuum gauge in line on
    both the right and left spindles. The "spindle vacuum off" alarms were
    happening at cycle start and we saw the vacuum go as low as 8 inHg when
    the load chucks checked for wafers but then it would go right back down
    to 24-25 range. The vacuum gauge is still in line on the left spindle
    for your viewing pleasure. 
    
    						B-shift
7.272TUB OVERFLOW ERRORSTRATA::WESTPHALWed Nov 13 1996 19:404
    CMP.A2 - Down for "Tub drain overflow" error.  The tub overflow sensor
             sensitivity got out of adjustment.  Adjusted the same and
    	     tested several times.  returned the tool to production.
    
7.273Polish Tub Water Level Alarms.ASDG::POIRIERThu Nov 14 1996 19:2811
    
    Polish Table Tub High Level Alarms:
    ----------------------------------
    The tool is intermittently alarming for the Polish Tub water level.
    The level sensor needs a sensitivity adjustment. We've been letting it
    alarm because MFG needs the tool and the tool is running GREAT at this
    time. "D" shift, please readjust this sensor at the first available 
    opportunity.
    
    
                                                            C Shift.
7.274Pad changeSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Nov 15 1996 06:207

            Tool failed for deviation and range, replaced pad.



                                          D shift
7.275Polish Tub Level Errors / Remote Vacuum PumpASDG::POIRIERFri Nov 15 1996 19:2718
                                                                   
    Polish Tub level Errors:
    -----------------------
    The system errored numerous times today for Polish Tub Level errors.
    Even though it errored repeatedly, it was less frequent than
    yesterday. The Polish Tub level sensor still needs a little more
    adjustment. We didn't adjust it today due to tool availability. 
    
    
    Remote vacuum cabinet pump:
    --------------------------
    The vacuum pump in the remote vac cabinet is blowing water out of the
    exhaust muffler. There may be a problem with the float switches in the 
    water collection tank in the cabinet. We need to verify the top and
    bottom float switches are working correctly at the first available
    opportunity.
    
                                                           C Shift. 
7.276Pad Replacement Due To Poor UniformityASDG::POIRIERSat Nov 16 1996 19:2310
    
    Poor Uniformity:
    ---------------
    A Pad change is currently taking place due to poor uniformity.
    The typical (salad bowl) profile has appeared once again. 
    Current status is bgn_repair.
    
    
                                                      C Shift. 
                                                         
7.277Furon's replacedFABSIX::M_RIDLONEliminate the obliviousTue Nov 19 1996 06:255
    Replaced the Furon slurry/di valves with 2 Burkett valves per pump.
    Calibrated slurry flows.
    
    				
    						B-shift w/ AQ
7.278Wafer Detected On Robot ErrorsASDG::POIRIERThu Nov 21 1996 19:379
    "Wafer Detected On Robot" errors:
    --------------------------------
    The system experienced numerous errors for "wafer detected on robot".
    This condition also knocked the tool out of auto cycle. This problem
    hasn't occured during the last four hours. If this problem persists
    into D Shift, check the robot vacuum tube.
    
                                                           C Shift.
    
7.279Robot issuesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Nov 23 1996 05:159
    

          The robot had been removed and inspected on C shift and
       re-installed.  when we came in, we checked all of the robot
       positions for verification, replaced pad, brushes and 
       carriers.  Tool didn't have any problems with the robot
       during break-ins, MQC's or product runs all morning.
     
                                          D shift
7.280info for 11/22 end effector vac. problemFABSIX::J_SWENSONSat Nov 23 1996 18:4216
    
    
     Checked stockroom for spare furon valve none in stock
    
     Rushed in valve, recieved 11/23
     
     Cleaned out valve as best as posible (44)
     
     Reassembled robot and valve
    
     Set end effector vac. switch.
    
     If the intermittent vac. problem comes back there is a new 
     Solenoid Valve in the Strausbaugh roll around
    
                              J.T./C-Shift
7.281Wafer breakage....SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Nov 30 1996 06:1410
          Tool crashed two product wafers in segment #3.
       The crash was caused from a wafer sliding out of
       the left carrier.  The only thing we found was the 
       left carrier expossure was at -4 to -7.  The uniformity
       was ok previous to the crash.  The crash form was posted
       and tool cleaned up.  MQC's will be starting shortly...


                                          D shift
7.282slurry loss detect installedFABSIX::B_FINNTue Dec 03 1996 19:257

	Slurry loss sensor install for "slurry #1 loss detect", on I/O 
input 29. It will inhibit the start of a cycle if slurry # 1 loss is detected,
and generate an "end cycle" if loss is detected during a run. 

					A shift
7.283Pad change due to MQC's results.SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Dec 05 1996 04:3810
    
    
          Product coming off tool with wacked out readings,
      8000A on the left with 900A on the right.  PSS had them
      run full MQC'S.  The tool was in spec for removal but
      out on uniformity, particles and scratches.  All in all,
      there was a problem with the lot but the tool needed TLC
      anyways due to the MQC results.  
    
                                          D shift
7.284Pad changeSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Dec 07 1996 01:207
               A pad change was needed due to failing MQC's
          for bad uniformity.  There was actually a bubble 
          around 12" around the center. 


                                          D shift
7.285POOR UNIFORMITYFABSIX::B_WESTPHALWed Dec 11 1996 19:224
CMP.A2 - Down for poor uniformity.  The pad and conditioner were changed and
	 the tool was cleaned up and returned to production.

7.286Robot crashSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Dec 12 1996 06:1913
    
         The robot went to unload the 2nd wafer of the two that it was
       placing into the receive elevator and didn't release it.  It didn't
       alarm and went to pick up the next set of wafers from the send
       elevator and crashed into the next wafer that was suppose to have
       been picked up.  The tool did not error for having a wafer on the
       robot before it loaded more wafers nor after the crash.  The robot 
       kept trying to drive to the elevator.  The tool was EMO'ed and 
       recovered without any problems.  Wafers were run with the NOTOUCH 
       recipe and the problem did not repeat.                    
     
                                       D shift
       
7.287Vacuum problemsFABSIX::J_SWENSONThu Dec 12 1996 19:2313
    
    	Replaced solenoid 44
    
        Replaced vacuum line to robot.
    
        Set vacuum switch.
    
        System as of 0715 in verify. Tool running dummy wafers. to verify
        fix.
    
    
                                     C-shift and J.Tooker
                                                         
7.288Uniformity, near-crash, vacuum errorSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Dec 13 1996 06:5517
               A pad change was needed due to failing MQC's
          for bad uniformity.  

               The tool then needed another pad change because of
          scratches and uniformity.  

               The tool finally passed MQC's and was running product 
          when a wafer slid out of the right carrier and the tool was
          EMO'ed before crashing.  The right carrier was replaced and
          MQC's needed to be run again.  While running MQC's, there
          were vacuum errors regarding the new carrier but it ended 
          up being the o0ring in the hub.  After the o-ring was
          replaced, MQC's were started over again.


                                          D shift
7.289Right Spindle Hub (GAP) Adjusted.ASDG::POIRIERFri Dec 13 1996 19:2310
    
    * Right Spindle Hub (GAP) Adjusted:
      --------------------------------
      The gap was measured at .020" on the tight side to .100" on
      the wide side. With the base of the hub swept in at .>.001",
      the new gap is .042" on the tight side and .100" on the wide side.
    
    
                                                         C Shift.
    
7.290Broken Wafers / Wafers Sticking To Pad - Suspect PadsASDG::POIRIERSat Dec 14 1996 19:3525
    
    Broken Wafers / Pads Suspected :
    ------------------------------
    The system broke 2 product wafers this morning. The Pad, Carriers &
    conditioning disk were changed. We stopped using the pads in the cmp
    room (3 different lots) and found pads with different lot #'s in the
    stock room. A pad with a different lot # was used as areplacement.
    
    The polish table controller amperage spiked to 48 amps during the down
    force ramp just prior to the crash. This unusually high amperage spike
    occured after 106 wafers were run on the pad. This supports our feeling
    that something is wrong with the pads.
    
    
    
    7:00 update:
    -----------
    The system just went down again for wafers sticking to the table after
    60 wafers were run on the pad. This tool had new carriers and a new
    conditioning disk installed this morning. We believe this sticking
    condition supports the BAD PAD theory.
    
                                                           C Shift.
      
    
7.291Broken wafer, carrier problemFABSIX::R_GEESun Dec 15 1996 19:3713
	A breakin wafer was left on the pad by the right carrier The 
	system never errored for lose of vacuum. The conditioner 
	crushed the wafer during the pad conditioning cycle. 
	Found a pinched line inside the carrier. Cleaned the system,
	performed the weekly p.m., replaced the pad and conditioner.
	Reconfigured the vacuum lines and turned the barbed fitting
	that is fixed to the top plate back to its normal 90% from 
	center position.

					A shift

	   

7.292ROBOT MIS-LOADSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed Dec 18 1996 21:2813

            The robot apparently went in to pick up a product 
        wafer from the right send elevator and the wafer went 
        between the end effector and its base.  No damage visible.  
        Ran (6) wafers at slots #1, #2, #10, #11, #24 and #25
        under NOTOUCH with no problems.  PSS will have evaluate 
        the wafers on hold from OPS.

                                  good to go...


                                          D shift
7.293Robot slamage....SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Dec 19 1996 05:1114

            The robot apparently went in to pick up a product 
        wafer from the right send elevator and the wafer went 
        between the end effector and its base.  No damage visible.  
        Ran (6) wafers at slots #1, #2, #10, #11, #24 and #25
        under NOTOUCH with no problems.  The two product wafers
        in question visibly looked ok but were send along with 
        remarks per PPS.

                                  good to go...


                                          D shift
7.294Right Carrier ProblemFABSIX::J_SWENSONThu Dec 19 1996 19:2512
    Right Carrier Problem
    
    - Carrier leaving wafer on pad
    
    - Carrier not unloading wafer in unload station
    
    - Replaced carrier
    
    - Returned back to production with no further problems
    
    
                           C-Shift
7.295Pad change - Bad unif.SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Dec 21 1996 00:508
         Tool failed MQC's for uniformity at 292 wafers.  A pad
      change was done and the tool was returned to OPS for break-in
      and MQC's....

  

                                          D shift
7.296Robot errors / ruptured tubing.ASDG::POIRIERFri Dec 27 1996 19:2515
    
  Robot errors:
  ------------
    The system errored for "no wafer found on robot". The 1/8' vacuum tubing 
    leading to the robot end effector was ruptured. The tubing
    was replaced and the vacuum switch was recalibrated.          
    
    The system went down again shortly thereafter for the same error. We
    discoverd tha the end effector disk was not seating correctly in the end
    effector after each wafer pick up. A slight adjustment in the 
    tubing length took care of the problem. A 50 wafer verification run
    has been started.  The system status is now in VERIFY.                                              
    
                                                                C Shift
                                                                    
7.297pad changeFABSIX::O_WHITFORDQuticherwhinin&#039;Fri Dec 27 1996 21:104
    Replaced pad as it had dried out extensively during robot
    troubleshooting. Conditioined and turned over for verification.
    
                                        Dave Whitford/D shift 
7.298Wafer dropped out of left carrier.FABSIX::R_POIRIERMon Dec 30 1996 07:333
    Dropped wafer out of Left carrier during polish.
    Replaced right carrier also, due to high count (380).
    
7.299tank filling problemFABSIX::J_SWENSONSat Jan 11 1997 18:1614
    
       Tanks not filling quickly
    _________________________________
    
     Di pressure 40
    
     cda 100 psi
    
     reset tool, tanks filled still slow 
     di line acting as if there is air within the line.
    
     returned tool for production.

                                  C-shift
7.300mqc failureFABSIX::M_RIDLONEliminate the obliviousMon Jan 13 1997 07:285
    Failed mqc's. Did the weekly w/ a pad and carrier change. NOTE: the
    conditioner disk had very little diamonds left on it. 
    
    
    					B-shift
7.301pad changeFABSIX::O_WHITFORDQuticherwhinin&#039;Sat Jan 18 1997 00:024
    Replace pad at 106 wafer count, LARGE bubble had developed since I
    installed this pad last night.... bummer.....		
    
                           Dave  
7.302Wafer Breakage / 2 Product Wafers ScrappedASDG::POIRIERSat Jan 18 1997 19:0745
     
    Wafer Breakage:
    --------------
    Two product wafers were lost during a crash this morning. The wafers
    broke dring the polish step. We were unable to determine the cause of
    th fly out. The pad was on the tool overnight with only 34 wafer run on
    it. Both carrier films were severly damaged making it impossible to
    measure the step height. (See crash report below)
    
    The pad, carriers & conditioning disk were replaced. The tool was
    sucessfully MQC'd and ran fine for the rest of the shift.  
     
  ---------------------------------------------------------------------------
    
 CMP.A1 & CMP.A2 WAFER CRASH FORM (to be filled out by MFG/EEG/PSS, together)

 1) System __CMP.A2_________
 2) Recipe ___PMD________
 3) Step # __2___
 4) Time into step __NA____
 5) Type of wafers that broke _PRODUCT_________
 6) Carrier which wafer slid out of __LEFT____
 7) Other carriers that were damaged __RIGHT______
 8) Alarms displayed by polisher _______NO VACUUM FOUND ON LEFT SPINDLE________
				 ______________________________________________
				 ______________________________________________
 9) Did any other polisher error for slurry or crash NO__, systems____________
10) # of wafers on pad _34___
11) # of wafers on carrier __216____
12) Was there any vibration prior to the crash __NO____
13) Was there plenty of slurry (for step 1&2) or water (step 3) present on
    the table __YES_______
14) Check step height of the carrier from question #6 (CARRIERS ALL CHEWED UP) 
                                                      
15) Please note any other unusual circumstances:
                               



16) Put this check sheet in CMP note #9

    
                                                             C Shift.
                            
7.303wafer detectLUDWIG::FINNMon Jan 20 1997 19:187
	Installed power, common and signal wires from new wafer detect I/O, 
        via conduit and flex track to Bridge Terminal strip # 20.
        Much more difficult than it sounds. Weekly pm completed.
	
				A shift
	
7.304pad changeFABSIX::M_RIDLONEliminate the obliviousTue Jan 21 1997 02:184
    Replaced pad due to mqc failure.
    
    
    					B-shift
7.305pad change and robot malfunctionFABSIX::J_SADINFreedom isn&#039;t free.Tue Jan 21 1997 16:5113
    
    
    	replaced pad again due to mqc failure. Went back to suba + machine
    stacked. Mqcs passed, but only after 24 breakin wafers. Also, machine
    errored for "left spindle vacuum not found" and then proceded to slam
    the robot arm in and out of the left load cassette (arm was sliding in
    between wafers and not sensing the wafer was present). This condition
    always seems to accompany the "left spindle vacuum not found" alarm
    (the alarm stops in just a few seconds without any operator
    intervention). Weird software glitch. Aborted process and reset left load
    elevator. Everything was ok from then on.
    
    jim
7.306"Wafer Not Found On Robot " - ErrorsASDG::POIRIERWed Jan 22 1997 19:2710
    
    Robot End Effector:
    ------------------
    "Wafer Not Found On Robot " Errors.
     The 1/8" vacumm tubing leading to the End Effector Puck was too
     tight. This prevented the Puck from seating in the End Effector.
     The tubing slack was adjusted and the vacuum errors ceased.
    
                                                           C Shift.
    
7.307ROBOTFABSIX::J_SWENSONThu Jan 23 1997 19:2811
    Wafer not found on Robot
    ------------------------
    
     Checked vac. switch cal looked good
     checked blow off for end effector pressure 20 psi o.k.
     checked end effector for defects o.k.
     ran 25 wafers with handling recipe no errors occured.
     returned to manufacturing.
    
    
                             C-Shift
7.308POOR UNIFORMITYFABSIX::B_WESTPHALFri Jan 24 1997 19:394
CMP.A2 - Down for failing MQCs for poor uniformity. Pad was changed.  
    	 The tool is still at verify.  Needs more breakin wafers.
    
7.309crash FABSIX::M_RIDLONEliminate the obliviousMon Jan 27 1997 05:5416
    Tool crashed with no apparent nor obvious reason. The wafer came out
    of the right carrier. Earlier in the evening the right carrier was
    sticking to the pad. The black oring and the platen screws were
    replaced but the problem persisted. The carriers were swapped which
    improved the release but the right still lagged by about a second. 
    Not sure why the right would be sticking to the pad with a good
    carrier. The cal/config numbers were checked and compared to A1 with
    nothing out of the ordinary found. To recover from the crash the
    carriers were flushed (autocycle, f8, f5) due to the right carrier
    sucking up wafer debris. And then the carriers, pad, and diamond disk
    were replaced. The disk was just replaced this AM by A-shift and
    already had signs of considerable wear (only 80 wafers on  pad).
    Weekly pm was going to be performed but A-shift beat us to it. 
    
    
    					B-shift
7.310wafer breakage investigation resultsSUBPAC::LANDRYMon Jan 27 1997 19:2538
    
The system was taken down this morning to investigate some wafer breakage
problems that have occurred over the last week or two.  Most of the wafer
breaks have been caused by the right carrier/spindle assembly.  Many items 
were checked out:

*  The left pressure gauge on top of the bridge was found to be rubbing
   the inner top of the system as the bridge moves.  This gauge displays
   the back pressure and vacuum read out.  The gauge was rotated to 
   alleviate this occurrance.  Prior to rotating, the atmospheric readout 
   numbers were off of the 14.7 target (the left ~15.0, the right ~14.4)

*  The back pressure PSIA readout was calibrated, final readouts achieved
   were 14.7 psia for both.  This was followed by the E-P regulator 
   calibration for back pressure.  The left side was about 1 psia low on 
   the high end.  Performed calibration on both sides.

*  Next, the spindle alignment was checked on both spindles.  Both were
   in spec, with NO adjustments made.

*  Both hubs were checked.  Both were fine, with NO adjustments made.

*  Down force calibration was then performed on both spindles.  The right
   side was real tight (.09% force error).  The left spindle was found to
   be at a force error of ~3.5%.  Calibrated the left side.

*  Slurry calibration was checked.  No problems found.

*  Load and Cleaning station alignments were performed.  The left side was
   fine.  However, the right load station was inconsistant in sitting in
   the ring properly.  Therefore, we made the mechanical adjustment to
   allow free movement in all directions as the carrier picked up a wafer.

*  A new pad was installed at the end of the above checks.  The HANDLING
   recipe was used to exercise the mechanical hand offs of the system.
   All ran fine.  Breakin wafers were then successfully run.  MQC passed.

                                      A shift and Jim T
7.311robotFABSIX::M_RIDLONEliminate the obliviousMon Jan 27 1997 20:509
    Tool put down for robot problems. The only thing that I could see wrong
    while testing the robot was that the wafer appeared to be dragging on
    the lower arm of the robot. Made a minor adjustment to the end effector
    to correct and tested the load/unload utilizing the same recipe. No
    other problems were seen. 
    
    
    
    					B-shift
7.312exLUDWIG::FINNWed Jan 29 1997 19:447
    
    
    ADDED CR-40 relay for wafer detect mod.
    
    
    			A shift
    
7.313Wafer breakage - Product....SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Jan 30 1997 08:479
    
    
           A product wafer slid out of the left carrier at lift-off
       a hit the right carrier which had not popped up yet.  There
       were some small bubbles noticed upon removal.  The tool was
       cleaned up and turned back over to OPS.  A crash report was
       posted in #9 with the lot/tool info.
    
                                            D shift
7.314Another one bites the dust!SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Jan 31 1997 05:0917
         
         The tool crashed two monitor wafers on the table and it
      was so bad, both wafers were turned into stone dust and the
      left carrier fell off the tool????  The tool was cleaned up
      and new pad and carriers and pad were installed.  The tool
      passed MQC's without any problems.

         After the robot dropped a product wafer off at the left 
      receive station, surface friction caused the wafer to come 
      back out with the robot.  When the robot went to swing around 
      to the load cassette, CRASH! SLAM! BAM!  No more wafer...  
    
          The robot was re-homed and the mess cleaned up the best it
      it could.  VERIFY wafers were run without any issues.  The
      tool was brought back up.

                                             D shift
7.315Pad changeSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Feb 01 1997 06:274
    
         Pad change was done due to bad unif. on MQC's.
    
                              D-shift
7.316Robot & Left Spindle Vacuum FailuresASDG::POIRIERSat Feb 01 1997 19:1635
    
    Vacuum Errors On Robot & Major Vacuum Loss On Left Spindle:
    ----------------------------------------------------------
    
    The Robot experienced vacuum loss at the end of "D" shift this 
    morning. This caused wafer breakage during handling. The Robot
    has been experiencing intermittent vacuum failures for the
    last few days. In addition, and more importantly, there has
    been two incidents in the last two days of THE LEFT CARRIER
    FALLING OFF THE HUB. (MAJOR VACUUM FAILURE)
    
    The left spindle rotory union appears to be the cause of the vacuum
    failure. It is intermittently leaking vacuum as it rotates. This
    lowers the vacuum at the hub and robot causing numerous vacuum 
    problems in the tool. 
    
    Both left and right spindle rotory unions were replaced in NOV 96.
    See note 7.268
    There is one union in stock at the present time. The Union o-ring does
    not come with the union. A SID sheet has been sent to Materials for
    the Strasbaugh supplied o-rings. We have located some that are close to
    the proper size. We will hand them off to "B" shift tonight. If they
    won't fit, call Jim Tooker at home. 
    
    The o-ring type is: 2-020 N70
    
    The o-ring size is: .864 ID  X .070 Width 
    
    
    NOTE: Remember to replace the small o-rings on the rotory union
          feed tube as well. We have plenty of these sized o-rings
          an will hand them off to "B" Shift.
    
                                                              C Shift.
    
7.317orings and monthlyFABSIX::M_RIDLONEliminate the obliviousSun Feb 02 1997 03:587
    Installed new o-rings in both rotary unions. Could not cause vacuum
    to fail as mentioned by C-shift in prior note. Monthly pm was
    completed. New pad installed and conditioned and tool given back to
    production.
    
    
    					B-shift
7.318unions leakFABSIX::M_RIDLONEliminate the obliviousTue Feb 04 1997 07:2911
    INtermittent alarms for left spindle vacuum not found. Checked rotary
    unions by pulling up on hose connections and vacuum dropped and caused
    said alarm (left for left, right for right). One union obtained from
    stock and one from Jim T.'s office. The new union went on the right.
    The one from Jim's office has a different kind of nut which has holes
    in the side preventing a vacuum seal. We tried the old union with new
    and improved orings but it was a no go. When tested it failed while the
    right one passed. This is over to A-shift as our time is about up. 
                                                          
    
    						B-shift
7.319Vacuum errorsFABSIX::B_FINNTue Feb 04 1997 19:2322
Continued rotary union repairs from B shift. Removed left rotary 
union,  removed large oring from base of rotary union reducer. This 
does not provide a seal. The two small orings below the threads 
provide the top seal. Measured male and female threads depth and 
length and confirmed this oring is not being compressed. Removed inner 
tube and found the lower oring damaged. This oring was split in two 
via its circumference so that we had two intact orings, neither 
providing a seal. Reinstalled orings and tube, installed new rotary union. 
Installed both left and right hubs, set both gaps and aligned with dial 
indicator.  

Ran 36 wafers with handling program and received a vacuum error. This 
error conincided with a loud whistle from the Vacuum pump cabinet.
This error has not reoccurred and seems VERY intermittent.
We have swapped sol 1 with sol 2 per eng request. None in stock, eng is 
has ordered 2. Due in fedx p1. Just completed 50 wafer verify run with 
handling program. Pad has been condition, but not yet broken in.
System remains in verify. 


					A shift
7.320valves installedFABSIX::J_SWENSONFri Feb 07 1997 19:478
    Vacuum valves installed in vac. cabinet
    A 50 wafer verification run will need to be run on the handling
    recipe before product can be run.
    
    Will pass on to D-Shift
    
    
                           C-Shift
7.321Leaving wafers on tableFABSIX::R_POIRIERMon Feb 10 1997 07:1417
	
	We had a problem with A2. Right carrier was leaving wafer
	on the pad at lift-off. I replaced both carriers because
	the wafer count was high (386). Started running break-ins,
	and after the 4th set of wafers, the left carrier left a
	wafer on the table. 

	The tool didn't alarm and the conditioning arm came over
	and broke the wafer. Replaced the left carrier, the pad
	and conditioning disk. Did weekly PM. 
    
    	Tool is in break-in/MQC's.      


					B Shift
        
7.322wafer breakageSUBPAC::LANDRYMon Feb 10 1997 19:255
    The system is currently in a pad change after a wafer breakage.  See
    breakage report for details.
    
                                          A shift
    
7.323CMP.A2 problems...loooong storyFABBIT::J_SADINFreedom isn&#039;t free.Tue Feb 11 1997 19:4136
    
    
    	Well, it's been a rough ride on this doggy today. Started off the
    day with both carriers sticking to pad for an extended amount of time
    after polish. It was necessary to stop all axis' to get carriers off
    the table. Initially changed the conditioner and performed an extended
    condition...no change in behavior (carriers still sticking). Checked
    carriers and found tubings bent over. Repaired carriers and tried
    again. Carriers were now lifting off! Ran 16wfrs and turned back over
    to manufacturing. The very first run resulted in a crash (wafers slid
    out of left carrier. left carrier stuck to pad without the wafer).
    Right carrier popped up ok. Screen alarm for this was "axis error"
    which refers to carrier rotation failure. This could be explained by
    the fact that the vacuum was on at the carrier (minus the wafer), hence
    the vacuum was trying to stop the table and carrier rotation. 
    
    	Cleaned up the mess, built two new carriers (installed on tool now)
    and tried again. First run we lost a wafer out of the left carrier as
    soon as it touched down on the pad (EMO'd tool before wafer crashed).
    We were running the NEWPAD recipe. Swapped carriers and ran wafers only
    to have the wafer slide out of the left spindle AGAIN (at touchdown).
    Doc swapped the two vacuum lines back to their proper positions (these
    are the lines that were switched last night). Retried running wafers
    and noticed that the backpressure on the left spindle was up at 67psia
    despite the fact that the program was not calling for it (program was
    calling for 14.8psia in step 2 of the NEWPAD recipe. right spindle was
    ok). 
    
    	B-shift will need to switch anything they changed last night back
    to the way it was previously. Vacuum lines are already done. 
    
    	Leaving the system as is. Will verbally pass off to B-shift. 
    
    
    	jim/doc/brendan 
               
7.324communicatingFABSIX::M_RIDLONEliminate the obliviousTue Feb 11 1997 23:4221
    Sorry for the confusion folks. Heads been up the ole poop shoot of
    late. Last night we swapped the vacuum lines (82 & 88, 24 & 26) and the
    backpressure control lines (c98 & c99) in the hopes of getting more
    info on the "wafer-being-left-on-table" problem we've been
    experiencing. In a sense we did get some more info. If you filter
    through the problems that were seen today the important data is that
    the wafer flyouts and being-left-on-the-pad problems switched to the
    left spindle. We swapped the vac lines and control lines back to normal
    and then just switched the left and right backpressure EP regulators.
    We'll monitor this and if the wafers continue to fly out of the left
    spindle than the EP regulator is at fault. NOTE: product will not be
    subjected in this short test, just breakins (new pad). One other thing
    to note, could be important, the BKPR readings in idle were reading
    14.5 to 14.6. All the recipes have Backpressure set to 14.8 (when not
    0.0) which is going to put 2 to 3 tenths of psia onto the wafer.
    Calibrated number 66 and 68 (left and right back pressure offset's) so
    that the idle reading is 14.8 psia.
    
    
    
    						B-shift
7.325more communicatingFABSIX::M_RIDLONEliminate the obliviousWed Feb 12 1997 05:488
    No fly outs or wafers left on the pad with the backpressure regulators
    swapped. We swapped them back and calibrated them using the procedure
    in the manual. Ran another boat of wafers using the PMD recipe and
    again we didn't have any problems or errors. Gave it to production to
    run MQC's and they passed on the first try. 
    
    
    						B-shift
7.326CMP.A2 passdownFABSIX::J_SADINFreedom isn&#039;t free.Wed Feb 12 1997 19:3419

    Lots of issues on this machine today. An operator put the tool down for
    having severe rem rate range after mqcs and on product (the tool was
    removing 1000� more on the right spindle than the left). EP regulator
    and bp valves fittings were checked and most all were found to be
    leaking.  All fittings were retaped and reinstalled (no leaks).
    Downforce cal was performed and the right spindle was found to be
    fluctuating ~10lbs at the higher cal pressures (left calibrated ok with
    no fluctuation). Determined downforce EP regulator to be bad and
    replaced the old style regulator with the new style. EP regulator has
    been installed and calibrated. 

        ONLY RIGHT SPINDLE DOWNFORCE NEEDS TO BE DONE! Left is ok and the
spreadsheet has been entered into Excel. 

        a-shift                    


7.327lost 100psi on the back pressure regulatorsSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Feb 13 1997 06:3022
          Picked up from A shift and completed the downforce
      on the right spindle.  It was fine...

          Started up the VERIFY and the tool crapped out after
      the 1st set of wafers.  The tool wasn't blowing off the 
      wafers into the load cups.  Tried looking into spindle flush
      to confirm there is a problem and the tool wouldn't flush
      either.  According to the schematics, there is suppose to 
      be constant 100 psi to the back pressure regulators in the
      pneumatic cabinet.  The 100 psi was missing.  We tried tracing
      the line but we lost it when it goes towards the electrical
      cabinet.  The line is suppose to be #1 but we have #114. 
      Both tools are the same way.  We spoke with Todd about it
      he seemed to thing it was a pinched line. We jumped the #114
      line and brought in another 100 psi.  The tool now ran fine.
      The line is still by-passed but was ok to run.  The riggers
      are suppose to be here at 8am for the upgrade.

 

                                             D shift
7.328Gearbox ReplacedASDG::POIRIERThu Feb 13 1997 19:3323
    
    Gearbox Replacement:
    -------------------
    The gearbox has been replaced and the table reinstalled. 
    
    Action plan for tonight:
    -----------------------
    1) Todd from Speedfam will stay here tonight until the install is
       complete.
    
    2) After final alignments are complete, put a new pad on table 
       and run 25 wafers using the BREAKIN recipe.
    
    3) After the Breakin recipe is run, turn the tool over to manuf 
       for MQC's.
    
                          
       Also,  Jim Tooker wants the filters in the Bridge changed.
    
       If any major problems develop, call Jim Tooker at home.
    
    
                                                          C Shift.
7.329Gearbox upgrade statusSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Feb 14 1997 05:2215


            The table is skewed left to right and the Rep.
        decided that the table will have to be pulled again
        in the morning.  None of the items that were listed
        in the C-shift passdown were able to get done.  Jim
        Tooker was notified and work will commence around
        8am. 

            Both 50 micron spindle filters were replaced.

            Area cleaned up of debris.
 
                                 D shift
7.330Can't get the Bridge to Table sweep alignment in.ASDG::POIRIERFri Feb 14 1997 19:3518
    
    Gearbox Upgrade Status:
    ----------------------
    We struggled all shift trying to get the sweep reading in. The Table
    was pulled again and the holes on the gearbox mounting bracket were 
    enlarged. The Table was reinstalled and the sweep reading is still
    .085". The Strasbaugh rep is planning to work into D Shift.
    
    We don't know if the sweep reading was in spec before the Table was 
    removed yesterday. The Bridge to Table sweep reading is something
    Strasbaugh started checking last year. I remember that this measurement
    was not taken when the tools were installed 3 years ago. Maybe we're
    chasing a phantom problem?  CMP.A2 is the first time the rep has
    encountered this problem in over 20 gearbox conversions. He is trying
    to decide how to proceed at the present time.  Passing on down to "D"
    shift. 
    
                                                        C Shift.
7.331A2 upgrade?SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Feb 15 1997 07:049
    
            Rep. left at 9pm...  left a note for C shift.
    
       We didn't even get a change change to see what was 
    needed for completion but from what the note on the tool 
    reads and everything being still in pieces, who know.  We
    don't even know if the Rep. is coming back.  
    
                                      
7.332Gearbox install updateASDG::POIRIERSat Feb 15 1997 19:4132
    
    Gearbox Install Update:
    ----------------------
    The Strasbaugh rep left Friday night at 9:00. He gave up on trying to get 
    the table sweep alignment in. He left instructions to perform our standard 
    table/spindle/hub method of alignment. We didn't have time to finish
    the alignments due to other priorities in the area. Here's the current
    status:
    
    1) Table alignment: currently at .0012"  OK
    
    2) Left Spindle alignment: currently at .001"   OK
    
    3) Right Spindle alignment: Currently at .020"  Spec .001
    
    4) Left & Right Hub alignments: Not Done        Spec .001
    
    Todd from Strasbaugh can be reached via his pager in Springfield Mass
    at 1-800-459-6922.
    
    In addition to the alignments, there are a few pipes & hoses that
    connect the chilled water to the water jacket not yet installed.
    Also, there are hoses and exhaust ducts that are not yet installed.
    
    One other thing, the mounting hardware for the slurry delivery tube 
    has been lost. It consistes of a nut and two washers. Take a look at
    A1 to see what sizes are needed. 
    
    
                                                             C Shift.
    
    
7.333Show me the moneyFABSIX::M_RIDLONEliminate the obliviousSun Feb 16 1997 07:3815
    To make a really long story short....the table run out alignment equals
    .001". Both the left and right spindles were checked and since we did
    some messin around with the table they needed some adjusting, not to
    worry though as they are both .001". All hoses and nuts and bolts and
    watchyamacallits that go underneath have been installed. The left and
    right hub run out alignments need to be performed as well as all panels
    reinstalled. The slurry delivery also needs to be reinstalled. The
    chiller was turned on and flow maxed out at 3.9gpm to check for leaks
    with none being found. We can see why the rep got a bit frustrated with
    this pig. Would have been nice if the table sweep alignment was done
    before the table was removed. Its a good thing that alignments not in
    the spec.
    
    
    					B-shift
7.334CMP.A2 hub runouts etc...FABSIX::J_SADINFreedom isn&#039;t free.Sun Feb 16 1997 19:448
    
    
    	Did hub-runouts and brought within .001. Replaced
    pad/carriers/conditoner disk and ran 25 wafers on NEWPAD recipe without
    incident. Turned over to production and machine passed mqcs.
    
    
    	a-shift
7.335RR stands for Reading RailroadFABSIX::M_RIDLONEliminate the obliviousTue Feb 18 1997 06:4910
    Failed mqc's. Left RR was 2200, right RR was 1100. Since we had the
    reverse situation last week we went ahead and rechecked the EP
    regulator cal to rule out drifting and redid the down force cal. Both
    EP's were 10psi low at 3140 and at 580 both read 9psi. We cal'd the
    high end to 80 psi following the book but left the low to 9psi. The
    down force cal for the right didn't need any adjustments but the left
    needed to be tweaked. It wasn't out by much at least not enough to
    account for the differenc in RR.    
    
    						B-shift
7.336table alignmentsFABSIX::B_FINNTue Feb 18 1997 19:0633
	System failed MQC's. Decision was made at the post morning 
meeting to correct the table tilt condition, (which has been at 64 mils 
low on load side), since the completion of the table's gear box. Before 
loosening any bolts we check the tilt by sweeping the table with a dial 
indicator attached to the spindle. We the rotated the the table 90' 
and sweeped again until we had measured the table in 360's. The tilt was 
continuously ~-64 to -68 mils.  Tim D. tried to contact Strasbaugh in 
N.Y.  We contacted Ca. and got the Eastcoast F.S. Sup. phone and 
pager. Tim spoke with him and explained the situation. They 
(Strasbaugh) have no one to send at this time.   

We tried to slightly loosen the gearbox bolts and jack up the load 
side of table, we improved to approximately -45. We tried loosening the 
the white exhaust box to pick up some play, but NFG. We 
tried installing a second jack on the far right side of the table to wedge the 
high end down whilst still raising the front we improved to -35. Base 
of drip pan began to bow as we jacked.  We installed boards and shims 
under the system and steel plates under the bottom jack to prevent this. We 
loosened the front jacking support bar and retried the above procedure. We 
improved to -27.  Unfortunately from this point on ...the whole table, 
both sides began to rise. 

Table center eye bolt for table removal could not be found in the room or 
Jim T's office.

All jacks have been removed for safety at passdown. Current sweep ,(bolts loose)
is -73 mils (hey we worked all day to get this!). 

There does not seem to be enough adjustment room.

The 6 bolts to the gear box are still loose.
Table rotation motor cables are still unplugged 
7.337Table statusSUBPAC::BJUBINVILLECMP Equipment EngineeringWed Feb 19 1997 05:5811
         how the heck does that thing come out?????  

         We got three of the bolts out for Tim but the
      the right top one and two front ones are still in.  
      We felt it was necessary to leave it in the condition 
      it is now rather than have hundreds of pounds of metal sitting
      on a scissors jack.  The locking bolts and supports
      are still in place.  
 
                                 D shift
7.338Gearbox Update / Mounting Screws Turned DownASDG::POIRIERWed Feb 19 1997 19:3318
    
    Gearbox Replacement Update:
    --------------------------
    The gearbox mounting screws were removed and turned down on the lathe.
    This allowed for greater left to right movement of the gearbox. The 
    bridge to table sweep alignment is now .0005". However, when the 
    table is rotated 90 deg, the sweep measurment increases to .009".
    
    Somehow we need to align the bridge to table while (rotating) the table 
    and sweeping the table (linear). We're much closer than we were
    yesterday, it appears that turning down the mounting screws worked.
    Passing on to "D" shift to finish the alignments and reassemble all the 
    remaining hoses and duct.
    
    
                                                             C Shift.
    
     
7.339Table, sweep and spindle alignmentsSUBPAC::BJUBINVILLECMP Equipment EngineeringThu Feb 20 1997 05:5823

      Pre-readings showed table sweep .001 on one direction
   and .009 at 90 degrees.  Table run-out was .050????  We
   wanted to get a baseline before we cranked some weight on
   the table from the spindles.  We put ~600 lbs from each 
   carrier at three locations to see if the table would
   actually settle.  It didn't move!  First we brought the
   table run-out in to .0005.  We got the sweep at both cross
   sections to .001 and .002.  We re-checked the table run-out
   and still had .0005.  We applied the 600lbs downforce again
   at three locations to see if the readings would change and
   they didn't.  The spindles were checked.  The left was out
   at .03 and brought in to .002.  The right spindle was at
   .05 and that was brought in to .001.  Both hubs were at
   .003 and brought in to .0005 each.  The tool was cleaned 
   up and all of the covers, plates and hoses were re-connected.  
   We ended replacing both carriers due to the right carrier 
   having vacuum problems and the left for sticking to the pad.  
   The first round of MQC's had R.R. at 1100, and 1300 with unif. 
   at 7% and 11%.  Things are looking good.
 
                                 D shift
7.340Unif. issues GONE!!!FABSIX::B_JUBINVILLEFri Feb 21 1997 06:1611
    
           At first change, the MQC results showed 8% on the left
       and 12% on the right.  We wanted to prove it was the carrier 
       and not the tool so we swapped the carriers from left to right
       and 'voila!' the left came in at 9.2% and the right at 7.6%.
       Just a precaution, a 2nd round of MQC's were run and tool
       passed with the same results.  Tool came back up to OPS.
    
                                             Good to go...  
    
                                                   D shift
7.341lousy conditionerFABSIX::M_RIDLONEliminate the obliviousWed Feb 26 1997 01:374
    Failed mqc's twice. Pad was changed and given back.
    
    
    					B-shift
7.342carrier oringsLUDWIG::FINNMon Mar 10 1997 19:195
    
    
    Right Carrier sticking to the pad. Orings changed.
    
    					A shift
7.343robotFABSIX::R_GEETue Mar 11 1997 19:288
	The robot was erroring for no wafer on robot whenever the was an
	empty slot between wafers. Adjusted cal. variable #118 from .88
	to .82 ran verification wafers on both elevators, all set.

					A shift

	   

7.344POOR UNIFORMITYFABSIX::B_WESTPHALSat Mar 15 1997 18:205
CMP.A2 - Down for failing MQCs for poor uniformity.  Cleaned up the process
	 area.  Replaced the conditioner and pad.  Conditioned the pad 10 times
	 and returned the tool for MQCs.  

7.345Carrier Vacuum ProblemsLUDWIG::STANDINGThu Mar 27 1997 14:2816
 CMP.A2    Vacuum System.  The left carrier had problems with maintaining
           carrier vacuum.  Swapped carriers left to right and the problem
           moved with the carrier.  So it was decided to replace the O-rings
           and the check valves in the carrier and reinstall.  After doing this
           and turning on the vacuum pump we noticed water dripping from the
           pump cabinet.  Found the water to be coming from the pump exhaust
           muffler.  Upon closer inspection we found the regulator to the pump
           cabinet reading zero lbs. this was due to the regulator being shut
           off.  Then checked the blown down tank and it was full of water,
           drained into a bucket.  Then checked the vacuum tank and found it to
           be completely full of water!  Well with the air shut off to the blow
           down tank it never drained, so the vacuum tank never drained either. 
           This could account for the vacuum problems.  Next we checked the
           vacuum pump for CMP.A1 and found the air to the regulator was shut
           off as well!  Both tanks were full of water.  Drained same. 
           Returned both tools to production.
7.346wafer crash a2/dummy wafers no product lostFABSIX::J_SWENSONThu Mar 27 1997 19:2724
    Wafer Crash 
   -------------
    
   - Send ahead wafer slid out of right carrier causing crash
    
   - Manufacturing noticed wafer had slid out and tool was EMOED
    
   - Noticing pads are having long breakin periods (40-50 wafers)
    
   - This will be the third pad in 24hrs for this tool.
    
   - Cleaning up tool, replacing pad,carriers and conditioning disk
    
   - Will be passing on to D-Shift for MQC's
    
                     FYI these are the lot #'s of the 2 previous pads
    
    
               - 6FN-JP2
    
               - 6HO-JP9
    
    
                                             C-Shift
7.347Carrier replaced - bad uniformitySUBPAC::BJUBINVILLECMP Equipment EngineeringFri Mar 28 1997 05:246
           The tool failed for uniformity on the left carrier
        which was replaced.  The tool finally passed MQC's after 
        that.

                                 D shift
7.3484" Disk/BP Cond'n testsFABSIX::B_POLLARDThu Apr 03 1997 18:5319
The Conditioning tests are continuing from day shift.  I ran 2 MQC's and 
a total of about 25 wafers on the MQCBP recipe.  For tonight I need 
to a MQC done when the current boat of dummies finishes (Eric will do the 
MQC).  Then I need 100 to 150 more dummy wafers cycled before the tool
is given back to production and then 1 FINAL MQC on the MQCBP recipe using
the EMQC wafers next to the tool.  Please print out the maps (if PSS can 
help with the final MQC that would be appreciated).  NOTE - the wafer counter
was NOT reset with this PAD so it should be over 250 by the END of the last
batch of 100 dummies.

When the Tool is given back over to Production - the following will be 
needed:
1. Save the cross-section of the PAD for Bryan.
2. EE will put the STD 6" disk back on along with correcting the offsets
   that were changed for the 4" disk.
3. A new pad will be needed with standard MQC's.

Bryan Pollard
7.3494" COnditioning Disk Experiment Ongoing.ASDG::POIRIERThu Apr 03 1997 19:1915
    
    4" Conditioning Disk Experiment Ongoing - Offset Coeff's Changed:
    ----------------------------------------------------------------
    Bryan Pollard is currently running an experiment with the 4"
    conditioning disk. This evaluation will run into "D" Shift
    tonight. When the experiment is completed, set the following
    offset coefficients to their origional settings:
    
    #126 Conditioner Offset Coefficient    96,500   (currently 94K)
    
    #71  Conditioner Dead Weight Coeff     Set Same as CMP.A1. 
                                           
    
    
                                                               C Shift.
7.350ENG work completedSUBPAC::BJUBINVILLECMP Equipment EngineeringFri Apr 04 1997 07:149
          The tool finished its ENG work.  The pad was removed 
      and a piece was cut out for Bryan.  Replaced a shredded o-ring   
      in the left hub.  Installed a new pad.  Re-installed 6"
      conditioner disk and changed back the variables.  Pad was
      broken-in and is BGN_MQC, waiting for the results.


                                 D shift
7.351CMP.A2 vacuum pump replacementFABSIX::J_SADINFreedom isn&#039;t free.Sun Apr 06 1997 20:3112
    
    
    	Left carrier fell off spindle during carrier clean step. Found some
    o-rings not seated properly and fixed those. Carrier fell off again and
    it was noticed that the vacuum pump was not reaching optimal vacuum
    (running constantly). Isolating the vacuum pump from the system made no
    change in the pumps behavior and it was therefore decided that the pump
    was bad. Replaced the old pump with one of the new ones in CMP and all
    is well! System is pumping down quickly and holding good vacuum. 
    
    
    	a-shift
7.352valves replacedFABSIX::B_FINNMon Apr 07 1997 20:2814
	System crashed in left receive station, cleaned up and ran test wafers
	found end effect very close of to of wafer while withdrawing on R 
        stroke. intermittently dragging wafers out of cassette and 
        tipping cassette over. Tried several cassette. Wafer was dragging on 
        the bottom of slots while being placed in cassette.
        Adjusted robot/elevator position and ran test wafers aok.

	And in other Unrelated news:

	Replaced both left and right backpressure/vacuum regulators and Valves 
        45 and 46, at EE's request. System is in its 24 hour MQC.  

					A shift
7.353Pad changeSUBPAC::BJUBINVILLECMP Equipment EngineeringThu Apr 10 1997 07:324
    
       Pad change due to MQC's  - bad unif. 
    
                                  D shift
7.354Scratches-uniformity.FABSIX::L_SHKOLNIKOVFri Apr 11 1997 07:377
            The tool failed MQC's for poor uniformity and scratches on the
        left carrier. The tool has been cleaned and left brush replaced.
        Tool passed MQC'c and released to manufacturing.
    
    
                                                     D-shift.
    
7.355System leaving wafers on pad at lift off.ASDG::POIRIERFri Apr 11 1997 20:259
    
    System Dropping Wafers On Pad:
    -----------------------------
    The tool has been leaving wafers on the pad this afternoon. The o-rings
    in both carriers have been replaced. We just ran 6 wafers with no
    problems. 
    
                                                        C Shift.
    
7.356robot bluesSUBPAC::LANDRYTue Apr 22 1997 20:307
    The robot was rather annoying this morning.  a couple of times, it
    errored for "robot not making left send".  The rails were cleaned (a
    bit too much grease was found).  Cleaned the lead screw as well.  It
    also have "fragments" near the end stroke of the travel.  It has been
    performing admirably since.
    
                                  A shift
7.357Conditioning Tests for TonightSTRATA::POLLARDThu Apr 24 1997 17:5420
Please continue the following Conditioning Tests on CMP.A2
** NOTE the Left Send Cass is having problems so unless EE gets it fixed
   only use the Right Send CAss.

1. Finish Current MQCBP run (Cassette has 4 dummies then 2 mon in slots 5,6
   which of course are in reverse order after).
   - Scrub wafers & Measure tx 
   - Pre readings on Filmthick.A1 ProcEng MQC prog with labels of 
     BP Cond'n Test3 L2nd (corresponds to SL2 after) & R2nd (corresponds to SL1)
   - be sure to save Delta measurements on tool
2. Run 25 dummies thru MQCBP recipe.
3. Do std MQC run & eval - save Delta results.
4. Cycle 75 more dummies thur MQCBP recipe.
5. Do std MQC run & eval - save Delta results. 
6. EE do Pad change & reinstall std 6" disk 
   EE also change  Var 126 "Prog Cond Offset" back to 96,500.
7. Do std startup procedures for New Pad.

Thanks,
Bryan P.
7.358SAVE the Pad - Please!!!STRATA::POLLARDThu Apr 24 1997 17:5822
Please continue the following Conditioning Tests on CMP.A2
** NOTE the Left Send Cass is having problems so unless EE gets it fixed
   only use the Right Send CAss.

1. Finish Current MQCBP run (Cassette has 4 dummies then 2 mon in slots 5,6
   which of course are in reverse order after).
   - Scrub wafers & Measure tx 
   - Pre readings on Filmthick.A1 ProcEng MQC prog with labels of 
     BP Cond'n Test3 L2nd (corresponds to SL2 after) & R2nd (corresponds to SL1)
   - be sure to save Delta measurements on tool
2. Run 25 dummies thru MQCBP recipe.
3. Do std MQC run & eval - save Delta results.
4. Cycle 75 more dummies thur MQCBP recipe.
5. Do std MQC run & eval - save Delta results. 
6. EE do Pad change & reinstall std 6" disk 
 ****PLEASE TO SURE TO SAVE A FULL CROSS-SECTION OF THE PAD ******
   EE also change  Var 126 "Prog Cond Offset" back to 96,500.
7. Do std startup procedures for New Pad.

Thanks,
Bryan P.

7.359Left send Elevator Needs Adj.FABSIX::J_SWENSONThu Apr 24 1997 19:5816
    Robot/Left Send Elevator
    ------------------------
    
    
    - Robot would not pick up wafers in left send elevator.
    
    - Moved cassette to right elevator robot picked up wafers and ran fine.
    
    - Tool has a warning for left send elevator and is up to ENG.
    
    - Looks like cal variable 80 will need to be adjusted.
    
    - Will pass on to D-shift
    
    
                              C-Shift
7.360left send elevator works again.FABSIX::L_SHKOLNIKOVFri Apr 25 1997 08:185
    Replaced the pad after engineering tests. Rebooted the tool and run 24
    wafers with notouch recipe to verify left send elevator. Did not have 
    any errors. Released tool to manufacturing.
    
                                                     D - shift.
7.361robot crashFABSIX::J_SWENSONFri Apr 25 1997 20:2624
    Robot crash
    ----------
    
    
    - robot crashed into right send elevator 
    
    - checked vac switch and re calibrated
    
    - checked end effector vac solenoid looked good
    
    - manually excercized robot and found alignments to be o.k.
    
    - found no root cause to the crash.
    
    - the original crash cassette is in the right send elevator
    
    - tool remains down in comets
    
    - these crashes are intermittent, we may be running worn cassettes
      it might be time to weed out cassettes in which we are having
      problems.
    
    
                        c-shift
7.362No crash.FABSIX::L_SHKOLNIKOVSat Apr 26 1997 08:306
    
    Regarding recent wafer crash in right send elevator we checked 
    robot and elevators alignments,removed and cleaned the robot arm
    and replaced vacuum line between robot and end effector.Verification
    run of 24 wafers has been completed without any errors. tool released to
    production.
7.363PAD ChangeLUDWIG::STANDINGSat Apr 26 1997 20:171
CMP.A2    Down for a pad change.
7.364robot arm problemsFABSIX::J_SADINFreedom isn&#039;t free.Tue Apr 29 1997 17:1911
    
    
    	Wafers were being dragged back out of both receive cassettes on
    CMP.A2. Found two allen screws on the back of the robot arm were loose and
    tightened them. Also adjusted robot top position to .63 (from .64)
    since the arm still appeared a bit close to the wafers on retraction.
    Ran verification wafers fine. Back to production.
    
    
    	a-shift
    
7.365Bad Uniformity - pad changeSUBPAC::BJUBINVILLECMP Equipment EngineeringWed Apr 30 1997 07:085
    
         tool failed MQC's for 20%+ unif.  Replaced pad, conditioner
       disk and tool came back in.  
    
                                         D shift
7.366Failed MQC'sFABSIX::R_POIRIERMon May 05 1997 07:018
    
    		Failed 24 hr MQC's. Replaced pad, conditioner
    		and left carrier (380 wafer count). Awaiting
    		MQc results.
    
    
    						B Shift
                                                  
7.367Engineering ExperimentFABSIX::R_POIRIERTue May 06 1997 07:3911
    

			Took tool at 5:00am and began engineering 
			experiment for Brian Pollard.  Put new pad
			and 4 inch conditioning disk. Ran special
			breakin recipie and special MQC recipie.


					B Shift

7.368Cond'n Tests for TonightLUDWIG::POLLARDTue May 06 1997 18:3231
We're rounding up the FINAL Cond'n Marathon on CMP.A2 - hopefully this will
help us Double the Pad LIFE.

PSS/OPs - Please continue the following Conditioning Tests on CMP.A2:

1. Cycle 50 Dummies (ECMP-XRSZ) thru recipe "XRSZBP2"
2. Do a  MQC run (wafers ECMP-MQC-1 - Need Pre Tx Readings) 
   using recipe MQCBP2.
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run as an MQC with Particles also.
   - Put in NOTes the Time of MQC and # Wafers on Pad.
3. Cycle 50 Dummies (ECMP-XRSZ) thru recipe "XRSZBP2"
4. Do a  MQC run (wafers ECMP-MQC-1 - Need Pre Tx Readings) 
   using recipe MQCBP2.
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run as an MQC with Particles also.
   - Put in NOTes the Time of MQC and # Wafers on Pad.

** Repeat the Above Steps Until the System has 500 wafers or the uniformity
   is > 15%.

5. If the system is needed back to Prod - the following will be done:
   EE do Pad change & reinstall std 6" disk 
 ****PLEASE TO SURE TO SAVE A FULL CROSS-SECTION OF THE PAD ******
   EE also change  Var 126 "Prog Cond Offset" back to 96,500.
6. Do std startup procedures for New Pad.

Thanks,
Bryan P.


7.369mqc test for brian pollardFABSIX::R_ANDERSONTue May 06 1997 20:351
    mqc wafer run at 7:30 p.m. wafer count on pad 160.  a-shift.
7.370Save the PAD! Condn & 18% tests on CMP.A2STRATA::POLLARDWed May 07 1997 19:2750
** MOST Important - SAVE THE PAD from the CMP.A2 Cond'n Tests!!!!!!!

Well the Marathon for the New Con'n recipe has gone very well.  We've had
425 wafers and the last uniformity's were 9.4% & 6.3%.  So I wanted to go
all the way to 500.  Then we're going to be converting to SC1 18% slurry from
a barrel early in the morning so the system will be at UP/ENG all night.

Here's the detailed plan:

1. Finish the current 24 dummies thru recipe "XRSZBP2"
2. Do a  MQC run (wafers ECMP-MQC - Need Pre Tx Readings) 
   using recipe MQCBP2.
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run (LVNE) as an MQC with Particles also.
   - Put in NOTes the Time of MQC and # Wafers on Pad.
3. Cycle 50 Dummies (ECMP-XRSZ) thru recipe "XRSZBP2"
4. Do a  final MQC run (wafers ECMP-MQC - Need Pre Tx Readings) 
   using recipe MQCBP2.
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run (LVNE) as an MQC with Particles also.
   - Put in NOTes the Time of MQC and # Wafers on Pad.
5. EE will REMOVE current Pad and SAVE IT Please (put label on it "Bryan P")!
 ** NOTE - Leave SETUP for 4" DISK as I will do New Slurry Tests with 
           the 4" Disk and appropriate Offsets.
6. EE will get 18% barrels of SC1 slurry outside Ashland's office.
7. EE will install slurry and calibrate flows.  Use small pump to keep
   slurry in barrell mixed.
8. EE will do 20 Cond'n Cycles with the Recipe - "NWPadBP3"
   (Note the Default # of Cond'n cycles is now 2).
9. Ops/PSS will do 20 Breakins with recipe "NwPadBP3".
10. Ops/PSS will do an MQC run with recipe "MQCBP2".
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run (LVNE) as an MQC with Particles also.
11. If uniformity is > 10% then do 12 more Bkins on "NwPadBP3" then go to
    #12 - otherwise just go to #12 with no additional Bkins.
12. Ops/PSS will do an MQC run with recipe "MQCBP2".
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run (LVNE) as an MQC with Particles also.
13. Ops/PSS will do an MQC run with recipe "MQCBP2".
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run (LVNE) as an MQC with Particles also.

I will continue testing to Optimize for particles/scratches in the morning.

NOTE this is very urgent as Lot JL1509 is on HOLD until I get an 18% process
in place to split it.

Thanks,
Bryan

7.37118% slurrySTRATA::FINNWed May 07 1997 21:185
	Ashland has mixed 100 gals of 18% slurry and it is outside their office
for tomorrows eng test.   (two 50 gal containers)

				A shift
7.372Update to Bryans gameplan. YIELD::SHONGThu May 08 1997 08:44140
Here's the detailed plan:

1. Finish the current 24 dummies thru recipe "XRSZBP2"
________________________________________________________________________________
Bryan, This portion of the test was done on A-Shift. I'm not sure who did it 
but when we came in they were running part 2.
________________________________________________________________________________

2. Do a  MQC run (wafers ECMP-MQC - Need Pre Tx Readings) using recipe MQCBP2.
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run (LVNE) as an MQC with Particles also.
   - Put in NOTes the Time of MQC and # Wafers on Pad.
________________________________________________________________________________
The thickness readings were taken but the Pre's were done on Thickness.A1(A-
shift). I'm not sure why they did that but they did. Removal rate and Std.Dev
was L=1376/10.4 and R=1187/6.6. Particle results were: 

	Left	Right
	----	-----
  7)	80	93
Sum)	835	840
Scr)	0mm	0mm

Results have been entered in Workstream. I've also placed a note in the notes-
file. The MQC's were completed at ~21:25. The count on the pad was 451 wafers 
after the MQC's. 

________________________________________________________________________________

3. Cycle 50 Dummies (ECMP-XRSZ) thru recipe "XRSZBP2"
________________________________________________________________________________
Done. That was completed at ~22:45. 
________________________________________________________________________________

4. Do a  final MQC run (wafers ECMP-MQC - Need Pre Tx Readings) 
   using recipe MQCBP2.
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run (LVNE) as an MQC with Particles also.
   - Put in NOTes the Time of MQC and # Wafers on Pad.
________________________________________________________________________________
This run was started at ~23:00. The PRE and POST readings for this MQC run 
was done on Thickness.A4. The PRE is called "PRE MQC 500 WAFERS" the post is 
called "POST MQCBP2 PART4". Thickness and Std.Dev readings were L=1253/11.2 and
R=1130/5.5. Particle results were::

	Left	Right
	----	-----
  7)	48	102
Sum)	389	653
Scr)	1.69mm	0mm

The file numbers for this run in the UV1050 are (Pre->L=2814, R=2815)(Post->
L=2816, R=2817). 

Pad count after these two MQC wafers were run was 503. I've just talked to EE 
and there going to go get the new Slurry drum. The time is 23:45.
________________________________________________________________________________

5. EE will REMOVE current Pad and SAVE IT Please (put label on it "Bryan P")!
 ** NOTE - Leave SETUP for 4" DISK as I will do New Slurry Tests with 
           the 4" Disk and appropriate Offsets.
6. EE will get 18% barrels of SC1 slurry outside Ashland's office.
7. EE will install slurry and calibrate flows.  Use small pump to keep
   slurry in barrell mixed.
8. EE will do 20 Cond'n Cycles with the Recipe - "NWPadBP3"
   (Note the Default # of Cond'n cycles is now 2).
________________________________________________________________________________
There was some concern as to weather the carriers should be changed with the 
pad. The counts were above 400. I had them changed with the pad. You didn't 
mention anything special about the carriers so I didn't think there would be 
a problem. 

The pad and carrier change was completed at ~1:30AM 5/8/97. EE have placed the 
pad in the backroom labeled HOLD FOR BRYAN POLLARD. 
________________________________________________________________________________

9. Ops/PSS will do 20 Breakins with recipe "NwPadBP3".
________________________________________________________________________________
Done. This was started at ~2:30AM. It finished at ~4:00AM. There were no 
alarms/problems with the breakins. 
________________________________________________________________________________

10. Ops/PSS will do an MQC run with recipe "MQCBP2".
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run (LVNE) as an MQC with Particles also.
________________________________________________________________________________
MQC was started at ~4:05AM. Wafers were polished with no problems(ie alarms)
Removal rate and Std.Dev was L=1385/17.8, R=1244/17.9. Particle results were: 

	Left	Right
	----	-----
  7)	79	75
Sum)	379	342
Scr)	0	1.48

The file numbers for this run in the UV1050 are (Pre->L=2818, R=2819)(Post->
L=2824, R=2825). 

Since the Std.Dev was high (> 10%) I'm going to step 11 and running 12 more 
breakin wafers. 
________________________________________________________________________________

11. If uniformity is > 10% then do 12 more Bkins on "NwPadBP3" then go to
    #12 - otherwise just go to #12 with no additional Bkins.
________________________________________________________________________________
Done. 12 more breakins were started at 4:45am. I let them run and then let the 
system sit in wet idle for ~30 minutes. 
________________________________________________________________________________

12. Ops/PSS will do an MQC run with recipe "MQCBP2".
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run (LVNE) as an MQC with Particles also.
________________________________________________________________________________
I ran two breakin wafers(NwPadBP3) after letting the system sit for ~30 minutes.
The pad count at this time is 36 wafers. Thickness and Std.Dev readings were 
L=1418/16.7 and R=1221/16.3. Particle results were::

	Left	Right
	----	-----
  7)	66	72
Sum)	369	365
Scr)	0	16.8(I've saved this wafer for you to look at. 

Instead of going right to another set of MQC's I requested another set of 
Breakin wafers. There's currently 12 going now. 
________________________________________________________________________________

13. Ops/PSS will do an MQC run with recipe "MQCBP2".
   - Save Delta Tx map in Thickness.A4.
   - Log OUT run (LVNE) as an MQC with Particles also.

I will continue testing to Optimize for particles/scratches in the morning.

NOTE this is very urgent as Lot JL1509 is on HOLD until I get an 18% process
in place to split it.

Thanks,
Bryan


7.373A2 plans for tonight - Split lot is runningSTRATA::POLLARDThu May 08 1997 19:1224
Finished up the Conditioning tests and all looked good (up to 500waf).

We've taken all day to get the 18% in spec for uniformity.  It was
very Center fast - it finally came in after we changed the Right Carrier
(which didn't improve all day as the Left one did).

Rich and I started the split lot - JL1509.  The send aheads are running on
CMP.A2 for the 18% split.  Rich will target the rest of the 18% split and
finish it up.  The std portion can be run on CMP.A1 (or CMP.A2 when everything
gets changed back to normal).

For EE to get CMP.A2 back to normal the following will be done:
1. Remove 18% slurry and reconnect STD slurry.
2. Change Pad (Please SAVE this PAD also as it acted weird!)
3. Change 4" Disk back to STD 6" Disk.
4. Change the following variables - 
   Var 126 "Prog Cond Offset" back to 96,500.
   Var 71 back to 1.55
7. Do std startup procedures for New Pad. - Bk-ins and MQC.

Thanks,
Bryan P.


7.374Eng. work completed, monthly completedSUBPAC::BJUBINVILLECMP Equipment EngineeringFri May 09 1997 06:0714
    
         Removed 18% barrel and hooked tool back up to house slurry.

         Completed monthly PM

         changed the variables back to normal per P.E. and put
         the 6" disc back on.  Unfortunately, all of the sweep 
         parameters and ID/OD parameters were changed and the 
         conditioner crashed into the slurry distribution tubes.  
         After the sweep variable were set to match A1, the tool 
         conditioned fine.  There was no apparent damage to the 
         conditioner or tubes.
    
        
7.375Scratches on Wafers / Cleaning Station Air MotorsASDG::POIRIERFri May 09 1997 20:338
    Scratches / Air Motors:
    ----------------------
    The system was scratching wafers this morning. Both cleaning station
    air motors were not rotating. Since there was only 1 motor in stock,
    the system will remain down until the 2nd motor arrives via FedEx 
    tomorrow morning.
    
                                                              C Shift.
7.376Sweep parametersSUBPAC::BJUBINVILLECMP Equipment EngineeringFri May 09 1997 20:428
    
        re - .374
    
            When you have a recipe called up, the sweep parameters 
         are set as well under Sweep Param.  All we had to do was
         select a normal 6" conditioner recipe such as MQC and the
         sweep recipe would have been fine.
                                               BJ
7.37718% Game PlanSTRATA::POLLARDTue May 13 1997 18:3230
Game plan for 18% Split lots JD1453 and JD1457.

Today we hooked up the 18% SC1 and tried to get it up with the old pad
but uniformity was 25% Edge Fast so we did a pad change.  After Bkins, the
first MQC was high at 16 and 17% Ctr Fast unif. (Removal Rates were ~ 1400A - 
particles good at < 100). So we did additional Bkins and another MQC.
The results of the 2nd MQC were also bad with 15% Ctr Slow unif.  So it
looks like we missed the whole window of opportunity.

So we're going to start over with a new pad and new carriers.
1. Regular slurry will be hooked back up to do Std Breakins and MQC.
2. Do Breakins and MQC until the Uniformity is in SPEC.
3. EE will hook up 18% slurry (keep same pad).
4. Do another std MQC with the 18% slurry.  If all is in spec then run
   the lots as described below (the uniformity can be as high as 11% on this
   18% run).

The 2 lots (JD1453, JD1457) need to be split on the CMP.A tools as 
follows (per Remark). Also note the lots need to be taken out of stores:

CMP.A1 - STD - wafers 1,2,5,6,9,10,13,14,17,18,21,22
CMP.A2 - 18% - wafers 3,4,7,8,11,12,15,16,19,20,23,24

NOTE - the wafers from JD1453 have been sorted with the 18% wafers in the
       cassette next to CMP.A2 and the std wafers in the POD.

Thanks,
Bryan


7.378Bad uniformity.... GONE!!SUBPAC::BJUBINVILLECMP Equipment EngineeringWed May 14 1997 07:1711
    
    
        Tool failed 3 times for uniformity with 12% slurry.
     The last straw was when the tool has 400A removal rate.
     Went to change the pad and it came off with two fingers.
     Replaced pad and the tool finally passed MQC's after 
     two rounds of MQC's.  The tool was then swapped over
     to 18% slurry and flow was calibrated.  It now sits
     in WET ILDE, ready for the next round of MQC's.
    
                                               D shift
7.379Results of Bryans gameplanYIELD::SHONGWed May 14 1997 08:2258
Game plan for 18% Split lots JD1453 and JD1457.

Today we hooked up the 18% SC1 and tried to get it up with the old pad
but uniformity was 25% Edge Fast so we did a pad change.  After Bkins, the
first MQC was high at 16 and 17% Ctr Fast unif. (Removal Rates were ~ 1400A - 
particles good at < 100). So we did additional Bkins and another MQC.
The results of the 2nd MQC were also bad with 15% Ctr Slow unif.  So it
looks like we missed the whole window of opportunity.

So we're going to start over with a new pad and new carriers.
1. Regular slurry will be hooked back up to do Std Breakins and MQC.
________________________________________________________________________________
Done. This was done on A-Shift and was ready for MQC's at ~8:30PM. 
________________________________________________________________________________

2. Do Breakins and MQC until the Uniformity is in SPEC.
________________________________________________________________________________
A full boat(24) of breakins were ran first. MQC's failed on the first try with 
high Std.Dev (~11%). I checked the map out and the tool is polishing edge fast. 
Particles were good. Another set of 12 breakins were ran. Second set of MQC's 
were ran and the removal rate for the second run was real bad ~350� for the 
right carrier. We took a look at the pad and noticed the top pad was not stick-
ing well to the bottom pad. It was changed again and another set of MQC's were 
started.

After the pad change we ran the conditioner for a while and then ran 10 break-
ins. MQC's were ran andd failed for high Std.Dev ~12% for the right carrier. 
Ran some more breakins and MQC's. This time the Std.Dev came in at ~6%L and 9%R.
I turned the tool over for the 18% slurry hook up.
________________________________________________________________________________

3. EE will hook up 18% slurry (keep same pad).
________________________________________________________________________________
Done. EE hooked the 18% slurry tank up and flushed the 12.8% line. 
________________________________________________________________________________

4. Do another std MQC with the 18% slurry.  If all is in spec then run
   the lots as described below (the uniformity can be as high as 11% on this
   18% run).
________________________________________________________________________________
I requested 6 breakins be ran prior to running the MQC monitors. There currently
in process. Results should be available after passdown. 
________________________________________________________________________________

The 2 lots (JD1453, JD1457) need to be split on the CMP.A tools as 
follows (per Remark). Also note the lots need to be taken out of stores:

CMP.A1 - STD - wafers 1,2,5,6,9,10,13,14,17,18,21,22
CMP.A2 - 18% - wafers 3,4,7,8,11,12,15,16,19,20,23,24

NOTE - the wafers from JD1453 have been sorted with the 18% wafers in the
       cassette next to CMP.A2 and the std wafers in the POD.

Thanks,
Bryan



7.380CMP.A2 Plans - Splits finishing upSTRATA::POLLARDWed May 14 1997 17:5330
The 18% splits for JD1453 and JD1457 have been completed.  
Send aheads for the std portion of JD1457 were processed and it was found
out that the wafers had already been polished on Saturday (they had scrapped
2w at that time due to overpolish) and now 2 more wafers were scrapped.
Rich is doing 100% readings to be sure the lot is set.
Also Rich will read the std portion of JD1453 to be sure it set - if not
he will process these wafers on CMP.A2 and then CMP.A2 will need to go
down as explained below.

On the Tool side of things CMP.A2 didn't do so Hot today either.  The 
Head-to-Head variation was very high on the 18% slurry-
  L- 1187A 9.8%  R- 1660A 6.1%
So once we converted back to std slurry another MQC was done - results were:
  L- 1167A 14.5% R- 1619A 9.8%

The carriers were swapped to see if the problem was just the carrier-
results were:
  L- 1696A 14.5% R- 1059A 9.6%

So unexpectedly the rates went with the carrier and the uniformity stayed 
with the splindle.  So AFTER the work with JD1453 is completed the following
needs to be done:
1. Std Pad Change
2. Change Both Carriers - and SAVE both Carriers with a LABEL to 
   HOLD for Jim Tooker.
3. Do Std Breakins and MQC's.

/Bryan Pollard

7.381wafers left on padFABSIX::J_SWENSONThu May 15 1997 20:2916
    Wafer being left on pad
    -----------------------
    
    
    - right carrier leaving wafers on pad.
    
      . found bad check valve (replaced)
    
      . replaced three plate screws
    
      . found white o-ring to be broken (replaced)
    
      . returned tool back to production.
    
    
                                          c-shift
7.382pad changeSUBPAC::BJUBINVILLECMP Equipment EngineeringFri May 16 1997 05:585
    
        Replaced pad because it failed MQC's for bad unif. and
     had a bubble
    
                                            D shift
7.383cmp.a2 pad/carrier changeFABSIX::J_SADINFreedom isn&#039;t free.Sun May 18 1997 19:5510
    
    
    	Performed weekly pm during routine 24hr pad change (after failed
    mqcs). Changed pad and carriers. Tool then proceeded to fail 3 sets of
    mqcs (high unif). Replaced pad with a fresh one and mqcs are running
    now.
    
    
    	a-team
    
7.384Important Work for CMP.A tools Sun Nite/Mon Morn.STRATA::POLLARDFri May 30 1997 17:4429
**** IMPORTANT Info for Sunday night B shift on CMP.A tools! *******

Please DO the following if a normal PAD change is required on a CMP.A tool
on Sunday Night/Monday morning AFTER 2am:
  (this is in preparation to convert to the new 4" conditioning recipes which
   will be turned on by TDCO Monday during the day)

1. Install the 4" Conditioning Disk in place of 6"
2. Change the following variables:  Var#126- 95,000     Var#71- 0.59
3. Install New Pad as normal.
4. Use Recipe "CRFSTBKN" to do 10 Conditioning Sweeps
   (set # Sweeps back to 2 when finished).
5. Use Recipe "CRFSTBKN" to do 10 Break-ins.
6. On CMP.A1 use Recipe "MQCBP" to do MQC.
   On CMP.A2 use Recipe "MQCBP3" to do MQC.
7. If MQC's pass leave up for Bryan to put TDCO on-line at 8am.
   If fails go to #8.
8. If MQC's failed then evaluate whether the removal rate is Center Fast or
   Center Slow:  
   - If it is Center Fast Repeat #4 and #5 and #6  - if it passes
     system will wait for Bryan to put on TDCO at 8am.  If it fails start
     with another Pad.
   - If it is Center Slow do 10 conditioning cycles with Recipe "CRSLWBKN"
     and do 10 Breakin wafers with the same recipe then repeat #6.  If
     it passes system will wait for Bryan to put on TDCO at 8am.  If it fails
     start with another Pad.

Bryan Pollard

7.385TDCO for New Cond'n & BreakinsFABSIX::B_POLLARDMon Jun 02 1997 16:57106
Please read the following TDCO which went on-line today for the New 
Conditioning cycles and the New Breakin Procedure for New Pads.
Thanks
Bryan Pollard

DISK$DCS2:[UDOC.QUAL]BK20-00000-0006.T   REV L    11-JUN-1996   PAGE  10 OF  13

| 6.1  Temporary D.C.O. Form

  Originator__Bryan Pollard____      TEMPORARY       TDCO # _________________
  Tel Ext__4036_Date__5/29/97__       DOCUMENT       Copy # _____   of ______
                                    CHANGE ORDER     Date Of Issue___________
_______________________________________________________________________________
 PROBLEM:                                             |1ST ISSUE APPROVAL STAMP
Conditioning part of all recipes have been updated for|EXPIRATION DATE
4" disk.  This change will extend the pad life from   |________
the typical 250-300 wafers up to 500 wafers.          |
Also changes have been made to break-in recipes.      |
______________________________________________________|_________________________
DESCRIPTION OF CHANGE:                                |2ND ISSUE APPROVAL STAMP
The conditioning sweep parameters (which are not      |EXPIRATION DATE 
typically seen by operators) have been updated for all|________
production recipes - MQC, PMD, ILD, STI.  The sweep   |
has also been changed from 1 to 2 & from Post         |
to Pre Condition.                                     |
                                                      |
The limit for Pad Life is extended to 500 wafers      |________________________
(with 24 hour MQC's in Spec - Section 4.1.11)         |3RD ISSUE APPROVAL STAMP
                                                      |EXPIRATION DATE 
Also changes have been to the Pad Breakin procedures -|________
these apply to section 3.1.1 and to various parts of  |
4.1 of the spec and are described on the second page. |
                                                      |
______________________________________________________|_______________________
DURATION OF TDCO OR NUMBER OF LOTS TO BE PROCESSED: 30 days
______________________________________________________________________________
DOCUMENT # / TITLE AFFECTED                  
HF6O-0069  /  STRAUSBAUGH 6DS-SP (CMP.A1, CMP.A2) Chemical Mechanical 
Planarization Operations Specification.
_______________________________________________________________________________
1ST ISSUE APPROVALS:| 2ND ISSUE APPROVAL | 3RD ISSUE APPROVAL | EARLY REMOVAL
                    |                    |                    | AUTHORIZATION
SIGNATURE/BADGE/DATE|SIGNATURE/BADGE/DATE|SIGNATURE/BADGE/DATE|SIGN./BADGE/DATE
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
IF LOT SPECIFIC: Upon completion of the lots, the originator must remove, sign, 
date this TDCO form and return it Document Control (HLO1-1/P5).

____________________              ____________              ___________________
    ORIGINATOR                        DATE                  DATE LOTS COMPLETED
                               ---------------
                               |d|i|g|i|t|a|l|
                               ---------------

DISK$DCS2:[UDOC.QUAL]BK20-00000-0006.T   REV L    11-JUN-1996   PAGE  10 OF  13

| 6.1  Temporary D.C.O. Form   
  
  Originator__Bryan Pollard____      TEMPORARY       TDCO # _________________
  Tel Ext__4036_Date__5/29/97__       DOCUMENT       Copy # _____   of ______
  Page 2                              CHANGE ORDER     Date Of Issue___________
_______________________________________________________________________________

The Break-In recipes and procedures for breaking in a New Pad have been altered-
the new procedure will be as follows (applies to section 4.1.6- 4.1.14):
  1. Run 10 Conditioning cycles with the recipe "CRFSTBKN" (stands for
     CenterFast Break-in - most new pads start with a Center Fast polish).
  2. Run 10 - 50K oxide wafers thru the recipe "CRFSTBKN".
  3. Do monitor run with recipe "MQC"
  4. If monitor fails uniformity then do the following:
     - If monitor is Center Fast - run 10 more conditioning cycles with 
       the recipe "CRFSTBKN" then do 10 more 50K oxide dummies with the same 
       recipe and repeat MQC.  If second MQC fails change the pad and start 
       over.  
     - If monitor is Center Slow - run 10 conditioning cycles with the recipe
       "CRSLWBKN" then do 10 more 50K oxide wafer with the same recipe and 
       repeat MQC. If second MQC fails change the pad and start over.
     - For either case, record in CMP Notesfile how many breakins are required 
       to get the uniformity in spec.

The Break-In recipes and procedures for breaking in a USED Pad have also been 
altered- the new procedure will be as follows (applies to section 4.1.6 - 
4.1.14):
   1. All breakins which previously were done with the "USEDPAD" recipe will
      be done with the "CRSLWBKN" recipe except if an MQC has failed
      due to being Center Fast then use the CRFSTBKN recipe.
   2. If a 24 hour MQC fails and is Center Slow then do 20 conditioning
      cycles with the "CRSLWBKN" recipe and repeat MQC - if it passes system
      is up if it fails then the system will require a pad change.
7.386set up for 4" conditioningFABSIX::J_SADINFreedom isn&#039;t free.Mon Jun 02 1997 20:039
    
    
    	set up for 4" conditioner. Just had to change the variables (126
    and 71) to support 4" conditioning. 
    
    
    	a-shift
    
    
7.387Please Save the Pads!STRATA::POLLARDWed Jun 04 1997 10:416
NOTE - During the time of the 4" Disk TDCO - Please SAVE a full-length 
       cross-section of all pads on the A tools.  Label them with the tool,
       # of wafers, Date, and "Hold for Bryan Pollard".

Thanks
Bryan Pollard
7.388Pad change - Bad uniformitySUBPAC::BJUBINVILLECMP Equipment EngineeringThu Jun 05 1997 07:135
    
          Pad change due to poor uniformity.  Pad was replaced
      and saved for Bryan.
    
                                     D shift